### v3

# A Comprehensive Study on Large Language Models for Mutation Testing

Abstract.
Large Language Models (LLMs) have recently been used to generate mutants in both research work and in industrial practice. However, there has been no comprehensive empirical study of their performance for this increasingly important LLM-based Software Engineering application. To address this, we report the results of a comprehensive empirical study over six different LLMs, including both state-of-the-art open- and closed-source models, on 851 real bugs drawn from two different Java real-world bug benchmarks. Our results reveal that, compared to existing rule-based approaches, LLMs generate more diverse mutants, that are behaviorally closer to real bugs and, most importantly, with higher fault detection. That is, 79.1% (for LLMs) vs. 41.6% (for rule-based); an increase of percentage points. Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by , , and percentage points respectively. These findings are immediately actionable for both research and practice. They allow practitioners to have greater confidence in deploying LLM-based mutation, while researchers now have a baseline for the state-of-the-art, with which they can research techniques to further improve effectiveness and reduce cost.
1. Introduction
For many years, mutation testing has been the topic of scientific research (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). Nevertheless, until the last five years, it had made little penetration into the industrial sector. In those most recent five years, it has become increasingly widely deployed in industry. For example, large tech sector companies such as Meta (Foster et al., 2025; Beller et al., 2021) and Google (Petrović et al., 2021) reported on results from deployment of mutation testing at their companies. In particular, results recently reported by Meta (Foster et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
This recent trend towards LLM-based mutant generation traces back to work on earlier learning-based approaches, which leverage deep learning models to generate mutations by learning patterns from historical bug fixes (Beller et al., 2021; Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., 2025). These methods aim to introduce mutations that better resemble real faults, potentially increasing their relevance for testing.
However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, there is to date, no large-scale comprehensive empirical study. Without such a study, we have no basis on which to deploy an industrial scale with confidence. Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutation generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
The generation of mutants, using any technique is challenging because of the following three important issues:
Generating effective mutations: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019). While existing methods have shown promise, there is still significant room for improvement in enhancing their fault-revealing potential (Degiovanni and Papadakis, 2022).
Generating valid mutations.: Mutants must be executable against test cases, which allows tests to assess the program’s behavior in the presence of mutations, to judge whether the mutants are killed. Ensuring executable mutations is challenging because mutations must remain syntactically and semantically valid, respect type constraints, preserve control and data flow, avoid breaking external dependencies, prevent unintended side effects, and align with test case assumptions to ensure meaningful evaluation (Jia and Harman, 2010).
Generating diverse and killable mutations: Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutations need to be diverse and cover the entire spectrum of code and program behaviors.
This paper reports on how well Large-Language Models (LLMs) handle these challenges, compared to more traditional rule-based approaches. We conducted a comprehensive comparison between six LLMs (three closed-source commercial and three open-source ones) and four existing mutation testing techniques with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014a) and 246 bugs from ConDefects (Wu et al., 2024)). Overall, our experiments required the generation of 611,100 mutations.
Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutation rate, equivalent mutation rate, and efficiency). Additionally, we study how different prompt engineering strategies and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat--16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs. This leads to a higher fault detection rate than traditional approaches. For example, the mutants generated by GPT-4o-Mini lead to a fault detection rate of 90.8%, while the mutants generated by LEAM, PIT, and Major are 63.9%, 40.1%, and 66.9%, respectively. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutation rate, and equivalent mutation rate. In particular, Major excels with a 97.6% compilability rate, 0% duplicate mutations, and 2.0% equivalent mutants, while GPT-4o exhibits a compilability rate of 76.4%, a duplicate mutation rate of 7.8%, and an equivalent mutation rate of 3.1%, respectively.
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that GPT-4o-Mini and DeepSeek exhibit the greatest diversity, newly introducing 49 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 21. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations. Moreover, the distributions of AST edit distance of mutations compared to the original code also present that LLMs generate more balanced and expressive mutations. For instance, almost 70% of the mutations of Major have the AST edit distance of one, while for GPT-4o, only 50% of its mutations have the same distance.
We also categorise the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall into 9 compilation error types with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. Moreover, we further analyze the code context of the non-compilable mutations generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features. This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, which could enhance the effectiveness of mutation testing using AI models.
Our implementation and the experimental data are available at our repo: https://github.com/da-123-snake/Mutantion-Generation-Study.
In summary, our paper makes the following main contributions:
We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our findings indicate that LLMs excel in generating diverse mutations that closely mimic real bugs.
We analyze the error types of non-compilable mutations and determine that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
We build a high-quality dataset of Java mutations comprehensively labeled with detailed quality and behavior metrics.
The remainder of this paper is organized as follows. Section 2 presents the design of our study, including the research questions, studied approaches, datasets, and metrics. Section 3 presents the empirical evaluation of these generation approaches and answers to each research question. Section 4 discusses the influences of different experimental setups and the implications summarized from the findings. Section 5 presents the threats to the validity of our study. Section 6 discusses the related work. Section 7 concludes this work.
2. Study Design
In this section, we introduce how our study is designed and conducted.
2.1. Overview and Research Questions
This paper intends to evaluate existing mutation generation approaches by answering the following research questions.
RQ1 (Effectiveness) How well do mutations generated by LLMs mimic real bugs compared to existing methods? This research question investigates the effectiveness of LLM-generated mutations in representing real bugs. Specifically, it evaluates whether these mutations exhibit similar characteristics to real faults in terms of detection capability, behavioral similarity, and diversity. This RQ is divided into the following sub-RQs:
RQ1.1: (Detectability) To what extent do bug-triggering tests successfully detect LLM-generated mutations?
RQ1.2 (Coupling) To what extent do LLM-generated mutations couple with real bugs?
RQ1.3 (Diversity) How diverse are LLM-generated mutations compared to traditional approaches?
RQ2 (Validity) How do LLMs compare to existing approaches in terms of validity? This research question explores the validity of mutations generated by LLMs compared to traditional approaches. Validity is assessed based on key factors such as compilability, duplication, and equivalence of generated mutations. By analyzing these aspects, we aim to determine whether LLMs produce valid and meaningful mutations that contribute to practical mutation testing. This RQ is divided into the following sub-RQs:
RQ2.1 (Compilability) What percentage of generated mutations are compilable across different approaches?
RQ2.2 (Duplication) What percentage of the generated mutations are duplicates of the original code across different approaches?
RQ2.3 (Equivalence) What percentage of generated mutations are equivalent across different approaches?
RQ3 (Efficiency) What are the time costs of different mutation generation approaches? This research question investigates the time costs of different mutation generation approaches. Time costs are associated with the efficiency and practicality of mutation generation approaches.
RQ4 (Compilation Error) What are the types of compilation errors of the non-compilable mutations of different approaches? This RQ aims to categorize and analyze the types of compilation errors that arise from non-compilable mutations generated by different approaches. Understanding these errors can guide mutation generation approaches to enhance their practical usability.
2.2. Studied Mutation Generation Approaches
Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking111https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutation generation approaches is for Java. We categorize these approaches into two groups: traditional approaches (i.e., those developed before the emergence of LLMs) and LLM-based approaches (i.e., approaches that leverage LLMs and are designed for interaction in human-like conversations), as shown below. Table 1 summarizes the mutation generation approaches covered by this study.
| Type | Name | Description |
| Traditional | PIT | Widely studied mutation testing framework with 29 mutation operators. |
| Major | Widely studied mutation testing framework with 11 mutation operators. | |
| LEAM | Custom-trained models learning from a large corpus of real bugs and patches. | |
| Bert | Mutation generation with the pre-trained model CodeBERT. | |
| LLM-based | GPT-3.5 | GPT-3.5-Turbo, whose based model is GPT, trained on data before 2021/09, released in 2021/03. |
| GPT-4o | The based model is GPT, trained on data before 2023/10, released in 2024/05. | |
| GPT-4o-M | GPT-4o-Mini, whose based model is GPT, trained on data before 2023/10, released in 2024/07. | |
| SC-16b | StarChat--16b, whose based model is StarCoder, released in 2023/06. | |
| CL-13b | CodeLlama-Instruct-13b, whose based model is Llama, released in 2023/08. | |
| DS-236b | DeepSeek-Coder-V2-236b, whose based model is DeepSeek, released in 2024/07. |
2.2.1. Traditional Approaches
For Java, there exist several mutation testing frameworks, such as JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), PIT (Coles et al., 2016), and Major (Just et al., 2011). As some tools are no longer maintained, following existing studies (Kaufman et al., 2022; Ojdanic et al., 2023; Kushigian et al., 2024), we employ PIT (Coles et al., 2016), Major (Just et al., 2011), LEAM (Tian et al., 2022), and Bert (Degiovanni and Papadakis, 2022). Major and PIT are widely studied rule-based approaches, while LEAM and Bert involve deep learning models.
PIT: PIT (Coles et al., 2016) is a widely used mutation testing tool on the Java bytecode level, which supports 29 well-designed mutation operators and four groups of active mutation operator sets. In our study, we employ PIT 1.14.4 and leverage the ALL group which activates all operators.
Major: Major (Just et al., 2011) is another high-impact mutation testing framework, that supports nine operators. In our study, we use Major 1.3.5 and involve all mutation operators.
LEAM: LEAM (Tian et al., 2022) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutations. Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutations.
Bert: Bert (Degiovanni and Papadakis, 2022) is the state-of-the-art mutation generation approaches based on the small-scale pre-trained model, Bert (Devlin et al., 2018). Equipped with the pre-trained model, Bert transforms the task of mutation generation to code completion under the surrounding context.
2.2.2. LLM-based Approaches
Our study involve multiple kinds of LLMs, including both open-source and closed-source models.
LLMs: Our study also comprehensively covers the most widely studied LLMs. Specifically, we select a representative set of LLMs with code capabilities both from commercial closed-source and open-source models. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5 (GPT-3.5-Turbo) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-M (GPT-4o-Mini) (Achiam et al., 2023)), and open-source LLMs, including SC-16b (StarChat--16b), CL-13b (CodeLlama-Instruct-13b), and DS-236b (DeepSeek-Coder-V2-236b). The details of these models are shown as Table 1. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 236 billion.
2.3. Datasets
| Dataset | Project | # of Bugs | Time Span |
| Defects4J (D4J) | Math | 106 | 2006/06/05 - 2013/08/31 |
| Lang | 65 | 2006/07/16 - 2013/07/07 | |
| Chart | 26 | 2007/07/06 - 2010/02/09 | |
| Time | 27 | 2010/10/27 - 2013/12/02 | |
| Closure | 133 | 2009/11/12 - 2013/10/23 | |
| Mockito | 38 | 2009/06/20 - 2015/05/20 | |
| Cli | 39 | 2007/05/15 - 2018/02/26 | |
| Codec | 18 | 2008/04/27 - 2017/03/26 | |
| Csv | 16 | 2012/03/27 - 2018/05/18 | |
| Gson | 18 | 2010/11/02 - 2017/09/21 | |
| JacksonCore | 26 | 2013/08/28 - 2019/04/05 | |
| Jsoup | 93 | 2011/07/02 - 2019/07/04 | |
| ConDefects (CD) | — | 246 | 2024/03/01 - 2024/06/30 |
| Total | — | 851 | 2006/07/16 - 2024/06/30 |
As one of the major tasks of our study is to evaluate the behavioral similarity between mutations and real-world bugs, we selected datasets that include real-world bugs along with their corresponding fault-revealing test cases. Specifically, we employ the datasets with the following properties:
The datasets should comprise Java programs, as existing methods are primarily based on Java, and we need to compare with them.
The bugs of the datasets should be real-world bugs so that we can compare the difference between mutations and real bugs.
Every bug in datasets has the correctly fixed version provided by developers so that we can mutate the fixed version and compare them with the corresponding real bugs.
Every bug is accompanied by at least one bug-triggering test because we need to measure whether the mutations affect the execution of bug-triggering tests.
To this end, we employ the Defects4J 2.0 (Just et al., 2014a) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 2. In total, we conducted experiments on 851 bugs.
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023; Just et al., 2014b; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains history bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from the Defects4J 2.0, which contains 605 bugs in total. However, from Table 1 and Table 2, we observe the periods of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder222https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs.
2.4. Evaluation Metrics
In this subsection, we present the evaluation metrics used in our study to assess the quality of the generated mutations comprehensively. We evaluate the mutations from multiple perspectives as before mentioned. Note that only relying on the mutation score is insufficient for a meaningful assessment. A higher mutation score does not necessarily indicate better quality mutations, as it only reflects the proportion of killed mutants rather than their effectiveness. However, we still report the mutation score of each mutation generation approach later in Section 3.
To better understand the metrics, we first illustrate the categories of mutations. Let all generated mutations be the set , where LLMs may generate syntactically incorrect mutations that can not pass compilation. Thus we denote the compilable mutations as the set . Note that the set contains two types of mutations that should be removed, the duplicate mutations (denoted as ) and equivalent mutations (denoted as ). Therefore, we have . The duplicate mutations refer to those identical to the original code or other mutations. In mutation testing, the ideal set of mutations to employ is from the set , excluding equivalent mutations that have no contributions to discovering the weakness of tests. However, determining the equivalent mutations set is undecidable, so we follow the common practice and use the set as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025).
2.4.1. RQ1: Effectiveness
Evaluating the effectiveness of mutation generation approaches requires assessing how well the generated mutations mimic real bugs. A key aspect of this evaluation is understanding whether LLM-generated mutations exhibit fault characteristics that align with real-world defects.
To quantify this effectiveness, we employ multiple metrics that capture different aspects of behavioral similarity between mutations and real bugs. These metrics assess whether bug-triggering tests can detect LLM-generated mutations (i.e., RQ1.1), how closely mutations resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutation generation techniques (i.e., RQ1.3). The metrics of these sub-RQs are listed as bellow.
RQ1.1: Real Bug Detection Rate: Following the approach of Just et al. (Just et al., 2014b), we evaluate the effectiveness of mutation generation by measuring its ability to detect real bugs using developer-written tests. For a given real bug (which may span multiple lines), we apply a mutation generation approach to create a set of mutations on the fixed version of the program. If the test cases that kill a mutation also fail on the real bug, we consider that the mutation generation approach has successfully “detected” this real bug. To quantify this effectiveness, we compute the Real Bug Detection Rate, which represents the proportion of real bugs for which at least one generated mutation is detected by the same test cases that reveal the real bug. This metric provides insight into how well LLM-generated mutations align with actual software defects in a testing context.
RQ1.2: Average Ochiai Coefficient: This metric estimates the semantic similarity between a mutation and its corresponding real bug based on their passing and failing test cases (Offutt and Untch, 2001). Ochiai coefficient measures how a mutation resembles its corresponding real bug. The Ochiai coefficient was originally introduced as a general similarity measure but later became widely adopted in spectrum-based fault localization (SBFL) (Zou et al., 2019). In this study, we use the Ochiai coefficient in its original form to assess how closely a mutation resembles its corresponding real bug, which is consistent with the study of Ojdanic et al. (Ojdanic et al., 2023).
More concretely, let and be a mutation and its corresponding real bug, and let and be the set of their failing test cases respectively. The coefficient of is computed by the following formula:
Therefore, if the Ochiai coefficient of is , it is semantically equal to its corresponding bug . Given a bug , its Ochiai coefficient is calculated as the mean Ochiai coefficient across all its corresponding de-duplicated mutations (i.e., ), shown as the following formula:
Given a mutation generation approach and a set of bugs , the Average Ochiai Coefficient of the approach is calculated as the mean Ochiai coefficient across all bugs, shown as the following formula:
RQ1.2: High Ochiai Bug Count: Following the study (Ojdanic et al., 2023), a bug and a set of mutations are considered to have high Ochiai if they satisfy . This threshold indicates a strong semantic similarity between the generated mutations and the real bug. For each mutation generation approach, we compute the number of bugs that meet this criterion.
RQ1.2: Coupling Rate: Given a real bug, one mutant is considered coupling if the set of test cases that kill the mutant overlaps with the bug-triggering tests. Note that each bug we analyzed has at least one failing test to reveal the bug, called bug-triggering test. The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutations (i.e., ). Note that Real Bug Detection Rate estimates the effectiveness via the proportion of detected bugs, which Coupling Rate estimates it via the proportion of mutations.
RQ1.2: Coupling Categorization: We categorize each generated mutation into distinct coupling scales based on their similarity to real bugs, following the classification proposed by (Gay and Salahirad, 2023), and analyze their distribution across different approaches. Specifically, given a compilable mutation (i.e., an element from the set ) on the fixed version of a real bug (which can be uncovered by at least one triggering test), we can classify it into the following six scales based on its behavior:
Strong Substitution: All triggering tests fail and no additional tests fail, which represents the mutation is a semantic replacement of the real bug under the given test suite.
Strong Substitution + Additional Tests Fail: All triggering tests fail but additional non-triggering tests fail.
Partial Substitution: Some triggering tests fail and no additional tests fail.
Partial Substitution + Additional Tests Fail: Some triggering tests fail and some additional non-triggering tests fail.
Not Substitution: Only non-triggering tests fail.
Not Detected: No tests fail (i.e., the mutation is alive under the given test suite).
Later, we will classify the mutations of each approach and analyze their distributions.
RQ1.3: AST Node Diversity: We prefer the set of mutations with diverse syntactic forms, i.e., method invocations, field accesses, etc. Therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. The approach with a greater variety of AST node types is considered superior, as it indicates a broader exploration of the syntactic space and enhances the likelihood of detecting weaknesses in the test suite.
RQ1.3: AST Edit Distance: The AST edit distance measures the structural difference between the original and mutated code by quantifying the number of tree transformation operations (i.e., insertions, deletions, and modifications) needed to convert one AST into another. This metric evaluates the extent of syntactic changes introduced by mutations. A higher edit distance suggests more significant syntactic changes.
2.4.2. RQ2: Validity
We consider three key metrics to evaluate the validity of different mutation generation approaches: generation efficiency, compilability, and the proportion of duplicate and equivalent mutations, providing a basic understanding of the validity of each approach.
RQ2.1: Compilability Rate: the proportion of mutations that successfully compile, is computed as:
Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher comparability rate.
RQ2.2: Duplicate Mutation Rate: the proportion of duplicate mutations, computed as:
LLMs sometimes generate duplicate mutations, i.e., mutations that are exactly the same as the original code. While traditional approaches rarely generate duplicate mutations, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data.
RQ2.3: Equivalent Mutation Rate: the proportion of equivalent mutation, computed as:
Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. We prefer the approach with a lower EMR.
2.4.3. RQ3: Efficiency
Efficiency describes how fast a mutation approach produces mutations. Mutation testing requires numerous mutations to evaluate the robustness of a test suite effectively. Therefore, the efficiency of mutation generation directly impacts the feasibility and scalability of mutation testing. To assess this, we compare these approaches based on time costs.
RQ3: Average Generation Time: measures the average time required to generate each mutation, providing a quantitative evaluation of the efficiency of each approach.
2.5. Mutation Generation
In this subsection, we discuss the setups for mutation generation approaches, especially the setups for LLM-based generation. We performed our experimental analysis on four rented cloud servers, each equipped with 2 NVIDIA GeForce RTX 3090 Ti, 168GB memory, and a 64 Core CPU Intel(R) Xeon(R) Platinum 8350C.
2.5.1. Mutation Generation via Traditional Approaches
For the well-established mutation generation approaches (i.e., PIT, Major, LEAM, and Bert), we activate all their mutation operators. Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutations within the context of real-world bugs. Note that for all mutation generation approaches, we employ the same setting for the code context.
2.5.2. Mutation Generation via LLMs
Recently, LLMs have shown great potential in numerous code-related tasks, to better understand their ability in mutation testing, we also study LLMs for generating mutations.
LLMs operate based on prompts, which consist of natural language instructions, often accompanied by code snippets. LLMs are driven by prompts, which are natural language instructions maybe with code. To effectively leverage LLMs for mutation generation, we design a well-crafted prompt template, shown as Figure 1. We follow the best practices suggesting prompts should comprise four aspects, namely, Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. In the Context, we clearly state that mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments. In addition, our prompts need few-shot examples, which should originate from real bugs enabling LLMs to learn how to mutate code from real bugs. This allows the model to understand the context and logic behind the changes, improving its ability to generate relevant and effective mutations.
To avoid the few-shot examples leaking information, beyond Defects4J and ConDefects, we employ another benchmark QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Referring to the study by Deng et al. (Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples. Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 3.
To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016). We use PIT-generated mutations with the universal set of operators and find that PIT generates 1.15 mutations per line of code. Therefore, based on the given code context, we generate one mutation per line. We explore the influence of context length on mutation generation in Section 4. Additionally, for the temperature, we use the default value for each LLM.
| Correct Version | Buggy Version |
| n = (n & (n - 1)); | n = (n ^ (n - 1)); |
| while (!queue.isEmpty()) | while (true) |
| return depth==0; | return true; |
ArrayList r = new ArrayList();
r.add(first).addll(subset); to_add(r); |
to_add.addAll(subset); |
| c = bin_op.apply(b,a); | c = bin_op.apply(a,b); |
| while(Math.abs(x-approx*approx)>epsilon) | while(Math.abs(x-approx)>epsilon) |
3. Evaluation Results
In this section, we aim to comprehensively evaluate these approaches via various aspects. Table 4 presents the overall performance of each generation approach. In terms of the number of mutations generated (Mutation Count), PIT produces the largest number of mutations at 340,728, significantly more than other approaches, while the remaining approaches range from about twenty thousand to thirty thousand. In terms of mutation score, the traditional rule-based approaches (PIT and Major) are over 0.5, LEAM and Bert are over 0.6, and the LLMs from the GPT family are all over 0.7.
Note that the number of generated mutations and mutation scores are not quality metrics for evaluating these approaches. A large number of mutations may include redundant or equivalent mutants that do not contribute to mutation scores, while a high mutation score may be influenced by the ease with which mutants are detected rather than their resemblance to real faults.
| Metric | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Mut. # | 340728 | 20831 | 36505 | 12946 | 31656 | 31945 | 34417 | 34741 | 28058 | 27614 |
| Mut. Score | 0.541 | 0.521 | 0.648 | 0.677 | 0.702 | 0.737 | 0.708 | 0.714 | 0.507 | 0.669 |
| Real Bug Detec. | 40.1% | 66.9% | 63.9% | 60.5% | 87.0% | 89.9% | 90.8% | 89.4% | 38.5% | 71.2% |
| Avg. Ochiai | 23.0% | 35.5% | 32.4% | 29.3% | 47.7% | 56.2% | 54.1% | 50.9% | 26.0% | 33.3% |
| Coupling Rate | 23.5% | 38.9% | 38.1% | 47.3% | 46.9% | 50.6% | 49.5% | 49.5% | 34.0% | 45.6% |
| Comp. Rate | 100% | 97.6% | 44.5% | 26.5% | 62.5% | 76.4% | 75.3% | 75.8% | 16.4% | 69.3% |
| Dup. Mut. Rate | 0.0% | 0.0% | 2.3% | 1.7% | 10.3% | 7.8% | 7.0% | 7.7% | 12.8% | 37.1% |
| A. G. Time (s) | 0.02 | 0.08 | 3.06 | 2.34 | 1.79 | 1.31 | 1.65 | 4.25 | 7.53 | 9.06 |
Note: The mutation generation approach achieving the best performance in terms of a metric is highlighted in grey background color. GPT-4o-M: GPT-4o-Mini. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat--16b. CL-13b: CodeLlama-Instruct-13b.
3.1. RQ1: Performance on Effectiveness
To measure the effectiveness of mutation generation approaches, we evaluate them with three sub-RQs, which estimate their effectiveness via real bug detectability, coupling, and diversity, respectively.
3.1.1. RQ1.1 Detectability
As illustrated in Section 2.4.1, we estimate real bug detectability via the following metric.
Real Bug Real Bug Detection Rate: The row Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutation generation approaches. Among the traditional approaches, PIT achieves a detection rate of 40.1%, while Major, LEAM, and Bert perform significantly better with a detection rate of 66.9%, 63.9%, and 60.5%, respectively. For LLM-based approaches, the results are significantly more effective than those of traditional ones. GPT-4o-Mini achieves the highest detection rate at 90.8%, closely followed by GPT-4o and DeepSeek, with detection rates of 89.9% and 89.4%, respectively. Meanwhile, GPT-3.5 and CodeLlama achieve moderate detection rates of 87.0% and 71.2%, respectively.
| Project | Bug # | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Chart | 26 | 19 | 17 | 16 | 10 | 22 | 22 | 23 | 23 | 9 | 8 |
| Lang | 65 | 23 | 51 | 27 | 38 | 52 | 55 | 56 | 57 | 11 | 41 |
| Time | 27 | 12 | 23 | 18 | 15 | 25 | 24 | 26 | 22 | 10 | 17 |
| Math | 106 | 40 | 82 | 44 | 65 | 92 | 94 | 96 | 95 | 58 | 77 |
| Mockito | 38 | 14 | 28 | 27 | 5 | 30 | 29 | 36 | 34 | 16 | 21 |
| Closure | 133 | 38 | 92 | 122 | 103 | 130 | 131 | 131 | 132 | 79 | 114 |
| Cli | 39 | 6 | 11 | 27 | 29 | 37 | 39 | 39 | 39 | 2 | 34 |
| Codec | 18 | 6 | 11 | 12 | 11 | 14 | 16 | 17 | 17 | 1 | 11 |
| Csv | 16 | 7 | 13 | 13 | 7 | 15 | 16 | 16 | 16 | 2 | 13 |
| Gson | 18 | 6 | 9 | 11 | 12 | 13 | 16 | 16 | 15 | 1 | 10 |
| JacksonCore | 26 | 1 | 17 | 17 | 18 | 23 | 23 | 21 | 21 | 1 | 17 |
| Jsoup | 93 | 48 | 77 | 90 | 86 | 93 | 92 | 93 | 90 | 57 | 87 |
| ConDefects | 246 | 121 | 138 | 120 | 116 | 194 | 208 | 203 | 200 | 81 | 156 |
| ALL | 851 | 341 | 569 | 544 | 515 | 740 | 765 | 773 | 761 | 328 | 606 |
Note: The mutation generation approaches that achieve the best performance in one project are highlighted in grey background color.
To better understand the detection of the generation approaches, we analyze their project-wise detectability, shown as Table 5. For individual projects, LLM-based approaches consistently achieve top performance across multiple projects. Particularly, GPT-4o-Mini outperforms other approaches in several projects, including Chart, Time, Mockito, and Jsoup, where it detects the highest number of bugs. Similarly, DeepSeek and GPT-4o demonstrate strong performance, achieving the best results in projects such as Lang, Math, and Closure.
3.1.2. RQ1.2 Coupling
To assess the degree of coupling, we perform comprehensive experiments that include both an overall metric evaluation and a project-wise analysis.
Average Ochiai Coefficient: The row Avg. Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measures the semantic distance between mutations and real bugs. Among the traditional approaches, Major achieves an Ochiai coefficient of 35.5%, higher than the others. LEAM and BERT achieve coefficients of 32.4% and 29.3%, respectively, indicating that the best-performing rule-based approach, Major, outperforms both of them. The LLM-based approaches demonstrate significantly higher Ochiai coefficients, with GPT-4o achieving the highest value of 56.2%, followed closely by GPT-4o-Mini at 54.1%, and DeepSeek at 50.9%. Overall, GPT-4o and GPT-4o-Mini exhibit their superiority in generating mutations with smaller semantic distances with real bugs.
High Ochiai Bug Count: Following the study of Ojdanic et al. (Ojdanic et al., 2023), the mutation set with an Ochiai coefficient over 0.8 is considered to be of high semantic similarity with its corresponding real bug. For each project, we count the number of bugs that achieved a high Ochiai coefficient by each approach, shown as Table 6. The overall trend is similar with Real Bug Detectability, which indicates that the LLM-based appraoches, GPT-4o and GPT-4o-Mini, achieve the highest number of bugs across most projects. DeepSeek also exhibits strong performance, closely following GPT-4o in many cases. All traditional approaches exhibit significantly fewer high Ochiai bugs. Interestingly, among traditional approaches, Major outperforms LEAM and BERT across most projects.
| Project | Bug # | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Chart | 26 | 9 | 9 | 3 | 5 | 10 | 13 | 12 | 12 | 7 | 3 |
| Lang | 65 | 15 | 34 | 13 | 18 | 32 | 35 | 37 | 34 | 12 | 23 |
| Time | 27 | 8 | 14 | 8 | 4 | 8 | 12 | 12 | 11 | 4 | 7 |
| Math | 106 | 20 | 42 | 13 | 27 | 46 | 56 | 48 | 47 | 19 | 33 |
| Mockito | 38 | 1 | 12 | 5 | 3 | 14 | 13 | 17 | 14 | 3 | 3 |
| Closure | 133 | 12 | 25 | 40 | 14 | 38 | 69 | 56 | 45 | 21 | 26 |
| Cli | 39 | 3 | 3 | 13 | 8 | 18 | 22 | 23 | 20 | 4 | 10 |
| Codec | 18 | 3 | 1 | 3 | 3 | 2 | 8 | 6 | 7 | 3 | 2 |
| Csv | 16 | 4 | 2 | 1 | 2 | 4 | 10 | 10 | 8 | 3 | 1 |
| Gson | 18 | 2 | 1 | 4 | 3 | 2 | 8 | 7 | 6 | 3 | 2 |
| JacksonCore | 26 | 2 | 3 | 9 | 8 | 11 | 14 | 12 | 11 | 5 | 6 |
| Jsoup | 93 | 16 | 6 | 32 | 20 | 46 | 39 | 39 | 41 | 25 | 19 |
| ConDefects | 246 | 101 | 150 | 132 | 134 | 175 | 179 | 181 | 177 | 112 | 148 |
| ALL | 851 | 196 | 302 | 276 | 249 | 406 | 478 | 460 | 433 | 221 | 283 |
Note: The mutation generation approaches that achieve the best performance in one project are highlighted in grey background color.
Coupling Rate: The overall Coupling Rate for each mutation generation approach is shown in the last row of Table 4. The traditional approaches, PIT, Major, LEAM, and Bert, exhibit lower coupling rates than LLM-based approaches, with coupling rates of 23.5%, 38.9%, 38.1%, and 47.3%, respectively. Among the LLM-based approaches, GPT-4o achieves the highest coupling rate, with a value of 50.6%. It is followed closely by GPT-4o-Mini and DeepSeek, with coupling rates of 49.5% and 49.5%, respectively. Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly GPT-4o, in generating highly coupled mutations that better mimic real bugs.
| Category | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Strong | 10.1% | 9.6% | 4.9% | 7.8% | 7.0% | 6.8% | 6.2% | 6.4% | 8.0% | 6.8% |
| Strong Addi. | 4.7% | 5.8% | 4.3% | 5.2% | 6.7% | 10.2% | 12.1% | 10.3% | 5.7% | 6.0% |
| Partial | 1.8% | 1.1% | 0.5% | 0.8% | 1.2% | 1.2% | 0.8% | 1.0% | 2.7% | 0.9% |
| Partial Addi. | 6.9% | 22.4% | 28.4% | 33.5% | 32.0% | 32.5% | 30.4% | 31.8% | 17.6% | 31.9% |
| Not Subs. | 30.6% | 13.2% | 26.7% | 20.4% | 23.3% | 23.1% | 21.3% | 21.9% | 16.7% | 21.3% |
| Not Detected | 45.9% | 47.9% | 35.2% | 32.3% | 29.8% | 26.3% | 29.2% | 28.6% | 49.3% | 33.1% |
Note: Strong - Strong Substitution, Strong Addi. - Strong Substitution + Additional Failing Tests, Partial - Partial Substitution, Partial Addi. - Partial Substitution + Additional Failing Tests, Not Subs. - Not Substitution.
Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutation of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected. Note that the first four categories are used to calculate the Coupling Rate. Table 7 presents the distribution of each generation approach. Among the traditional approaches, PIT and Major exhibit a significantly higher proportion of mutants in the Not Detected category, with rates of 45.9% and 47.9%, respectively, indicating that a substantial number of their generated mutants fail to couple with real bugs. However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutations. In contrast, LLM-based approaches such as GPT-4o and GPT-4o-Mini, demonstrate stronger coupling capabilities across multiple categories. These approaches achieve significantly higher rates in the Partial Substitution + Additional Tests Fail category, with values of 32.5% and 30.4%, respectively. Furthermore, they outperform traditional methods in the Strong Substitution + Additional Tests Fail category, with GPT-4o and GPT-4o-Mini achieving 10.2% and 12.1%, respectively. Overall, while the traditional rule-based approaches generate a larger proportion of strong substitution mutations, LLM-based approaches demonstrate a stronger ability to generate diverse and highly coupled mutations.
3.1.3. RQ1.3 Diversity
We explore the syntactic diversity of mutations generated by each approach. Note that PIT works at the Java byte code level, we thus skip it in this RQ.
AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang333https://pypi.org/project/javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 illustrates the number of introduced AST node types, with traditional approaches highlighted in green and LLM-based approaches in blue. The figure reveals that LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches. The best LLM-based approach, DeepSeek, exhibits the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types. Additionally, all six LLMs are significantly higher than all other approaches in AST node types. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 3), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
| Top K | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Del. | 15.3% | 48.9% | 0.4% | 5.3% | 2.1% | 0.5% | 1.0% | 0.3% | 0.1% |
| 1st | ST(78.6%) | LT(23.2%) | LT(29.4%) | LT(39.2%) | LT(45.4%) | LT(43.3%) | LT(38.8%) | LT(22.7%) | LT(47.8%) |
| 2nd | LT(21.4%) | MR(14.4%) | SE(25.7%) | BO(17.3%) | BO(16.6%) | BO(21.6%) | BO(19.1%) | BO(13.8%) | BO(25.7%) |
| 3rd | —— | CT(9.6%) | MR(24.5%) | MI(9.9%) | CC(8.8%) | CC(7.6%) | MI(10.1%) | MR(13.1%) | MI(9.3%) |
Del.: The proportion of deletion mutations. LT: Literal, BO: BinaryOperation, MI: MethodInvocation, MR: MemberReference, SE: StatementExpression, ST: Statement.
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the top three most frequent AST node types. For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutation ratio, with 48.9%, followed by Major at 15.3%. In contrast, all LLM-based approaches generate deletion mutations at a rate below 6%, with three models producing less than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations. As shown later, we find that deletions are one of the key factors contributing to the lower compilation success rate.
Regarding AST node types, the most frequent new AST node type of all the LLM-based approaches is Literal (e.g., ). However, for the traditional approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutations are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion.
AST Edit Distance: To quantify the extent of syntactic modifications introduced by each mutation generation approach, we compute the AST edit distance for each mutation using Gumtree (Falleri et al., 2014). The distribution of edit distances for each approach is presented in Table 9, providing insights into the complexity and diversity of the generated mutations.
| Distance | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| 1 | 69.2% | 58.5% | 73.8% | 52.4% | 50.0% | 47.3% | 45.1% | 36.0% | 46.4% |
| 2 | 27.8% | 23.6% | 8.9% | 36.4% | 38.8% | 41.9% | 44.2% | 39.1% | 46.6% |
| 3 | 1.8% | 7.3% | 2.7% | 6.8% | 7.2% | 6.3% | 6.5% | 8.0% | 3.8% |
| 4 | 0.6% | 3.4% | 14.2% | 2.4% | 2.2% | 2.5% | 2.6% | 6.0% | 1.8% |
| 5 | 0.4% | 1.0% | 0.1% | 0.9% | 0.8% | 0.8% | 0.8% | 3.7% | 0.5% |
| 6 | 0.2% | 6.1% | 0.4% | 1.0% | 1.0% | 1.2% | 0.8% | 7.2% | 1.0% |
From the table, we can find that all traditional approaches predominantly generate small-scale changes. Specifically, the proportion of mutations with an edit distance of 1 for Major, LEAM, and Bert is 69.2%, 58.5%, and 73.8%, respectively, indicating they involve minimal syntactic changes to the original code in the majority of cases. In contrast, LLM-based approaches exhibit a broader range of edit distances, reflecting their capability to introduce more complex transformations. For example, GPT-4o exhibits a more balanced distribution, with 38.8% of its mutations having an edit distance of 2, 7.2% with a distance of 3, and 2.2% with a distance of 4, indicating it generates both minor and moderate code changes.
3.2. RQ2: Performance on Validity
To evaluate the validity of mutation generation approaches, we consider the aspects that may impact their validity, i.e., comparability rate, duplicate mutation rate, and equivalent mutation rate. This RQ is divided into the following sub-RQs.
3.2.1. RQ2.1: Compilability
The row Comp. Rate of Table 4 presents the overall Compilability Rate (CR) of each approach. Among the traditional approaches, the rule-based approaches, PIT and Major, demonstrate superior compilability, with PIT achieving a perfect 100% , and Major following closely at 97.60%. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true). PIT uses the mutation operators more restrictedly and previously filters out non-compilable mutations. In contrast, LEAM and Bert, have rates at 35.0% and 22.5%, respectively. For LLM-based approaches, there is a considerable variation in compilability. GPT-4o achieves the highest compilability among LLMs at 76.40%, followed closely by GPT-4o-Mini and DeepSeek, with rates of 75.30% and 75.80%, respectively. Other LLMs such as GPT-3.5 and CodeLlama show moderate s of 62.50% and 71.20%, respectively, whereas models like StarChat fall behind with a of 38.50%. Compared to small models (i.e., LEAM and Bert), GPTs and DeepSeek show significantly higher compilation success rates. For example, the compilability rate of GPT-4o is over 30% higher than that of LEAM and Bert.
3.2.2. RQ2.2: Duplicate Mutation Rate
As aforementioned, duplicate mutations refer to those that are redundant or identical to the original code, providing no meaningful contribution to the mutation testing process. Therefore, a lower Duplicate Mutation Rate () is desirable, as it indicates a more efficient mutation generation process with minimal redundancy. The row Dup. Mut. Rate of Table 4 presents the results.
Rule-based approaches, PIT and Major, achieve a perfect 0% . This is expected, as these approaches apply well-defined mutation rules that ensure each generated mutant introduces a meaningful change to the program. Other traditional approaches, i.e., LEAM and Bert, show relatively low s of 2.30% and 1.70%, respectively. In contrast, LLMs generate a significant portion of duplicate mutations, exhibiting a higher . For the LLMs from OpenAI, GPT-3.5, GPT-4o, and GPT-4o-Mini exhibit moderate s of 10.30%, 7.80%, and 7.00%, respectively. The open-source LLMs such as StarCoder and DeepSeek demonstrate higher s of 12.80% and 7.70%, respectively, while CodeLlama shows the highest at 37.10%.
3.2.3. RQ2.3: Equivalent Mutations
Given the undecidable nature of identifying equivalent mutations, we manually evaluate each mutation’s equivalence, following the methodology of the existing study (Kushigian et al., 2024). For each mutation generation approach, we randomly sample a subset of compilable mutations, ensuring a 95% confidence level with a 10% margin of error.
Two authors independently evaluate the equivalence of the selected mutations to minimize subjectivity. The procedure for checking a mutation involves first transplanting it into the original project and executing all tests. If it passes all tests, the author then manually verifies whether the mutation is semantically equivalent. Note that PIT does not generate source-level mutations, so its mutations must be manually extracted from Java bytecode and transformed back into source-level representations, which is significantly more time-consuming.
In total, 959 mutations were labeled, 508 of which were filtered by tests (i.e., killed), and the remaining 451 were manually checked. The entire process took approximately ten hours per person. The manual verification requires checking the code context, which may involve multiple classes and methods. For example, to check a mutation that replaces an integer with a method invocation, where the method is invoked by the this object but defined in the parent class, we should review the parent class file and verify whether its return value is consistently equal to the integer. We measure the consistency between the two authors by Cohen’s kappa coefficient () (Viera et al., 2005), achieving an overall value of 0.8, which indicates a high level of agreement. In cases of disagreement, both authors conducted thorough discussions to resolve contradictions and finally reach consistent labels. The labeling results are available at our repo.
| Metric | PIT | Major | LEAM | Bert | GPT-3.5 | GPT-4o | GPT-4o-Mini | DS-236b | SC-16b | CL-13b |
| #Sampled | 97 | 98 | 96 | 94 | 96 | 96 | 96 | 96 | 94 | 96 |
| #Test Filtered | 26 | 52 | 48 | 53 | 60 | 64 | 59 | 63 | 32 | 51 |
| #Eq. Mut | 1 | 2 | 2 | 6 | 4 | 3 | 3 | 5 | 10 | 6 |
| EMR | 1.0% | 2.0% | 2.1% | 6.4% | 4.2% | 3.1% | 3.1% | 5.2% | 10.6% | 6.3% |
Table 10 presents the estimated Equivalent Mutation Rate (), reporting the number of sampled mutations (#Sampled), the number of equivalent mutations we identified (#Eq. Mut), and the for each approach. Among the traditional approaches, PIT achieves the lowest of 1.0%, followed by Major with an of 2.0%, indicating that rule-based techniques produce fewer equivalent mutations. In contrast, the LLM-based approaches exhibit higher s, with variations across different models. StarChat has the highest s of 10.6%, while GPT-4o, GPT-4o-Mini, and DeepSeek show relatively lower s, ranging from 3.1% to 5.2%. Overall, although the OpenAI models such as GPT-4o and GPT-4o-Mini, tend to have slightly higher s than PIT and Major, the increase remains modest and within an acceptable range.
Note that the estimated of Major given by the recent study of Kushigian et al. (Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutation for manual judgment. We adopt the same experimental process and the result of Major is 2.0%, which is consistent with theirs considering the randomness introduced by sampling.
3.3. RQ3: Performance on Efficiency
To evaluate the performance on efficiency, we record the average generation time for each approach.
Average Generation Time: The row A. G. Time of Table 4 presents the results of Average Generation Time. Among the traditional appraoches, PIT demonstrates the fastest generation time at 0.02 seconds per mutation, followed by Major, which takes 0.08 seconds, making them the most efficient approaches. In contrast, LEAM and Bert take 3.06 seconds and 2.34 seconds, respectively. Among the LLMs, the LLMs of OpenAI (i.e., GPT-3.5, GPT-4o, and GPT-4o-Mini) show relatively faster generation, ranging from 1.31 to 1.79 seconds. While the open-source LLMs StarChat, CodeLlama, and DeepSeek require 7.53, 9.06, and 4.25 seconds, respectively. The results suggest that the rule-based approaches are the most time-efficient for generating mutations, which are tens to hundreds of times faster than other approaches.
3.4. RQ4: Categorizations of Non-Compilable Mutations
To answer this RQ, we first identify the error types of non-compilable mutations and then analyze the surrounding code context types that are error-prone to generate them.
3.4.1. Categorizations of compilation error types
Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 3.2, LLMs generate many non-compilable mutations. This RQ analyzes the types of errors and potential root causes of these non-compilable mutations. Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT models, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015). From the manual analysis of these non-compilable mutations, we identified nine distinct error types, shown as follows.
Usage of Unknown Methods: this type of error is due to invoking unknown methods under the code context.
Code Structural Destruction: this type of error is due to illegal code structures, such as unmatched parentheses or braces.
Incorrect Method Parameters: this type of error is due to invoking a method with illegal parameters, including invoking by wrong numbers of parameters or unmatched type.
Usage of Unknown Variables: this type of error is due to using unknown variables under the code context.
Usage of Unknown Types: this type of error is due to using unknown type names under the code context.
Type Mismatch: this type of error is due to using illegal types when applying a certain operation, such as using objects to perform arithmetic operations.
Incorrect Initialization: this type of error is due to using illegal forms of constrictors or deleting necessary initialization statements.
Incorrect Location: This type of error is due to a mutation being applied to an incorrect code location.
Incorrect Exceptions: this type of error is due to throwing illegal exceptions.
| ID | Error Type | % | ID | Error Type | % |
| 1 | Usage of Unknown Methods | 27.3% | 6 | Type Mismatch | 7.6% |
| 2 | Code Structural Destruction | 22.9% | 7 | Incorrect Initialization | 3.1% |
| 3 | Incorrect Method Parameters | 12.8% | 8 | Incorrect Location | 3.1% |
| 4 | Usage of Unknown Variables | 11.2% | 9 | Incorrect Exceptions | 2.3% |
| 5 | Usage of Unknown Types | 9.6% | — | — | — |
Table 11 presents the results. From the table, we observe that the most frequent error type, Usage of Unknown Methods, accounts for 27.3% of total errors, highlighting the challenge of hallucination in LLM-generated code (Huang et al., 2023). The second most common error, Code Structural Destruction, makes up 22.9%, indicating that maintaining syntactic correctness remains a persistent difficulty for LLMs. These findings underscore the need for further improvements in LLMs to enhance their reliability and accuracy in code generation.
3.4.2. Categorizations of code context of non-compilable mutations
To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by all nine mutation generation approaches in Table 4, as shown in Figure 3. For most approaches (except Major), the code locations with MemberReference and MethodInvocation are the top-2 prevent AST node types. For example, for GPT LLMs and DeepSeek, there about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compilable mutations. The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 1.11% (DeepSeek), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% (Bert), and 16.61% (Major) of all their non-compilable mutations, respectively. Thus, for LLMs, deletion is not the major reason for non-compilation. Thus for LLMs, deletion is not the major reason for non-compilation.
4. Discussion
In this section, we explore the impact of different experiment setups for LLMs, then we discuss the implications.
4.1. Sensitivity to Experimental Settings
We discuss the experiment setups that may affect the evaluation results, particularly the setups related to LLMs.
4.1.1. Setup of the Prompts
Prompts play a critical role in the performance of LLMs. Therefore, we aim to investigate how different prompts influence mutation generation effectiveness. Due to budget constraints, we selected a subset of 306 bugs from our dataset, comprising 10 randomly sampled bugs from each Defects4J project and all bugs from the ConDefects dataset.
For the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the open-source DeepSeek-Coder-V2-236b, then switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get four types of prompts, shown as follows:
Prompt 1 (P1): The Default Prompt. P1 is the default prompt, shown as Figure 1.
Prompt 2 (P2): P1 without Few-Shot Examples. P2 is based on P1 by removing all the few-shot examples.
Prompt 3 (P3): P2 without Whole Java Method. P3 is based on P2 by removing the surrounding Java method code, only keeping the target code element to be mutated.
Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of corresponding unit tests of the target method into Context.
| Metric | GPT-3.5 | DS-236b | ||||||
| P1 | P2 | P3 | P4 | P1 | P2 | P3 | P4 | |
| Mut. # | 3681 | 3688 | 3350 | 3210 | 3803 | 3845 | 3802 | 3804 |
| Mut. Score | 0.773 | 0.785 | 0.842 | 0.801 | 0.784 | 0.778 | 0.793 | 0.814 |
| Real Bug Detec. | 79.4% | 78.4% | 72.2% | 72.5% | 84.6% | 84.0% | 81.4% | 82.6% |
| Avg. Ochiai | 63.1% | 61.8% | 60.8% | 62.1% | 67.3% | 64.1% | 65.0% | 63.7% |
| Coupling Rate | 52.1% | 52.2% | 47.4% | 49.6% | 53.1% | 51.9% | 52.9% | 50.7% |
| Compilability Rate | 66.9% | 66.8% | 60.0% | 61.2% | 79.1% | 78.9% | 74.9% | 78.2% |
| Dup. Mut. Rate | 11.4% | 13.3% | 14.9% | 12.7% | 10.9% | 9.4% | 9.8% | 12.7% |
| Avg. Gen. Time | 1.789 | 1.782 | 1.809 | 1.884 | 4.252 | 4.254 | 4.199 | 4.397 |
Table 12 presents a comparative evaluation of different prompts for GPT-3.5 and DeepSeek across multiple metrics. Among the four prompts, P1 consistently demonstrates superior performance across both models, achieving the highest Compilability Rates (66.9% for GPT-3.5 and 79.1% for DeepSeek) and the best Real Bug Detectability Rates (79.4% and 84.6%, respectively). P1 also yields the highest Avgerage Ochiai Coefficient across both models. These results overall indicate that providing LLMs with well-structured and appropriate information significantly enhances their performance (e.g., P1), whereas excessive information can lead to a decline in effectiveness (e.g., P4).
We further explore the performance similarity of different prompts. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the top part of Table 14. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the -values are all below 0.05.
4.1.2. Experiment Setup of the Context Length
| Metric | 1-Line Context | 2-Line Context | 3-Line Context |
| Real Bug Dectc. | 85.0% | 83.3% | 83.3% |
| Avg. Ochiai | 33.3% | 35.0% | 30.0% |
| Coupling Rate | 36.6% | 35.2% | 37.1% |
| Compilability Rate | 62.9% | 63.2% | 67.2% |
| Dup. Mut. Rate | 10.8% | 12.5% | 11.9% |
| Settings | Spea. Coef. | -value | Pear. Coef. | -value |
| GPT-3.5 P1 v.s. GPT-3.5 P2 | 0.9762 | 3.31E-05 | 0.9999 | 1.35E-31 |
| GPT-3.5 P1 v.s. GPT-3.5 P3 | 0.9040 | 2.00E-03 | 0.9999 | 8.66E-23 |
| GPT-3.5 P1 v.s. GPT-3.5 P4 | 0.9520 | 2.60E-04 | 0.9999 | 1.45E-25 |
| DS-236b P1 v.s. DS-236b P2 | 1.0000 | 0.00E+00 | 0.9999 | 1.94E-31 |
| DS-236b P1 v.s. DS-236b P3 | 0.9999 | 3.31E-05 | 0.9999 | 3.71E-30 |
| DS-236b P1 v.s. DS-236b P4 | 0.9762 | 3.31E-05 | 0.9999 | 1.83E-27 |
| 1-Line v.s. 2-Line | 0.9762 | 3.31E-05 | 0.9980 | 2.12E-08 |
| 2-Line v.s. 3-Line | 0.9762 | 3.31E-05 | 0.9997 | 1.04E-10 |
| 1-Line v.s. 3-Line | 0.9524 | 2.60E-04 | 0.9978 | 2.69E-08 |
| QB-6a (Default) v.s. QB-3 | 0.9701 | 6.55E-05 | 0.9996 | 1.70E-10 |
| QB-6a (Default) v.s. QB-3b | 0.9762 | 3.31E-05 | 0.9984 | 1.00E-08 |
| QB-6a (Default) v.s QB-9 | 0.9762 | 3.31E-05 | 0.9985 | 7.98E-09 |
| QB-6a (Default) v.s D4J-6 | 0.9524 | 2.60E-04 | 0.9980 | 2.13E-08 |
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples. D4J-6 involves random 6 examples from Defects4J.
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 13. Similarly, the middle part of Table 14 shows the performance similarity among different contexts. We can see that their mutual similarity exceeds 0.95 and the -values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
4.1.3. Impact of Different Few-Shot Examples
| Metric | QB-6a (Default) | QB-3 | QB-6b | QB-9 | D4J-6 |
| Real Bug Det. | 0.917 | 0.933 | 0.917 | 0.917 | 0.900 |
| Avg. Ochiai | 0.467 | 0.450 | 0.434 | 0.434 | 0.367 |
| Coupling Rate | 0.136 | 0.128 | 0.132 | 0.121 | 0.139 |
| Comp. Rate | 65.7% | 62.7% | 63.9% | 61.8% | 62.1% |
| Dup Mut. Rate | 13.7% | 12.8% | 12.8% | 13.8% | 12.5% |
| Equivalent Mutation Rate | 1.3% | 1.6% | 1.3% | 1.1% | 1.6% |
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples. D4J-6 involves random 6 examples from Defects4J.
To measure the impact of different few-shot examples on the results, we conduct comparative experiments with various examples. First, we randomly select 3, 6, and 9 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples. Additionally, despite the risk of data leakage, we select 6 examples from the Defects4J dataset. The results are presented in Table 15.
We further compute their similarity, shown as the bottom part of Table 14. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the -values are all below 0.05. Therefore, we can conclude that although different few-shot examples have perturbation, their performance is very similar.
4.1.4. Performance of Using the Same Number of Mutations
In Section 3, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of existing study (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by Bert (Degiovanni and Papadakis, 2022), which is 11397. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 16. Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
| Metric | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Real Bug Det. | 31.6% | 68.8% | 64.6% | 66.0% | 82.2% | 87.4% | 88.6% | 87.0% | 33.9% | 68.5% |
| Coupling Rate | 17.4% | 45.4% | 37.2% | 43.6% | 45.6% | 48.8% | 47.9% | 47.9% | 27.1% | 43.6% |
| Ochiai Coefficient | 12.6% | 22.2% | 14.1% | 19.0% | 26.5% | 36.0% | 31.8% | 34.1% | 15.4% | 16.9% |
| Comp. Rate | — | 96.9% | 54.1% | 26.1% | 61.2% | 75.5% | 74.3% | 74.6% | 12.4% | 69.5% |
| Dup Mut. Rate | — | 0.0% | 2.6% | 1.9% | 9.9% | 7.7% | 7.1% | 7.6% | 9.6% | 38.3% |
4.2. Implications
Based on our findings, we discuss the pros and cons of traditional and LLM-based mutation generation approaches, and suggest potential improvements.
4.2.1. Advantages of LLM-based Approaches
Our study demonstrates that LLMs can generate high-quality mutations for mutation testing at an acceptable cost. The key advantage of LLMs lies in their ability to produce mutations with stronger coupling relationships with real-world bugs, which is crucial for improving test effectiveness.
Additionally, LLMs could generate syntactically diverse mutations compared with traditional approaches, which is potentially the reason for generating high-coupling mutations. For example, Listing 1 presents two compilable mutations generated by GPT-3.5 for Chart-1, which are theoretically beyond the capabilities of traditional approaches. In Mutation-1, the LLM infers a relationship between variables (rightBlock implies the existence of leftBlock) and modifies the variable accordingly. This mutation is not killed by the original bug-triggering test, highlighting its potential for guiding test enhancement. Without natural language understanding (e.g., recognizing the correspondence between “right” and “left”), traditional approaches, both rule-based approaches and the approaches based on small-scale models, struggle to generate such context-aware mutations. In Mutation-2, the LLM transforms an else branch into an else if condition by synthesizing a logical condition using the integer variable seriesCount from the context, demonstrating an ability to introduce structurally meaningful changes that traditional approaches cannot achieve.
Additionally, our study explores how different prompt designs, model choices, and contextual information influence mutation generation performance. We find that selecting models with stronger code generation capabilities (e.g., GPT-4o) and designing prompts carefully are critical to achieving optimal results.
4.2.2. Advantages of Traditional Approaches
Despite the advantages of LLM-generated mutations, traditional approaches still exhibit strengths in reliability and efficiency. They produce fewer non-compilable and duplicate mutations, with PIT achieving a 100% compilability rate and a 0% duplicate mutation rate, while Major follows closely with a 97.6% compilability rate and 0% duplicate mutation rate. In contrast, LLM-based approaches exhibit significantly lower compilability and higher duplicate mutation rates. Additionally, overall, traditional rule-based approaches generate a slightly lower equivalent mutation rate, further avoiding redundant costs. This suggests that traditional approaches remain highly practical when strict compilability and low redundancy are prioritized.
4.2.3. Challenges and Future Directions
Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate substantial room for improvement. One major limitation is the high rate of non-compilable mutations, which limits their practical application in mutation testing. Our analysis reveals that common causes include syntactic errors in generated expressions, incorrect method invocations, and type mismatches. To address these challenges, future research should explore integrating LLM-based approaches with program analysis and repair techniques to refine mutation generation. Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutations before execution can enhance their reliability.
5. Threats to Validity
In this section, we illustrate the threats to the validity of our study.
The selected programming language, mutation generation approaches, datasets, and LLMs could be a threat to the validity of our results. To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek).
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014a) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024) which includes programs and faults that were made after the release time of the LLMs we use and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. On Defects4J, the results are GPT-3.5 (133), GPT-4o (63), GPT-4o-Mini (138), DeepSeek (66), StarChat (13), CodeLlama (28), LEAM (287), Bert (12), and Major (22), respectively. On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek (32), StarChat (35), CodeLlama (56), LEAM (56), Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except from GPTs and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with -value less than 0.05, indicating their performance is similar on the two datasets.
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompts, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar. One potential threat to validity is that the metrics used in this study may not fully reflect the actual bug detection in practice. Our evaluation is conducted on the fixed versions of programs, rather than the faulty versions where real-world testing would typically be applied, which may introduce potential differences. In other words, our study operates under the clean program assumption, which may not fully capture the dynamics of bug detection in practical scenarios. However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs,
The subjective nature of human decisions in labeling for equivalent mutations and non-compilation errors is another potential threat. To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutation. The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two annotators.
6. Related Work
In this section, we introduce the related work of our study.
6.1. Mutation Testing and Mutation Generation
Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978) one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014b; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
All these tasks require high-utility mutations. This is because it is required to compile and execute every mutation with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to couple with real faults/bugs to allow these techniques to be effective.
Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to rule-based. For example, one widely used mutation operator is to change one arithmetic binary operator with another one (e.g., ). During the last two decades, the majority of existing mutation testing approaches have been rule-based (Jia and Harman, 2010; Amalfitano et al., 2022). Researchers primarily focus on approaches in Java (e.g., PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), IBIR (Khanfir et al., 2023), etc.) and C/C++ (e.g., Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), WinMut (Wang et al., 2021), etc.), leveraging the mature ecosystems and tooling chains of these languages for mutation testing.
Recently, with the emergence of machine learning techniques, researchers proposed using deep learning models to generate mutations. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. However, these approaches suffer from the challenges in generating syntactic correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations. Perhaps the closest to our work is that of Degiovanni et al. who proposed Bert (Degiovanni and Papadakis, 2022). Bert generates mutation by masking the code elements to be mutated and then predicting the mutation via Bert (Devlin et al., 2018).
Several recent studies have also explored the use of LLMs for mutation generation. Tip et al. concurrently worked on using LLMs for mutation generation for JavaScript, but their approach neither included a comprehensive comparison with existing techniques nor did they utilize only three open-source LLMs (Tip et al., 2024). Similarly, Endres et al. employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches. Dakhel et al.(Dakhel et al., 2024) and Ibrahimzada et al.(Ibrahimzada et al., 2023) also proposed generating artificial bugs using LLMs. These approaches mainly evaluate from a certain perspective, such as mutation score or program repair. Compared with these approaches, our study is more comprehensive and evaluates existing mutation generation approaches from many aspects. Specifically, we measure the behavioral similarity between real bugs and mutations using established metrics, evaluate the syntactic diversity of mutations, and analyze the non-compilable mutations produced by LLMs. Our findings illustrate both the strengths and limitations of LLMs for mutation generation, and highlight directions for future improvement.
6.2. LLMs for Code
Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023). Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes. LLMs are driven by instructions in natural languages, i.e., prompts, which guide the models to perform target tasks and are crucial to performance. The design of prompts plays a crucial role in obtaining accurate and useful responses from these models.
Closed-source LLMs are proprietary models developed and maintained by companies or organizations, often offering advanced features and requiring a subscription or payment for access, while open-source LLMs are freely available models that allow users to view and modify them. For the models used in our study, GPT-3.5-Turbo, GPT-4o, and GPT-4o-Mini are closed-source LLMs released by OpenAI (Achiam et al., 2023), while DeepSeek-Coder, StarChat, and CodeLlama are open-source. An open-source LLM is usually fine-tuned against a base model for specified tasks. DeepSeek-Coder (Zhu et al., 2024) is based on DeepSeek. CodeLlama (Roziere et al., 2023) is derived from Llama (Touvron et al., 2023) by fine-tuning with extensive code datasets. StarChat is built on StarCoder (Li et al., 2023a) to enhance conversational abilities.
7. Conclusion
In this paper, we systematically investigate the performance of existing mutation generation appraoches across multiple dimensions, including their effectiveness, validity, and efficiency. Our study conducts extensive and detailed comparative experiments, covering both widely studied rule-based mutation generation approaches and LLM-based methods. Our evaluation results highlight that LLMs significantly outperform traditional approaches in generating diverse mutations that closely mimic real bugs. However, the results also expose key limitations of LLMs, particularly their tendency to generate a high proportion of non-compilable mutations. Therefore, we analyze the error types of the LLM-generated non-compilable mutations and identify member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations. Based on our findings, we emphasize the need for further research to enhance LLM-based mutation testing, addressing its current shortcomings while leveraging its strengths.

### v4

@@
 However, there has been no comprehensive empirical study of their performance for this increasingly important LLM-based Software Engineering application.
-To address this, we report the results of a comprehensive empirical study over six different LLMs, including both state-of-the-art open- and closed-source models, on 851 real bugs drawn from two different Java real-world bug benchmarks.
+To address this, we conduct a comprehensive empirical study evaluating BugFarm and LLMorpheus (the two state-of-the-art LLM-based approaches), alongside seven LLMs using our newly designed prompt, including both leading open- and closed-source models, on 851 real bugs from two Java real-world bug benchmarks.
 Our results reveal that, compared to existing rule-based approaches, LLMs generate more diverse mutants, that are behaviorally closer to real bugs and, most importantly, with higher fault detection.
-That is, 79.1% (for LLMs) vs.
-41.6% (for rule-based); an increase of percentage points.
-Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by , , and percentage points respectively.
+That is, 87.98% (for LLMs) vs.
+41.64% (for rule-based); an increase of percentage points.
+Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by , , and percentage points, respectively.
 These findings are immediately actionable for both research and practice.
@@
 Nevertheless, until the last five years, it had made little penetration into the industrial sector.
-In those most recent five years, it has become increasingly widely deployed in industry.
-For example, large tech sector companies such as Meta (Foster et al., 2025; Beller et al., 2021) and Google (Petrović et al., 2021) reported on results from deployment of mutation testing at their companies.
-In particular, results recently reported by Meta (Foster et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
+In the most recent five years, it has become increasingly widely deployed in industry.
+For example, large tech sector companies such as Meta (Harman et al., 2025; Beller et al., 2021) and Google (Petrović et al., 2021b; Petrović and Ivanković, 2018) reported on results from the deployment of mutation testing at their companies.
+In particular, results recently reported by Meta (Harman et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
 This recent trend towards LLM-based mutant generation traces back to work on earlier learning-based approaches, which leverage deep learning models to generate mutations by learning patterns from historical bug fixes (Beller et al., 2021; Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., 2025).
 These methods aim to introduce mutations that better resemble real faults, potentially increasing their relevance for testing.
-However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, there is to date, no large-scale comprehensive empirical study.
+However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, to date, no large-scale comprehensive empirical study.
 Without such a study, we have no basis on which to deploy an industrial scale with confidence.
 Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutation generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
-The generation of mutants, using any technique is challenging because of the following three important issues:
-Generating effective mutations: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019).
+The generation of mutants, using any technique, is challenging because of the following three important issues:
+Generating effective mutations: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019; Petrović et al., 2021a).
 While existing methods have shown promise, there is still significant room for improvement in enhancing their fault-revealing potential (Degiovanni and Papadakis, 2022).
@@
 At the same time, mutations need to be diverse and cover the entire spectrum of code and program behaviors.
-This paper reports on how well Large-Language Models (LLMs) handle these challenges, compared to more traditional rule-based approaches.
-We conducted a comprehensive comparison between six LLMs (three closed-source commercial and three open-source ones) and four existing mutation testing techniques with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014a) and 246 bugs from ConDefects (Wu et al., 2024)).
-Overall, our experiments required the generation of 611,100 mutations.
+This paper examines how well Large-Language Models (LLMs) address these challenges, in comparison to more traditional rule-based approaches.
+We conducted a comprehensive empirical study which studies BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025) (i.e., the state-of-the-art LLM-based mutation generation approaches), alongside seven LLMs adopting our newly designed prompt, covering both leading open- and closed-source models, and four existing mutation testing techniques (i.e., Major (Just et al., 2011), PIT (Coles et al., 2016), LEAM (Tian et al., 2022), and BERT (Degiovanni and Papadakis, 2022)) with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014b) and 246 bugs from ConDefects (Wu et al., 2024)).
+Overall, our primary experiments required the generation of 666,148 mutations.
 Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutation rate, equivalent mutation rate, and efficiency).
-Additionally, we study how different prompt engineering strategies and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat--16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task.
+Additionally, we study how different prompt engineering strategies, turbulence on prompt, and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), DeepSeek-V3 (Liu et al., 2024), StarChat--16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task.
+We also explore the potential of using LLMs for generating high-order mutations.
 Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
 Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs.
-This leads to a higher fault detection rate than traditional approaches.
-For example, the mutants generated by GPT-4o-Mini lead to a fault detection rate of 90.8%, while the mutants generated by LEAM, PIT, and Major are 63.9%, 40.1%, and 66.9%, respectively.
+This results in a higher fault detection rate compared to traditional approaches.
+For instance, the mutants generated by DeepSeek-V3-761b and GPT-4o-Mini achieve fault detection rates of 91.1 % and 90.8%, respectively, whereas those generated by LEAM, PIT, and Major yield substantially lower rates of 63.9%, 40.1%, and 66.9%, respectively.
 Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutation rate, and equivalent mutation rate.
@@
 Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types).
-In particular, our results show that GPT-4o-Mini and DeepSeek exhibit the greatest diversity, newly introducing 49 different AST node types.
+In particular, our results show that the DeepSeek models exhibit the greatest diversity, newly introducing 49 different AST node types.
 In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 21.
+The existing LLM-based approach, BugFarm, introduces 31 node types, higher than the traditional approaches as well.
 The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations.
-Moreover, the distributions of AST edit distance of mutations compared to the original code also present that LLMs generate more balanced and expressive mutations.
-For instance, almost 70% of the mutations of Major have the AST edit distance of one, while for GPT-4o, only 50% of its mutations have the same distance.
-We also categorise the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall into 9 compilation error types with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones.
+Moreover, the distributions of AST edit distance of mutations compared to the original code also show that LLMs generate more balanced and expressive mutations.
+For instance, almost 70% of the mutations of Major have an AST edit distance of one, whereas for GPT-4o, only 50% of its mutations have the same distance.
+We also categorise the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall into 9 compilation error types, with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones.
 Moreover, we further analyze the code context of the non-compilable mutations generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features.
-This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, which could enhance the effectiveness of mutation testing using AI models.
+This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, thereby enhancing the effectiveness of mutation testing using AI models.
+We analyze the alive (i.e., survive) mutations of each approach, which are important to strengthen the test case.
+For one alive non-equivalent mutation, it can be (1) not covered, (2) covered but no state infection, or (3) state infection without detection.
+We analyze the distribution of alive mutations across different approaches and find that LLMs based on our prompt are more likely to generate uncovered mutations, which consequently remain alive, providing more valuable guidance for enhancing unit tests.
 Our implementation and the experimental data are available at our repo: https://github.com/da-123-snake/Mutantion-Generation-Study.
@@
 We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods.
+Our study not only covers other existing LLM-based mutation generation approaches, but also introduces a new effective prompt template.
 Our findings indicate that LLMs excel in generating diverse mutations that closely mimic real bugs.
 We analyze the error types of non-compilable mutations and determine that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
-We build a high-quality dataset of Java mutations comprehensively labeled with detailed quality and behavior metrics.
+We construct a high-quality dataset of Java mutations, comprehensively annotated with detailed quality and behavior metrics, which can also serve as a valuable resource for fine-tuning in future research.
 The remainder of this paper is organized as follows.
@@
 RQ2.3 (Equivalence) What percentage of generated mutations are equivalent across different approaches?
-RQ3 (Efficiency) What are the time costs of different mutation generation approaches?
-This research question investigates the time costs of different mutation generation approaches.
-Time costs are associated with the efficiency and practicality of mutation generation approaches.
+RQ3 (Efficiency) What are the general costs of different mutation generation approaches?
+This research question investigates the general costs of different mutation generation approaches.
+We consider time and token costs in this RQ, which are associated with the efficiency and practicality of mutation generation approaches.
 RQ4 (Compilation Error) What are the types of compilation errors of the non-compilable mutations of different approaches?
@@
 Understanding these errors can guide mutation generation approaches to enhance their practical usability.
+RQ5 (Alive Mutations) What are the distributions of alive mutations across different approaches?
+This RQ aims to investigate the underlying reasons why mutations remain alive, by categorizing them into three cases following the existing study (Just et al., 2014a): (1) uncovered by tests, (2) covered but without infecting the program states, and (3) injecting the states but remaining undetected by tests.
+By examining how these cases are distributed across different approaches, we can better understand the effectiveness of generated mutations and their potential value for guiding test suite improvement.
 2.2. Studied Mutation Generation Approaches
-Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking111https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutation generation approaches is for Java.
+Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking111https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutation generation approaches are for Java.
 We categorize these approaches into two groups: traditional approaches (i.e., those developed before the emergence of LLMs) and LLM-based approaches (i.e., approaches that leverage LLMs and are designed for interaction in human-like conversations), as shown below.
+The details of the settings for each generation approach are illustrated in Section 2.5.
 Table 1 summarizes the mutation generation approaches covered by this study.
@@
 | BERT | Mutation generation with the pre-trained model CodeBERT. | |
-| LLM-based | GPT-3.5 | GPT-3.5-Turbo, whose based model is GPT, trained on data before 2021/09, released in 2021/03. |
-| GPT-4o | The based model is GPT, trained on data before 2023/10, released in 2024/05. | |
-| GPT-4o-M | GPT-4o-Mini, whose based model is GPT, trained on data before 2023/10, released in 2024/07. | |
-| SC-16b | StarChat--16b, whose based model is StarCoder, released in 2023/06. | |
-| CL-13b | CodeLlama-Instruct-13b, whose based model is Llama, released in 2023/08. | |
-| DS-236b | DeepSeek-Coder-V2-236b, whose based model is DeepSeek, released in 2024/07. |
+| LLM-based | BF (GPT-4o) | The existing LLM-based approach BugFarm with GPT-4o. |
+| LLMph (GPT-4o) | The existing LLM-based mutation generation approach LLMorpheus with GPT-4o. | |
+| LLMut (GPT-3.5) | Our prompt with GPT-3.5-Turbo, trained on data before 2021/09, released in 2021/03. | |
+| LLMut (GPT-4o) | Our prompt with GPT-4o, trained on data before 2023/10, released in 2024/05. | |
+| LLMut (GPT-4o-M) | Our prompt with GPT-4o-Mini, trained on data before 2023/10, released in 2024/07. | |
+| LLMut (SC-16b) | Our prompt with StarChat--16b, whose base model is StarCoder, released in 2023/06. | |
+| LLMut (CL-13b) | Our prompt with CodeLlama-Instruct-13b, whose base model is Llama, released in 2023/08. | |
+| LLMut (DS-236b) | Our prompt with DeepSeek-Coder-V2-236b, whose base model is DeepSeek, released in 2024/07. | |
+| LLMut (DS-671b) | Our approach with DeepSeek-V3-761b, whose base model is DeepSeek, released in 2024/12. |
 2.2.1. Traditional Approaches
@@
 Major: Major (Just et al., 2011) is another high-impact mutation testing framework, that supports nine operators. In our study, we use Major 1.3.5 and involve all mutation operators.
-LEAM: LEAM (Tian et al., 2022) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutations.
+LEAM: LEAM (Tian et al., 2022, [n.
+d.]) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutations.
 Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutations.
@@
 2.2.2. LLM-based Approaches
-Our study involve multiple kinds of LLMs, including both open-source and closed-source models.
-LLMs: Our study also comprehensively covers the most widely studied LLMs.
-Specifically, we select a representative set of LLMs with code capabilities both from commercial closed-source and open-source models.
-Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5 (GPT-3.5-Turbo) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-M (GPT-4o-Mini) (Achiam et al., 2023)), and open-source LLMs, including SC-16b (StarChat--16b), CL-13b (CodeLlama-Instruct-13b), and DS-236b (DeepSeek-Coder-V2-236b).
+Our study incorporates two existing LLM-based mutation generation approaches (i.e., BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025)), and introduces a new prompt design for mutation generation, which will be presented later.
+Existing LLM-based Approaches: We adopt two the state-of-the-art LLM-based mutation generation approaches, BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025).
+BugFarm trains a small-scale model for selecting lines for mutation, and then utilizes LLMs to mutate the target line.
+LLMorpheus performs mutation given the whole method as context.
+Note that LLMorpheus was originally designed for JavaScript.
+Upon examining its prompt, we found it to be language-agnostic.
+Therefore, we reimplemented the code-extraction component for Java (such as parsing ASTs and extracting the code context for mutation) while retaining the original prompt.
+For both approaches, we employ GPT-4o (Achiam et al., 2023), the current most widely used model, as their backend LLM.
+LLMs with Our Prompt: Different from the existing LLM-based approaches, we propose a new prompt engineering for mutation generation, shown as Figure 1.
+For ease of presentation, in the following part of this paper, we call our prompt LLMut.
+To craft effective prompts, we follow the best practices, which suggest that prompts should comprise four aspects: Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024).
+In the Instructions, we direct the LLMs to generate mutants for the target code element.
+The placeholder MUT_NUMBER specifies the number of mutations required, while the placeholder CODE_ELEMENT specifies the target code snippet to be mutated.
+In the Context, we clearly state that a mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark.
+In the Input Data, we specify the code element to be mutated and the number of mutants to generate.
+In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments.
+In addition, our prompts need few-shot examples, which should originate from real bugs, enabling LLMs to learn how to mutate code from real bugs.
+We further explore the different prompt settings in Section 4.1 and the impact of prompt turbulence in Section 4.2.
+To avoid the few-shot examples leaking information, beyond the evaluation datasets (i.e., as shown later, they are Defects4J and ConDefects), we employ another benchmark, QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b).
+Referring to the study by Deng et al.
+(Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples.
+Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt.
+To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected.
+If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 2.
+As LLM capabilities continue to advance, the optimal number of few-shot examples may vary.
+We therefore investigate the impact of different numbers of few-shot examples in Section 4.1.3.
+| Correct Version | Buggy Version |
+| n = (n & (n - 1)); | n = (n ˆ (n - 1)); |
+| while (!queue.isEmpty()) | while (true) |
+| return depth==0; | return true; |
+ArrayList r = new ArrayList();
+r.add(first).addll(subset); to_add(r); |
+to_add.addAll(subset); |
+| c = bin_op.apply(b,a); | c = bin_op.apply(a,b); |
+| while(Math.abs(x-approx*approx)>epsilon) | while(Math.abs(x-approx)>epsilon) |
+To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016).
+We use PIT-generated mutations with the universal set of operators and find that PIT generates 1.15 mutations per line of code.
+Therefore, based on the given code context, we generate one mutation per line.
+We explore the influence of context length on mutation generation in Section 4.
+Additionally, for the temperature, we use the default value for each LLM.
+Compared with the prompts used in BugFarm222https://github.com/Intelligent-CAT-Lab/BugFarm and LLMorpheus333https://github.com/githubnext/llmorpheus, our prompt design offers several advantages:
+it incorporates real-world bugs as few-shot examples, which have been shown to significantly enhance LLM performance in many code-related tasks;
+it determines the number of generated mutants based on the size of the target code, rather than on a fixed number;
+it specifies the mutation location using the target statement itself rather than line numbers, enabling more precise localization (Fan et al., 2023a); and
+it employs JSON output formatting, which is widely used in LLM fine-tuning and downstream tasks.
+Based on LLMut, our study comprehensively covers the most widely studied LLMs.
+Specifically, we select a representative set of LLMs with code capabilities, both from commercial closed-source and open-source models.
+Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5 (GPT-3.5-Turbo) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-M (GPT-4o-Mini) (Achiam et al., 2023)), and open-source LLMs, including SC-16b (StarChat--16b), CL-13b (CodeLlama-Instruct-13b), DS-236b (DeepSeek-Coder-V2-236b), and DS-671b (DeepSeek-V3-671b).
 The details of these models are shown as Table 1.
 Note that OpenAI has not officially released the parameter size of their closed-source models.
-The parameter size of the open-sourced models we studied ranges from 13 billion to 236 billion.
+The parameter size of the open-sourced models we studied ranges from 13 billion to 671 billion.
+Moreover, the selected LLMs represent a balance between general-purpose models (e.g., GPT-4o and DeepSeek-V3-671b), which are trained for a wide range of natural language and reasoning tasks, and code-specialized models (e.g., CodeLlama-Instruct-13b and DeepSeek-V2-Coder-236b), which are further fine-tuned on large-scale code corpora.
+This selection allows us to evaluate mutation generation across both broadly capable models and those explicitly optimized for programming tasks, ensuring the conclusions are representative of the current landscape of LLMs.
 2.3. Datasets
@@
 Every bug is accompanied by at least one bug-triggering test because we need to measure whether the mutations affect the execution of bug-triggering tests.
-To this end, we employ the Defects4J 2.0 (Just et al., 2014a) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 2.
+To this end, we employ the Defects4J 2.0 (Just et al., 2014b) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 3.
 In total, we conducted experiments on 851 bugs.
-Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023; Just et al., 2014b; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains history bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs.
-We select the most widely used 12 projects from the Defects4J 2.0, which contains 605 bugs in total.
-However, from Table 1 and Table 2, we observe the periods of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage.
+Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023; Just et al., 2014c; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains historical bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs.
+We select the most widely used 12 projects from Defects4J 2.0, which contain 605 bugs in total.
+However, from Table 1 and Table 3, we observe that the periods of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage.
 Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns.
-ConDefects consists of tasks from AtCoder222https://atcoder.jp programming contest.
+ConDefects consists of tasks from AtCoder444https://atcoder.jp programming contest.
 To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs.
+Among them, only the most recent model, DeepSeek-V3-731b, cannot guarantee data leakage-free, since it is released after the ConDefects dataset.
 2.4. Evaluation Metrics
 In this subsection, we present the evaluation metrics used in our study to assess the quality of the generated mutations comprehensively.
-We evaluate the mutations from multiple perspectives as before mentioned.
+We evaluate the mutations from multiple perspectives, as mentioned before.
 Note that only relying on the mutation score is insufficient for a meaningful assessment.
@@
 Let all generated mutations be the set , where LLMs may generate syntactically incorrect mutations that can not pass compilation.
-Thus we denote the compilable mutations as the set .
+Thus, we denote the compilable mutations as the set .
 Note that the set contains two types of mutations that should be removed, the duplicate mutations (denoted as ) and equivalent mutations (denoted as ).
@@
 The duplicate mutations refer to those identical to the original code or other mutations.
-In mutation testing, the ideal set of mutations to employ is from the set , excluding equivalent mutations that have no contributions to discovering the weakness of tests.
+In mutation testing, the ideal set of mutations to employ is from the set , excluding equivalent mutations that have no contribution to discovering the weakness of tests.
 However, determining the equivalent mutations set is undecidable, so we follow the common practice and use the set as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025).
+Therefore, let denote the set of killed mutants among the compilable and non-duplicate ones (i.e., ).
+The mutation score is then computed as follows:
+Note that traditional mutation generation approaches, such as Major and PIT, produce very few duplicate mutations (i.e., is 0).
+Hence, early studies (Kintis et al., 2017) defined equivalent mutations as those semantically equivalent to the original code, which may implicitly include , which is an empty set.
+However, LLM-based approaches tend to generate a large number of non-mutated duplicates.
+Moreover, while duplicate mutations can be easily filtered, identifying equivalent mutations remains challenging (Tian et al., 2024).
+Therefore, in this paper, we refine the definition of equivalent mutations to be those that are syntactically different but semantically equivalent to the original code.
 2.4.1. RQ1: Effectiveness
@@
 These metrics assess whether bug-triggering tests can detect LLM-generated mutations (i.e., RQ1.1), how closely mutations resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutation generation techniques (i.e., RQ1.3).
-The metrics of these sub-RQs are listed as bellow.
+The metrics of these sub-RQs are listed below.
 RQ1.1: Real Bug Detection Rate: Following the approach of Just et al.
-(Just et al., 2014b), we evaluate the effectiveness of mutation generation by measuring its ability to detect real bugs using developer-written tests.
+(Just et al., 2014c), we evaluate the effectiveness of mutation generation by measuring its ability to detect real bugs using developer-written tests.
 For a given real bug (which may span multiple lines), we apply a mutation generation approach to create a set of mutations on the fixed version of the program.
@@
 (Ojdanic et al., 2023).
-More concretely, let and be a mutation and its corresponding real bug, and let and be the set of their failing test cases respectively. The coefficient of is computed by the following formula:
+More concretely, let and be a mutation and its corresponding real bug, and let and be the set of their failing test cases, respectively. The coefficient of is computed by the following formula:
 Therefore, if the Ochiai coefficient of is , it is semantically equal to its corresponding bug .
@@
 The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutations (i.e., ).
-Note that Real Bug Detection Rate estimates the effectiveness via the proportion of detected bugs, which Coupling Rate estimates it via the proportion of mutations.
+Note that Real Bug Detection Rate and Coupling Rate measure effectiveness from different perspectives.
+Real Bug Detection Rate evaluates effectiveness at the bug level, i.e., the proportion of bugs that can be revealed by at least one mutant.
+In contrast, Coupling Rate evaluates effectiveness at the mutation level, i.e., the proportion of mutants whose failing tests overlap with bug-triggering tests.
+For instance, a mutation approach may generate a large number of high-quality mutants for only a few specific bugs, but few mutants for others.
+This can result in a high Coupling Rate (since the generated mutants are strongly aligned with the bug-triggering tests of those few bugs) but a low Real Bug Detection Rate (since many other bugs remain undetected).
+Conversely, another approach may generate mutants that cover a wide range of bugs, thereby achieving a high Real Bug Detection Rate.
+However, if the mutants for each bug exhibit only weak coupling with the bug-triggering tests, the overall Coupling Rate will remain low.
+Therefore, these two metrics are complementary, with each reflecting only one aspect of mutation effectiveness.
 RQ1.2: Coupling Categorization: We categorize each generated mutation into distinct coupling scales based on their similarity to real bugs, following the classification proposed by (Gay and Salahirad, 2023), and analyze their distribution across different approaches.
@@
 Therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes.
+In other words, we identify the unique AST node types introduced by each approach’s generated mutants, relative to the set of node types present in the original program.
 The approach with a greater variety of AST node types is considered superior, as it indicates a broader exploration of the syntactic space and enhances the likelihood of detecting weaknesses in the test suite.
@@
 A higher edit distance suggests more significant syntactic changes.
+The AST Edit Distance metric is well-suited for structured code, as it captures syntactic differences at the tree level rather than merely comparing token sequences, as in BLEU and similar metrics.
+However, it still has inherent limitations, such as its inability to fully reflect the impact of mutations on program behavior.
+We further discuss these limitations later in Section 5.
 2.4.2. RQ2: Validity
 We consider three key metrics to evaluate the validity of different mutation generation approaches: generation efficiency, compilability, and the proportion of duplicate and equivalent mutations, providing a basic understanding of the validity of each approach.
-RQ2.1: Compilability Rate: the proportion of mutations that successfully compile, is computed as:
-Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher comparability rate.
+RQ2.1: Compilability Rate: the proportion of mutations that successfully compile, which is computed as:
+Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher compilability rate.
 RQ2.2: Duplicate Mutation Rate: the proportion of duplicate mutations, computed as:
-LLMs sometimes generate duplicate mutations, i.e., mutations that are exactly the same as the original code.
+LLMs sometimes generate duplicate mutations, i.e., mutations that are either identical to the original code or identical to each other.
 While traditional approaches rarely generate duplicate mutations, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data.
+After eliminating duplicate mutations, all remaining mutations are guaranteed to be syntactically distinct both from the original code and from each other.
 RQ2.3: Equivalent Mutation Rate: the proportion of equivalent mutation, computed as:
-Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. We prefer the approach with a lower EMR.
+Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score.
+We prefer the approach with a lower EMR.
+Note that while equivalent mutations are also considered semantically duplicate, they are not easily identifiable.
+Therefore, we distinguish them from other duplicate mutations, which can be filtered syntactically.
 2.4.3. RQ3: Efficiency
@@
 To assess this, we compare these approaches based on time costs.
-RQ3: Average Generation Time: measures the average time required to generate each mutation, providing a quantitative evaluation of the efficiency of each approach.
-2.5. Mutation Generation
+RQ3.1: Average Generation Time: measures the average time required to generate each mutation, providing a quantitative evaluation of the efficiency of each approach.
+RQ3.2: Average Token: measures the average number of consumed tokens per mutation, reflecting both the financial cost (due to token-based pricing) and the computational time cost (since longer token sequences typically require more processing time).
+Since input and output token costs differ across models, we calculate them separately.
+2.5. Mutation Generation Setups
 In this subsection, we discuss the setups for mutation generation approaches, especially the setups for LLM-based generation.
-We performed our experimental analysis on four rented cloud servers, each equipped with 2 NVIDIA GeForce RTX 3090 Ti, 168GB memory, and a 64 Core CPU Intel(R) Xeon(R) Platinum 8350C.
-2.5.1. Mutation Generation via Traditional Approaches
 For the well-established mutation generation approaches (i.e., PIT, Major, LEAM, and BERT), we activate all their mutation operators.
 Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutations within the context of real-world bugs.
-Note that for all mutation generation approaches, we employ the same setting for the code context.
-2.5.2. Mutation Generation via LLMs
-Recently, LLMs have shown great potential in numerous code-related tasks, to better understand their ability in mutation testing, we also study LLMs for generating mutations.
-LLMs operate based on prompts, which consist of natural language instructions, often accompanied by code snippets.
-LLMs are driven by prompts, which are natural language instructions maybe with code.
-To effectively leverage LLMs for mutation generation, we design a well-crafted prompt template, shown as Figure 1.
-We follow the best practices suggesting prompts should comprise four aspects, namely, Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024).
-In the Instructions, we direct the LLMs to generate mutants for the target code element.
-In the Context, we clearly state that mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark.
-In the Input Data, we specify the code element to be mutated and the number of mutants to generate.
-In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments.
-In addition, our prompts need few-shot examples, which should originate from real bugs enabling LLMs to learn how to mutate code from real bugs.
-This allows the model to understand the context and logic behind the changes, improving its ability to generate relevant and effective mutations.
-To avoid the few-shot examples leaking information, beyond Defects4J and ConDefects, we employ another benchmark QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b).
-Referring to the study by Deng et al.
-(Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples.
-Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt.
-To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected.
-If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 3.
-To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016).
-We use PIT-generated mutations with the universal set of operators and find that PIT generates 1.15 mutations per line of code.
-Therefore, based on the given code context, we generate one mutation per line.
-We explore the influence of context length on mutation generation in Section 4.
-Additionally, for the temperature, we use the default value for each LLM.
-| Correct Version | Buggy Version |
-| n = (n & (n - 1)); | n = (n ^ (n - 1)); |
-| while (!queue.isEmpty()) | while (true) |
-| return depth==0; | return true; |
-ArrayList r = new ArrayList();
-r.add(first).addll(subset); to_add(r); |
-to_add.addAll(subset); |
-| c = bin_op.apply(b,a); | c = bin_op.apply(a,b); |
-| while(Math.abs(x-approx*approx)>epsilon) | while(Math.abs(x-approx)>epsilon) |
+Note that we supply the same code snippets as context across all mutation generation approaches to ensure a fair comparison.
+For all LLM-based approaches, we use the default settings of each model (e.g., token limits and temperature).
+For models accessible via Web API (e.g., GPT series and DeepSeek models), we rely on their official APIs.
+For the remaining open-source LLMs, we perform inference on four rented cloud servers, each equipped with dual NVIDIA GeForce RTX 3090 Ti GPUs, 168 GB of memory, and a 64-core Intel(R) Xeon(R) Platinum 8350C processor.
 3. Evaluation Results
@@
 A large number of mutations may include redundant or equivalent mutants that do not contribute to mutation scores, while a high mutation score may be influenced by the ease with which mutants are detected rather than their resemblance to real faults.
-| Metric | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
-| Mut. # | 340728 | 20831 | 36505 | 12946 | 31656 | 31945 | 34417 | 34741 | 28058 | 27614 |
-| Mut. Score | 0.541 | 0.521 | 0.648 | 0.677 | 0.702 | 0.737 | 0.708 | 0.714 | 0.507 | 0.669 |
-| Real Bug Detec. | 40.1% | 66.9% | 63.9% | 60.5% | 87.0% | 89.9% | 90.8% | 89.4% | 38.5% | 71.2% |
-| Avg. Ochiai | 23.0% | 35.5% | 32.4% | 29.3% | 47.7% | 56.2% | 54.1% | 50.9% | 26.0% | 33.3% |
-| Coupling Rate | 23.5% | 38.9% | 38.1% | 47.3% | 46.9% | 50.6% | 49.5% | 49.5% | 34.0% | 45.6% |
-| Comp. Rate | 100% | 97.6% | 44.5% | 26.5% | 62.5% | 76.4% | 75.3% | 75.8% | 16.4% | 69.3% |
-| Dup. Mut. Rate | 0.0% | 0.0% | 2.3% | 1.7% | 10.3% | 7.8% | 7.0% | 7.7% | 12.8% | 37.1% |
-| A. G. Time (s) | 0.02 | 0.08 | 3.06 | 2.34 | 1.79 | 1.31 | 1.65 | 4.25 | 7.53 | 9.06 |
-Note: The mutation generation approach achieving the best performance in terms of a metric is highlighted in grey background color.
-GPT-4o-M: GPT-4o-Mini.
-DS-236b: DeepSeek-Coder-V2-236b.
-SC-16b: StarChat--16b.
-CL-13b: CodeLlama-Instruct-13b.
+| Name | #Mut. | Mut. S. | R. B. Detec. | Avg. Ochiai | Coup. Rate | Comp. Rate | Dup. Rate | Time |
+| PIT | 340728 | 0.541 | 40.1% | 23.0% | 23.5% | 100% | 0.0% | 0.02 |
+| Major | 20831 | 0.521 | 66.9% | 35.5% | 38.9% | 97% | 0.0% | 0.08 |
+| LEAM | 36505 | 0.648 | 63.9% | 32.4% | 38.1% | 44.5% | 2.3% | 3.06 |
+| BERT | 12946 | 0.677 | 60.5% | 29.3% | 47.3% | 26.5% | 1.7% | 2.34 |
+| BF (GPT-4o) | 13068 | 0.762 | 58.6% | 17.7% | 47.9% | 74.1% | 7.5% | 5.35 |
+| LLMph (GPT-4o) | 20043 | 0.754 | 68.5% | 24.1% | 51.7% | 67.8% | 7.1% | 3.41 |
+| LLMut (GPT-3.5) | 31656 | 0.702 | 87.0% | 47.7% | 46.9% | 62.5% | 10.3% | 1.79 |
+| LLMut (GPT-4o) | 31945 | 0.737 | 89.9% | 56.2% | 50.6% | 76.4% | 7.8% | 1.31 |
+| LLMut (GPT-4o-M) | 34417 | 0.708 | 90.8% | 54.1% | 49.5% | 75.3% | 7.0% | 1.65 |
+| LLMut (DS-761b) | 33596 | 0.749 | 91.1% | 59.8% | 50.8% | 77.6% | 7.5% | 2.57 |
+| LLMut (DS-236b) | 34741 | 0.714 | 89.4% | 50.9% | 49.5% | 75.8% | 7.7% | 4.25 |
+| LLMut (SC-16b) | 28058 | 0.507 | 38.5% | 21.7% | 34.0% | 16.4% | 12.8% | 7.53 |
+| LLMut (CL-13b) | 27614 | 0.669 | 71.2% | 33.3% | 45.6% | 69.3% | 37.1% | 9.06 |
+Note: The mutation generation approach achieving the best performance in terms of a metric is highlighted in a grey background color.
+BF: BugFarm. LLMph: LLMorpheus. LLMut: Our new prompt.
+GPT-4o-M: GPT-4o-Mini. DS-761b: DeepSeek-V3-761b. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat--16b. CL-13b: CodeLlama-Instruct-13b.
 3.1. RQ1: Performance on Effectiveness
@@
 As illustrated in Section 2.4.1, we estimate real bug detectability via the following metric.
-Real Bug Real Bug Detection Rate: The row Real Bug Detec.
-of Table 4 presents the proportion of real bugs that can be detected across different mutation generation approaches.
+Real Bug Detection Rate: The column Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutation generation approaches.
 Among the traditional approaches, PIT achieves a detection rate of 40.1%, while Major, LEAM, and BERT perform significantly better with a detection rate of 66.9%, 63.9%, and 60.5%, respectively.
 For LLM-based approaches, the results are significantly more effective than those of traditional ones.
-GPT-4o-Mini achieves the highest detection rate at 90.8%, closely followed by GPT-4o and DeepSeek, with detection rates of 89.9% and 89.4%, respectively.
+The existing LLM-based approaches, BugFarm and LLMorpheus, achieve a rate of 58.6% and 68.5%, respectively.
+For the LLMs based on our prompting, DeepSeek-V3-761b achieves the highest detection rate at 91.1%, closely followed by GPT-4o-Mini, GPT-4o and DeepSeek-Coder-V2-236b, with detection rates of 90.8%, 89.9%, and 89.4%, respectively.
 Meanwhile, GPT-3.5 and CodeLlama achieve moderate detection rates of 87.0% and 71.2%, respectively.
-| Project | Bug # | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
-| Chart | 26 | 19 | 17 | 16 | 10 | 22 | 22 | 23 | 23 | 9 | 8 |
-| Lang | 65 | 23 | 51 | 27 | 38 | 52 | 55 | 56 | 57 | 11 | 41 |
-| Time | 27 | 12 | 23 | 18 | 15 | 25 | 24 | 26 | 22 | 10 | 17 |
-| Math | 106 | 40 | 82 | 44 | 65 | 92 | 94 | 96 | 95 | 58 | 77 |
-| Mockito | 38 | 14 | 28 | 27 | 5 | 30 | 29 | 36 | 34 | 16 | 21 |
-| Closure | 133 | 38 | 92 | 122 | 103 | 130 | 131 | 131 | 132 | 79 | 114 |
-| Cli | 39 | 6 | 11 | 27 | 29 | 37 | 39 | 39 | 39 | 2 | 34 |
-| Codec | 18 | 6 | 11 | 12 | 11 | 14 | 16 | 17 | 17 | 1 | 11 |
-| Csv | 16 | 7 | 13 | 13 | 7 | 15 | 16 | 16 | 16 | 2 | 13 |
-| Gson | 18 | 6 | 9 | 11 | 12 | 13 | 16 | 16 | 15 | 1 | 10 |
-| JacksonCore | 26 | 1 | 17 | 17 | 18 | 23 | 23 | 21 | 21 | 1 | 17 |
-| Jsoup | 93 | 48 | 77 | 90 | 86 | 93 | 92 | 93 | 90 | 57 | 87 |
-| ConDefects | 246 | 121 | 138 | 120 | 116 | 194 | 208 | 203 | 200 | 81 | 156 |
-| ALL | 851 | 341 | 569 | 544 | 515 | 740 | 765 | 773 | 761 | 328 | 606 |
+| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Gs. | Csv | Ja. | Js. | ConD. | All |
+| #Bug | 26 | 65 | 27 | 106 | 38 | 133 | 39 | 18 | 16 | 18 | 26 | 93 | 246 | 851 |
+| PIT | 19 | 23 | 12 | 40 | 14 | 38 | 6 | 6 | 7 | 6 | 1 | 48 | 121 | 341 |
+| Major | 17 | 51 | 23 | 82 | 28 | 92 | 11 | 11 | 13 | 9 | 17 | 77 | 138 | 569 |
+| LEAM | 16 | 27 | 18 | 44 | 27 | 122 | 27 | 12 | 13 | 11 | 17 | 90 | 120 | 544 |
+| BERT | 10 | 38 | 15 | 65 | 5 | 103 | 29 | 11 | 7 | 12 | 18 | 86 | 116 | 515 |
+| BF (GPT-4o) | 14 | 37 | 15 | 57 | 28 | 93 | 23 | 6 | 8 | 11 | 7 | 83 | 117 | 499 |
+| LLMph (GPT-4o) | 17 | 51 | 20 | 61 | 31 | 121 | 30 | 12 | 8 | 12 | 9 | 87 | 124 | 583 |
+| LLMut (GPT-3.5) | 22 | 52 | 25 | 92 | 30 | 130 | 37 | 14 | 15 | 13 | 23 | 93 | 194 | 740 |
+| LLMut (GPT-4o) | 22 | 55 | 24 | 94 | 29 | 131 | 39 | 16 | 16 | 16 | 23 | 92 | 208 | 765 |
+| LLMut (GPT-4o-M) | 23 | 56 | 26 | 96 | 36 | 131 | 39 | 17 | 16 | 16 | 21 | 93 | 203 | 773 |
+| LLMut (DS-761b) | 24 | 61 | 27 | 101 | 35 | 131 | 39 | 13 | 16 | 16 | 25 | 92 | 195 | 775 |
+| LLMut (DS-236b) | 23 | 57 | 22 | 95 | 34 | 132 | 39 | 17 | 16 | 15 | 21 | 90 | 200 | 761 |
+| LLMut (SC-16b) | 9 | 11 | 10 | 58 | 16 | 79 | 2 | 1 | 2 | 1 | 1 | 57 | 81 | 328 |
+| LLMut (CL-13b) | 8 | 41 | 17 | 77 | 21 | 114 | 34 | 11 | 13 | 10 | 17 | 87 | 156 | 606 |
 Note: The mutation generation approaches that achieve the best performance in one project are highlighted in grey background color.
+BF: BugFarm. LLMph: LLMorpheus. LLMut: our prompt. Moc.: Mockito. Clo.: Closure. Co.: Codec. Gs.: Gson. Ja.: JacksonCore. Js.: Jsoup. ConD: ConDefects.
 To better understand the detection of the generation approaches, we analyze their project-wise detectability, shown as Table 5.
 For individual projects, LLM-based approaches consistently achieve top performance across multiple projects.
-Particularly, GPT-4o-Mini outperforms other approaches in several projects, including Chart, Time, Mockito, and Jsoup, where it detects the highest number of bugs.
-Similarly, DeepSeek and GPT-4o demonstrate strong performance, achieving the best results in projects such as Lang, Math, and Closure.
+Notably, DeepSeek-V3-761b detects the largest number of bugs in most Defects4J projects and achieves the highest overall bug detection count.
+Similarly, DeepSeek-Coder-V2-236b, GPT-4o, and GPT-4o-Mini demonstrate strong performance, achieving the best results in projects such as Mockio, Csv, and ConDefects.
 3.1.2. RQ1.2 Coupling
 To assess the degree of coupling, we perform comprehensive experiments that include both an overall metric evaluation and a project-wise analysis.
-Average Ochiai Coefficient: The row Avg.
+Average Ochiai Coefficient: The column Avg.
 Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measures the semantic distance between mutations and real bugs.
@@
 LEAM and BERT achieve coefficients of 32.4% and 29.3%, respectively, indicating that the best-performing rule-based approach, Major, outperforms both of them.
-The LLM-based approaches demonstrate significantly higher Ochiai coefficients, with GPT-4o achieving the highest value of 56.2%, followed closely by GPT-4o-Mini at 54.1%, and DeepSeek at 50.9%.
-Overall, GPT-4o and GPT-4o-Mini exhibit their superiority in generating mutations with smaller semantic distances with real bugs.
+Existing other LLM-based approaches achieve an overall lower Ochiai coefficient of 17.7% and 24.1%, respectively.
+The LLM-based approaches using our prompt achieve notably higher Ochiai coefficients.
+DeepSeek-V3-761b attains the highest value at 59.8%, followed by GPT-4o at 56.2% and GPT-4o-Mini at 54.1%.
+The LLMs based on our prompt outperform other existing methods, not only traditional approaches but also existing LLM-based ones, showing their superiority in generating mutations that are semantically closer to real bugs.
 High Ochiai Bug Count: Following the study of Ojdanic et al.
@@
 For each project, we count the number of bugs that achieved a high Ochiai coefficient by each approach, shown as Table 6.
-The overall trend is similar with Real Bug Detectability, which indicates that the LLM-based appraoches, GPT-4o and GPT-4o-Mini, achieve the highest number of bugs across most projects.
-DeepSeek also exhibits strong performance, closely following GPT-4o in many cases.
-All traditional approaches exhibit significantly fewer high Ochiai bugs.
+The overall trend is similar to Real Bug Detectability, which indicates that the LLM DeepSeek-V3-761b, based on our prompt, achieves the highest number of bugs across most projects.
+The closed-source LLMs GPT-4o and GPT-4o-Mini also exhibit strong performance, closely following DeepSeek-V3-761b in many cases.
+Compared with our approach, all traditional approaches exhibit significantly fewer high Ochiai bugs.
 Interestingly, among traditional approaches, Major outperforms LEAM and BERT across most projects.
-| Project | Bug # | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
-| Chart | 26 | 9 | 9 | 3 | 5 | 10 | 13 | 12 | 12 | 7 | 3 |
-| Lang | 65 | 15 | 34 | 13 | 18 | 32 | 35 | 37 | 34 | 12 | 23 |
-| Time | 27 | 8 | 14 | 8 | 4 | 8 | 12 | 12 | 11 | 4 | 7 |
-| Math | 106 | 20 | 42 | 13 | 27 | 46 | 56 | 48 | 47 | 19 | 33 |
-| Mockito | 38 | 1 | 12 | 5 | 3 | 14 | 13 | 17 | 14 | 3 | 3 |
-| Closure | 133 | 12 | 25 | 40 | 14 | 38 | 69 | 56 | 45 | 21 | 26 |
-| Cli | 39 | 3 | 3 | 13 | 8 | 18 | 22 | 23 | 20 | 4 | 10 |
-| Codec | 18 | 3 | 1 | 3 | 3 | 2 | 8 | 6 | 7 | 3 | 2 |
-| Csv | 16 | 4 | 2 | 1 | 2 | 4 | 10 | 10 | 8 | 3 | 1 |
-| Gson | 18 | 2 | 1 | 4 | 3 | 2 | 8 | 7 | 6 | 3 | 2 |
-| JacksonCore | 26 | 2 | 3 | 9 | 8 | 11 | 14 | 12 | 11 | 5 | 6 |
-| Jsoup | 93 | 16 | 6 | 32 | 20 | 46 | 39 | 39 | 41 | 25 | 19 |
-| ConDefects | 246 | 101 | 150 | 132 | 134 | 175 | 179 | 181 | 177 | 112 | 148 |
-| ALL | 851 | 196 | 302 | 276 | 249 | 406 | 478 | 460 | 433 | 221 | 283 |
+| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Gs. | Csv | Ja. | Js. | ConD. | All |
+| #Bug | 26 | 65 | 27 | 106 | 38 | 133 | 39 | 18 | 16 | 18 | 26 | 93 | 246 | 851 |
+| PIT | 9 | 15 | 8 | 20 | 1 | 12 | 3 | 3 | 4 | 2 | 2 | 16 | 101 | 196 |
+| Major | 9 | 34 | 14 | 42 | 12 | 25 | 3 | 1 | 2 | 1 | 3 | 6 | 150 | 302 |
+| LEAM | 3 | 13 | 8 | 13 | 5 | 40 | 13 | 3 | 1 | 4 | 9 | 32 | 132 | 276 |
+| BERT | 5 | 18 | 4 | 27 | 3 | 14 | 8 | 3 | 2 | 3 | 8 | 20 | 134 | 249 |
+| BF (GPT-4o) | 9 | 15 | 3 | 13 | 1 | 13 | 3 | 2 | 1 | 4 | 2 | 14 | 71 | 151 |
+| LLMph (GPT-4o) | 6 | 25 | 7 | 24 | 6 | 22 | 9 | 2 | 3 | 3 | 5 | 12 | 81 | 205 |
+| LLMut (GPT-3.5) | 10 | 32 | 8 | 46 | 14 | 38 | 18 | 2 | 4 | 2 | 11 | 46 | 175 | 406 |
+| LLMut (GPT-4o) | 13 | 35 | 12 | 56 | 13 | 69 | 22 | 8 | 10 | 8 | 14 | 39 | 179 | 478 |
+| LLMut (GPT-4o-M) | 12 | 37 | 12 | 48 | 17 | 56 | 23 | 6 | 10 | 7 | 12 | 39 | 181 | 460 |
+| LLMut (DS-761b) | 13 | 40 | 15 | 62 | 21 | 79 | 27 | 8 | 10 | 12 | 17 | 59 | 146 | 509 |
+| LLMut (DS-236b) | 12 | 34 | 11 | 47 | 14 | 45 | 20 | 7 | 8 | 6 | 11 | 41 | 177 | 433 |
+| LLMut (SC-16b) | 7 | 12 | 4 | 19 | 3 | 21 | 4 | 3 | 3 | 3 | 5 | 25 | 112 | 221 |
+| LLMut (CL-13b) | 3 | 23 | 7 | 33 | 3 | 26 | 10 | 2 | 1 | 2 | 6 | 19 | 148 | 283 |
 Note: The mutation generation approaches that achieve the best performance in one project are highlighted in grey background color.
+BF: BugFarm. LLMph: LLMorpheus. LLMut: our prompt. Moc.: Mockito. Clo.: Closure. Co.: Codec. Gs.: Gson. Ja.: JacksonCore. Js.: Jsoup. ConD: ConDefects.
 Coupling Rate: The overall Coupling Rate for each mutation generation approach is shown in the last row of Table 4.
 The traditional approaches, PIT, Major, LEAM, and BERT, exhibit lower coupling rates than LLM-based approaches, with coupling rates of 23.5%, 38.9%, 38.1%, and 47.3%, respectively.
-Among the LLM-based approaches, GPT-4o achieves the highest coupling rate, with a value of 50.6%.
-It is followed closely by GPT-4o-Mini and DeepSeek, with coupling rates of 49.5% and 49.5%, respectively.
-Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly GPT-4o, in generating highly coupled mutations that better mimic real bugs.
-| Category | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
-| Strong | 10.1% | 9.6% | 4.9% | 7.8% | 7.0% | 6.8% | 6.2% | 6.4% | 8.0% | 6.8% |
-| Strong Addi. | 4.7% | 5.8% | 4.3% | 5.2% | 6.7% | 10.2% | 12.1% | 10.3% | 5.7% | 6.0% |
-| Partial | 1.8% | 1.1% | 0.5% | 0.8% | 1.2% | 1.2% | 0.8% | 1.0% | 2.7% | 0.9% |
-| Partial Addi. | 6.9% | 22.4% | 28.4% | 33.5% | 32.0% | 32.5% | 30.4% | 31.8% | 17.6% | 31.9% |
-| Not Subs. | 30.6% | 13.2% | 26.7% | 20.4% | 23.3% | 23.1% | 21.3% | 21.9% | 16.7% | 21.3% |
-| Not Detected | 45.9% | 47.9% | 35.2% | 32.3% | 29.8% | 26.3% | 29.2% | 28.6% | 49.3% | 33.1% |
+Among the LLM-based approaches, LLMorpheus achieves the highest coupling rate, with a value of 51.7%.
+It is followed closely by DeepSeek-V3-761b and GPT-4o-Mini based on our prompt, with coupling rates of 50.8% and 50.6%, respectively.
+Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly LLMorpheus, DeepSeek-V3-761b, and GPT-4o, in generating overall highly coupled mutations that better mimic real bugs.
+| Category | Strong | Strong Addi | Partial | Partial Addi | Not Subs. | Not Detected |
+| PIT | 10.1% | 4.7% | 1.8% | 6.9% | 30.6% | 45.9% |
+| Major | 9.6% | 5.8% | 1.1% | 22.4% | 13.2% | 47.9% |
+| LEAM | 4.9% | 4.3% | 0.5% | 28.4% | 26.7% | 35.2% |
+| BERT | 7.8% | 5.2% | 0.8% | 33.5% | 20.4% | 32.3% |
+| BugFarm (GPT-4o) | 8.4% | 5.4% | 0.6% | 33.6% | 28.3% | 23.8% |
+| LLMorpheus (GPT-4o) | 9.3% | 4.9% | 2.1% | 35.4% | 23.6% | 24.6% |
+| LLMut (GPT-3.5) | 7.0% | 6.7% | 1.2% | 32.0% | 23.3% | 29.8% |
+| LLMut (GPT-4o) | 6.8% | 10.2% | 1.2% | 32.5% | 23.1% | 26.3% |
+| LLMut (GPT-4o-M) | 6.2% | 12.1% | 0.8% | 30.4% | 21.3% | 29.2% |
+| LLMut (DS-761b) | 11.2% | 4.6% | 1.9% | 33.1% | 24.1% | 25.1% |
+| LLMut (DS-236b) | 6.4% | 10.3% | 1.0% | 31.8% | 21.9% | 28.6% |
+| LLMut (SC-16b) | 8.0% | 5.7% | 2.7% | 17.6% | 16.7% | 49.3% |
+| LLMut (CL-13b) | 6.8% | 6.0% | 0.9% | 31.9% | 21.3% | 33.1% |
 Note: Strong - Strong Substitution, Strong Addi.
@@
 Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutation of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected.
-Note that the first four categories are used to calculate the Coupling Rate.
+Note that the first four categories are used to calculate the Coupling Rate, indicating these mutations semantically overlap with their corresponding real bugs.
 Table 7 presents the distribution of each generation approach.
@@
 However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutations.
-In contrast, LLM-based approaches such as GPT-4o and GPT-4o-Mini, demonstrate stronger coupling capabilities across multiple categories.
-These approaches achieve significantly higher rates in the Partial Substitution + Additional Tests Fail category, with values of 32.5% and 30.4%, respectively.
-Furthermore, they outperform traditional methods in the Strong Substitution + Additional Tests Fail category, with GPT-4o and GPT-4o-Mini achieving 10.2% and 12.1%, respectively.
+In contrast, LLM-based approaches such as DeepSeek-V3-761b, GPT-4o, and GPT-4o-Mini, demonstrate stronger coupling capabilities across multiple categories.
+For instance, DeepSeek-V3-761b achieves the highest proportion in the Strong category at 11.2%, while GPT-4o-Mini leads in the Strong + Additional Tests Fail category, with 12.1%.
 Overall, while the traditional rule-based approaches generate a larger proportion of strong substitution mutations, LLM-based approaches demonstrate a stronger ability to generate diverse and highly coupled mutations.
@@
 AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element).
-For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang333https://pypi.org/project/javalang, to observe which new AST node types are introduced.
+For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang555https://pypi.org/project/javalang, to observe which new AST node types are introduced.
 A larger number of newly introduced AST node types indicates richer diversity.
@@
 The figure reveals that LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches.
-The best LLM-based approach, DeepSeek, exhibits the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types.
-Additionally, all six LLMs are significantly higher than all other approaches in AST node types.
+The best LLM-based approaches, DeepSeek-V3-761b and DeepSeek-Coder-V2-236b, exhibit the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types.
+Additionally, all six LLMs based on our prompt are significantly higher than all other approaches in AST node types.
+The existing LLM-based approaches, BugFarm and LLMorpheus, introduce 31 and 21 new AST node types, respectively.
+While BugFarm outperforms traditional approaches, it still lags behind ours.
 We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 2), as shown by the “Examples” in Figure 2.
@@
 This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
-| Top K | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
-| Del. | 15.3% | 48.9% | 0.4% | 5.3% | 2.1% | 0.5% | 1.0% | 0.3% | 0.1% |
-| 1st | ST(78.6%) | LT(23.2%) | LT(29.4%) | LT(39.2%) | LT(45.4%) | LT(43.3%) | LT(38.8%) | LT(22.7%) | LT(47.8%) |
-| 2nd | LT(21.4%) | MR(14.4%) | SE(25.7%) | BO(17.3%) | BO(16.6%) | BO(21.6%) | BO(19.1%) | BO(13.8%) | BO(25.7%) |
-| 3rd | —— | CT(9.6%) | MR(24.5%) | MI(9.9%) | CC(8.8%) | CC(7.6%) | MI(10.1%) | MR(13.1%) | MI(9.3%) |
-Del.: The proportion of deletion mutations. LT: Literal, BO: BinaryOperation, MI: MethodInvocation, MR: MemberReference, SE: StatementExpression, ST: Statement.
-We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the top three most frequent AST node types.
+| Approach | Deletion | 1st | 2nd | 3rd |
+| Major | 15.3% | ST (78.6%) | LT (21.4%) | —— |
+| LEAM | 48.9% | LT (23.2%) | MR (14.4%) | CT (9.6%) |
+| BERT | 0.4% | LT (29.4%) | SE (25.7%) | MR (24.5%) |
+| BugFarm (GPT-4o) | 5.5% | LT (41.9%) | BO (12.2%) | MI (9.0%) |
+| LLMorpheus (GPT-4o) | 0.3% | LT (24.0%) | BO (23.7%) | MR (15.8%) |
+| LLMut (GPT-3.5) | 5.3% | LT (39.2%) | BO (17.3%) | MI (9.9%) |
+| LLMut (GPT-4o) | 2.1% | LT (45.4%) | BO (16.6%) | CC (8.8%) |
+| LLMut (GPT-4o-M) | 0.5% | LT (43.3%) | BO (21.6%) | CC (7.6%) |
+| LLMut (DS-761b) | 3.8% | LT (44.2%) | BO (15.6%) | CC (8.6%) |
+| LLMut (DS-236b) | 1.0% | LT (38.8%) | BO (19.1%) | MI (10.1%) |
+| LLMut (SC-16b) | 0.3% | LT (22.7%) | BO (13.8%) | MR (13.1%) |
+| LLMut (CL-13b) | 0.1% | LT (47.8%) | BO (25.7%) | MI (9.3%) |
+Deletion: The proportion of deletion mutations. BO: BinaryOperation. CC: ClassCreator. CT: Cast. LT: Literal. MI: MethodInvocation. MR: MemberReference. SE: StatementExpression. ST: Statement.
+We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the proportion of the mutations deleting the target code element and the top three most frequent AST node types.
 For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutation ratio, with 48.9%, followed by Major at 15.3%.
@@
 The distribution of edit distances for each approach is presented in Table 9, providing insights into the complexity and diversity of the generated mutations.
-| Distance | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
-| 1 | 69.2% | 58.5% | 73.8% | 52.4% | 50.0% | 47.3% | 45.1% | 36.0% | 46.4% |
-| 2 | 27.8% | 23.6% | 8.9% | 36.4% | 38.8% | 41.9% | 44.2% | 39.1% | 46.6% |
-| 3 | 1.8% | 7.3% | 2.7% | 6.8% | 7.2% | 6.3% | 6.5% | 8.0% | 3.8% |
-| 4 | 0.6% | 3.4% | 14.2% | 2.4% | 2.2% | 2.5% | 2.6% | 6.0% | 1.8% |
-| 5 | 0.4% | 1.0% | 0.1% | 0.9% | 0.8% | 0.8% | 0.8% | 3.7% | 0.5% |
-| 6 | 0.2% | 6.1% | 0.4% | 1.0% | 1.0% | 1.2% | 0.8% | 7.2% | 1.0% |
+| Distance | 1 | 2 | 3 | 4 | 5 | 6 |
+| Major | 69.2% | 27.8% | 1.8% | 0.6% | 0.4% | 0.2% |
+| LEAM | 58.5% | 23.6% | 7.3% | 3.4% | 1.0% | 6.1% |
+| BERT | 73.8% | 8.9% | 2.7% | 14.2% | 0.1% | 0.4% |
+| BugFarm (GPT-4o) | 44.7% | 43.5% | 8.3% | 1.6% | 0.6% | 1.4% |
+| LLMorpheus (GPT-4o) | 54.2% | 25.7% | 14.6% | 3.7% | 0.9% | 0.8% |
+| LLMut (GPT-3.5) | 52.4% | 36.4% | 6.8% | 2.4% | 0.9% | 1.0% |
+| LLMut (GPT-4o) | 50.0% | 38.8% | 7.2% | 2.2% | 0.8% | 1.0% |
+| LLMut (GPT-4o-M) | 47.3% | 41.9% | 6.3% | 2.5% | 0.8% | 1.2% |
+| LLMut (DS-761b) | 46.1% | 44.2% | 6.8% | 1.5% | 0.4% | 0.9% |
+| LLMut (DS-236b) | 45.1% | 44.2% | 6.5% | 2.6% | 0.8% | 0.8% |
+| LLMut (SC-16b) | 36.0% | 39.1% | 8.0% | 6.0% | 3.7% | 7.2% |
+| LLMut (CL-13b) | 46.4% | 46.6% | 3.8% | 1.8% | 0.5% | 1.0% |
 From the table, we can find that all traditional approaches predominantly generate small-scale changes.
@@
 In contrast, LLM-based approaches exhibit a broader range of edit distances, reflecting their capability to introduce more complex transformations.
-For example, GPT-4o exhibits a more balanced distribution, with 38.8% of its mutations having an edit distance of 2, 7.2% with a distance of 3, and 2.2% with a distance of 4, indicating it generates both minor and moderate code changes.
+For example, DeepSeek-V3-761b exhibits a more balanced distribution, with 44.2% of its mutations having an edit distance of 2, 7.2% with a distance of 3, and 2.2% with a distance of 4, indicating it generates both minor and moderate code changes.
 3.2. RQ2: Performance on Validity
-To evaluate the validity of mutation generation approaches, we consider the aspects that may impact their validity, i.e., comparability rate, duplicate mutation rate, and equivalent mutation rate.
+To evaluate the validity of mutation generation approaches, we consider the aspects that may impact their validity, i.e., compilability rate, duplicate mutation rate, and equivalent mutation rate.
 This RQ is divided into the following sub-RQs.
 3.2.1. RQ2.1: Compilability
-The row Comp.
-Rate of Table 4 presents the overall Compilability Rate (CR) of each approach.
+The column Comp.
+Rate of Table 4 presents the overall Compilability Rate () of each approach.
 Among the traditional approaches, the rule-based approaches, PIT and Major, demonstrate superior compilability, with PIT achieving a perfect 100% , and Major following closely at 97.60%.
@@
 For example, the Java compiler’s analyzer rejects constructs like while(true).
-PIT uses the mutation operators more restrictedly and previously filters out non-compilable mutations.
+PIT uses the mutation operators more restrictively and previously filters out non-compilable mutations.
 In contrast, LEAM and BERT, have rates at 35.0% and 22.5%, respectively.
 For LLM-based approaches, there is a considerable variation in compilability.
-GPT-4o achieves the highest compilability among LLMs at 76.40%, followed closely by GPT-4o-Mini and DeepSeek, with rates of 75.30% and 75.80%, respectively.
-Other LLMs such as GPT-3.5 and CodeLlama show moderate s of 62.50% and 71.20%, respectively, whereas models like StarChat fall behind with a of 38.50%.
-Compared to small models (i.e., LEAM and Bert), GPTs and DeepSeek show significantly higher compilation success rates.
+The existing LLM-based approaches, BugFarm and LLMorpheus, achieve relatively high rates of 74.1% and 67.8%, respectively.
+Among the LLMs using our prompt, DeepSeek-V3-761b attains the highest at 77.6%, followed closely by GPT-4o (76.4%), DeepSeek-Coder-V2-236b (75.8%), and GPT-4o-Mini (75.3%).
+Compared to small models (i.e., LEAM and BERT), GPTs and DeepSeek models show significantly higher compilation success rates.
 For example, the compilability rate of GPT-4o is over 30% higher than that of LEAM and BERT.
@@
 Therefore, a lower Duplicate Mutation Rate () is desirable, as it indicates a more efficient mutation generation process with minimal redundancy.
-The row Dup.
+The column Dup.
 Mut.
@@
 This is expected, as these approaches apply well-defined mutation rules that ensure each generated mutant introduces a meaningful change to the program.
-Other traditional approaches, i.e., LEAM and Bert, show relatively low s of 2.30% and 1.70%, respectively.
+Other traditional approaches, i.e., LEAM and BERT, show relatively low s of 2.3% and 1.7%, respectively.
 In contrast, LLMs generate a significant portion of duplicate mutations, exhibiting a higher .
-For the LLMs from OpenAI, GPT-3.5, GPT-4o, and GPT-4o-Mini exhibit moderate s of 10.30%, 7.80%, and 7.00%, respectively.
-The open-source LLMs such as StarCoder and DeepSeek demonstrate higher s of 12.80% and 7.70%, respectively, while CodeLlama shows the highest at 37.10%.
+For the LLMs from OpenAI, GPT-3.5, GPT-4o, and GPT-4o-Mini exhibit moderate s of 10.3%, 7.8%, and 7.0%, respectively.
+For the existing LLM-based approaches, BugFarm and LLMorpheus, exhibit similar s of 7.5% and 7.1%, respectively.
+The open-source LLMs such as StarCoder and DeepSeek-Coder-V2-236b demonstrate higher s of 12.8% and 7.7%, respectively, while CodeLlama shows the highest at 37.1%.
 3.2.3. RQ2.3: Equivalent Mutations
@@
 Note that PIT does not generate source-level mutations, so its mutations must be manually extracted from Java bytecode and transformed back into source-level representations, which is significantly more time-consuming.
-In total, 959 mutations were labeled, 508 of which were filtered by tests (i.e., killed), and the remaining 451 were manually checked.
-The entire process took approximately ten hours per person.
+In total, 1247 mutations were labeled, 709 of which were filtered by tests (i.e., killed), and the remaining 538 were manually checked.
+The entire process took approximately twelve hours per person.
 The manual verification requires checking the code context, which may involve multiple classes and methods.
@@
 The labeling results are available at our repo.
-| Metric | PIT | Major | LEAM | Bert | GPT-3.5 | GPT-4o | GPT-4o-Mini | DS-236b | SC-16b | CL-13b |
-| #Sampled | 97 | 98 | 96 | 94 | 96 | 96 | 96 | 96 | 94 | 96 |
-| #Test Filtered | 26 | 52 | 48 | 53 | 60 | 64 | 59 | 63 | 32 | 51 |
-| #Eq. Mut | 1 | 2 | 2 | 6 | 4 | 3 | 3 | 5 | 10 | 6 |
-| EMR | 1.0% | 2.0% | 2.1% | 6.4% | 4.2% | 3.1% | 3.1% | 5.2% | 10.6% | 6.3% |
+| Approach | #Sampled | #Test Filtered | #Eq. Mut. | EMR |
+| PIT | 97 | 26 | 1 | 1.0% |
+| Major | 96 | 52 | 2 | 2.1% |
+| LEAM | 96 | 48 | 2 | 2.1% |
+| BERT | 94 | 64 | 3 | 6.4% |
+| BugFarm (GPT-4o) | 96 | 67 | 3 | 3.1% |
+| LLMorpheus (GPT-4o) | 96 | 69 | 4 | 4.2% |
+| LLMut (GPT-3.5) | 96 | 60 | 4 | 4.2% |
+| LLMut (GPT-4o) | 96 | 64 | 3 | 3.1% |
+| LLMut (GPT-4o-M) | 96 | 59 | 3 | 3.1% |
+| LLMut (DS-761b) | 96 | 65 | 6 | 6.3% |
+| LLMut (DS-236b) | 96 | 63 | 5 | 5.2% |
+| LLMut (SC-16b) | 94 | 32 | 10 | 10.6% |
+| LLMut (CL-13b) | 96 | 51 | 6 | 6.3% |
 Table 10 presents the estimated Equivalent Mutation Rate (), reporting the number of sampled mutations (#Sampled), the number of equivalent mutations we identified (#Eq.
 Mut), and the for each approach.
-Among the traditional approaches, PIT achieves the lowest of 1.0%, followed by Major with an of 2.0%, indicating that rule-based techniques produce fewer equivalent mutations.
+Among the traditional approaches, PIT achieves the lowest of 1.0%, followed by Major with an of 2.1%, indicating that rule-based techniques produce fewer equivalent mutations.
 In contrast, the LLM-based approaches exhibit higher s, with variations across different models.
-StarChat has the highest s of 10.6%, while GPT-4o, GPT-4o-Mini, and DeepSeek show relatively lower s, ranging from 3.1% to 5.2%.
-Overall, although the OpenAI models such as GPT-4o and GPT-4o-Mini, tend to have slightly higher s than PIT and Major, the increase remains modest and within an acceptable range.
+StarChat has the highest s of 10.6%, while GPT-4o, GPT-4o-Mini, and DeepSeek-Coder-V2-236b show relatively lower s, ranging from 3.1% to 5.2%.
+Overall, although the OpenAI models such as GPT-4o and GPT-4o-Mini tend to have slightly higher s than PIT and Major, the increase remains modest and within an acceptable range.
 Note that the estimated of Major given by the recent study of Kushigian et al.
-(Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutation for manual judgment.
-We adopt the same experimental process and the result of Major is 2.0%, which is consistent with theirs considering the randomness introduced by sampling.
+(Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutations for manual judgment.
+We adopt the same experimental process and the result of Major is 2.0%, which is consistent with theirs, considering the randomness introduced by sampling.
+Moreover, the definition of equivalent mutations in this paper is the set of syntactically different but semantically equivalent mutations, which may differ from existing studies (Tip et al., 2024).
 3.3. RQ3: Performance on Efficiency
 To evaluate the performance on efficiency, we record the average generation time for each approach.
-Average Generation Time: The row A.
-G.
-Time of Table 4 presents the results of Average Generation Time.
+Average Generation Time: The column Time of Table 4 presents the results of Average Generation Time in seconds.
 Among the traditional appraoches, PIT demonstrates the fastest generation time at 0.02 seconds per mutation, followed by Major, which takes 0.08 seconds, making them the most efficient approaches.
 In contrast, LEAM and BERT take 3.06 seconds and 2.34 seconds, respectively.
-Among the LLMs, the LLMs of OpenAI (i.e., GPT-3.5, GPT-4o, and GPT-4o-Mini) show relatively faster generation, ranging from 1.31 to 1.79 seconds.
-While the open-source LLMs StarChat, CodeLlama, and DeepSeek require 7.53, 9.06, and 4.25 seconds, respectively.
+The existing LLM-based approaches, Among the LLMs, the LLMs of OpenAI (i.e., GPT-3.5, GPT-4o, and GPT-4o-Mini) show relatively faster generation, ranging from 1.31 to 1.79 seconds.
+The open-source LLMs StarChat, CodeLlama, and DeepSeek require 7.53, 9.06, and 4.25 seconds, respectively.
 The results suggest that the rule-based approaches are the most time-efficient for generating mutations, which are tens to hundreds of times faster than other approaches.
+Note that the time costs of self-deployed open-source models typically depend on the types of GPUs used for execution.
+Additionally, the response time of models accessed via network APIs may vary across service providers and can be further affected by network fluctuations.
+We further discuss these potential impacts in the Section 5.
+| Approach | Avg. Input Token | Avg. Output Token |
+| BugFarm (GPT-4o) | 200 | 368 |
+| LLMorpheus (GPT-4o) | 370 | 74 |
+| LLMut (GPT-3.5) | 126 | 59 |
+| LLMut (GPT-4o) | 126 | 59 |
+| LLMut (GPT-4o-M) | 126 | 59 |
+| LLMut (DS-761b) | 126 | 58 |
+| LLMut (DS-236b) | 126 | 59 |
+| LLMut (SC-16b) | 126 | 63 |
+| LLMut (CL-13b) | 126 | 58 |
+Average Token: Table 11 presents the average number of input and output tokens used by each LLM-based approach to generate one mutation.
+The existing LLM-based approaches, BugFarm and LLMorpheus, consume 200 and 370 input tokens, and 368 and 74 output tokens per mutation, respectively.
+LLMs using our prompt consume substantially fewer tokens, averaging 126 input tokens and 59–63 output tokens per mutation.
+As all LLMs share identical inputs, their input token counts remain the same.
 3.4. RQ4: Categorizations of Non-Compilable Mutations
@@
 Code Structural Destruction: This type of error is due to illegal code structures, such as unmatched parentheses or braces.
-Incorrect Method Parameters: this type of error is due to invoking a method with illegal parameters, including invoking by wrong numbers of parameters or unmatched type.
-Usage of Unknown Variables: this type of error is due to using unknown variables under the code context.
-Usage of Unknown Types: this type of error is due to using unknown type names under the code context.
+Incorrect Method Parameters: This type of error is due to invoking a method with illegal parameters, including invoking with the wrong number of parameters or mismatched types.
+Usage of Unknown Variables: This type of error is due to using unknown variables in the code context.
+Usage of Unknown Types: This type of error is due to using unknown type names in the code context.
 Type Mismatch: This type of error is due to using illegal types when applying a certain operation, such as using objects to perform arithmetic operations.
-Incorrect Initialization: this type of error is due to using illegal forms of constrictors or deleting necessary initialization statements.
+Incorrect Initialization: This type of error is due to using illegal forms of constructors or deleting necessary initialization statements.
 Incorrect Location: This type of error is due to a mutation being applied to an incorrect code location.
@@
 To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by all nine mutation generation approaches in Table 4, as shown in Figure 3.
-For most approaches (except Major), the code locations with MemberReference and MethodInvocation are the top-2 prevent AST node types.
-For example, for GPT LLMs and DeepSeek, there about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation.
+For most approaches (except Major and BugFarm), the code locations with MemberReference and MethodInvocation are the top-2 most-prevent AST node types.
+For example, for GPT and DeepSeek models, about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation.
 This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references.
-If any required method or member is missing or incorrectly specified, it can easily lead to non-compilable mutations.
+If any required method or member is missing or incorrectly specified, it can easily lead to non-compileable mutations.
 The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure.
-Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 1.11% (DeepSeek), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% (Bert), and 16.61% (Major) of all their non-compilable mutations, respectively.
+Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 3.4% (DeepSeek-V3-761b), 1.11% (DeepSeek-Coder-V2-236b), 3.4% (BugFarm), 0.8% (LLMorpheus), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% (BERT), and 16.61% (Major) of all their non-compilable mutations, respectively.
 Thus, for LLMs, deletion is not the major reason for non-compilation.
-Thus for LLMs, deletion is not the major reason for non-compilation.
+3.5. RQ5: Distributions of Alive Mutations
+As shown in Table 4, the mutation scores of rule-based approaches (i.e., PIT and Major) are lower than LLMs, indicating their mutations are harder to detect.
+Therefore, we intend to analyze the reasons for survival mutations.
+To address this RQ, we adopt a setting similar to the analysis of equivalent mutations.
+Specifically, for each approach, we randomly sample its alive mutations and manually classify them.
+We employ the categories of alive mutation causes defined in the study by Just et al.
+(Just et al., 2014a) as follows.
+(1) Not covered, i.e., the mutation is located in code never executed by the test suite. We refer to the mutation set of this category as .
+(2) Covered but no state infection, i.e., the mutation is executed but does not propagate any state change. We refer to the mutation set of this category as .
+(3) State infection without detection, i.e., the mutation alters the program state but any test oracle does not observe the deviation. We refer to the mutation set of this category as .
+These categories are mutually exclusive, with the following relationships: , , and .
+This classification enables us to analyze the underlying reasons for alive mutations and further evaluate the extent to which different approaches contribute to strengthening the test suite.
+Given an alive mutation, we first determine whether it is equivalent; if so, it is discarded.
+For non-equivalent mutations, we then check whether they are uncovered.
+If a mutation is covered, we further examine whether it alters the program state, i.e., whether any memory values are modified.
+Finally, the remaining alive mutations that are covered but do not lead to detectable differences are classified as not detected.
+This stepwise process ensures that every alive mutation is uniquely assigned to one category (equivalent, uncovered, not injecting state, or not detected).
+| Approach | #Sampled | #Eq. Mut. | NC | NI | ND | NC% | NI% | ND% |
+| PIT | 96 | 3 | 43 | 27 | 23 | 46.24% | 29.03% | 24.73% |
+| Major | 96 | 4 | 42 | 30 | 20 | 45.65% | 32.61% | 21.74% |
+| LEAM | 95 | 7 | 59 | 16 | 13 | 67.05% | 18.18% | 14.77% |
+| BERT | 88 | 8 | 37 | 24 | 19 | 46.25% | 30.00% | 23.75% |
+| BugFarm (GPT-4o) | 94 | 9 | 62 | 11 | 12 | 72.94% | 12.94% | 14.12% |
+| LLMopheus (GPT-4o) | 95 | 11 | 26 | 44 | 14 | 30.95% | 52.38% | 16.67% |
+| LLMut (GPT-3.5) | 95 | 12 | 66 | 12 | 5 | 79.52% | 14.46% | 6.02% |
+| LLMut (GPT-4o) | 95 | 11 | 53 | 16 | 15 | 63.10% | 19.05% | 17.86% |
+| LLMut (GPT-4o-M) | 95 | 12 | 51 | 15 | 17 | 61.45% | 18.07% | 20.48% |
+| LLMut (DS-761b) | 95 | 12 | 45 | 16 | 22 | 54.22% | 19.28% | 26.51% |
+| LLMut (DS-236b) | 95 | 13 | 54 | 15 | 13 | 65.85% | 18.29% | 15.85% |
+| LLMut (SC-16b) | 93 | 16 | 62 | 11 | 4 | 80.52% | 14.29% | 5.19% |
+| LLMut (CL-13b) | 96 | 11 | 68 | 12 | 5 | 80.00% | 14.12% | 5.88% |
+As shown in Table 13, the distributions of alive mutations differ significantly across approaches, where the columns NC%, NI%, and ND% represent the proportions of , , and within all non-equivalent mutations, respectively.
+First, LLM-based methods exhibit a substantially higher proportion of (i.e., not covered) mutations, meaning that more of their alive mutations reside in previously uncovered code lines.
+For instance, among the alive mutations generated by GPT-3.5 and BugFarm, 79.52% and 72.94% fall into the category, respectively, whereas PIT and Major exhibit much lower proportions of 46.24% and 45.65%.
+This characteristic highlights the potential of LLMs in guiding test suite augmentation by directing attention to untested program areas.
+Second, traditional approaches, especially rule-based methods such as PIT and Major, produce a larger fraction of (i.e., not injected) and (i.e., not detected) mutations compared to LLM-based methods.
+For instance, PIT and Major exhibit proportions of 29.03% and 32.61%, respectively, whereas all LLMs based on our prompt remain below 20%.
+These types of alive mutations require not only generating inputs that execute the mutated code but also propagating and detecting subtle semantic differences, which is inherently more challenging than increasing coverage.
+Therefore, the overall implication is that LLM-generated alive mutations are more directly beneficial for strengthening unit tests, as they emphasize extending coverage, while rule-based approaches tend to stress the harder task of state propagation and detection.
 4. Discussion
@@
 Due to budget constraints, we selected a subset of 306 bugs from our dataset, comprising 10 randomly sampled bugs from each Defects4J project and all bugs from the ConDefects dataset.
-For the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the open-source DeepSeek-Coder-V2-236b, then switch different prompt templates.
-More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get four types of prompts, shown as follows:
+For the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the recent open-source model DeepSeek-V3-761b, then switch different prompt templates.
+More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get five types of prompts, shown as follows:
 Prompt 1 (P1): The Default Prompt. P1 is the default prompt, shown as Figure 1.
@@
 Prompt 3 (P3): P2 without Whole Java Method. P3 is based on P2 by removing the surrounding Java method code, only keeping the target code element to be mutated.
-Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of corresponding unit tests of the target method into Context.
-| Metric | GPT-3.5 | DS-236b | ||||||
-| P1 | P2 | P3 | P4 | P1 | P2 | P3 | P4 | |
-| Mut. # | 3681 | 3688 | 3350 | 3210 | 3803 | 3845 | 3802 | 3804 |
-| Mut. Score | 0.773 | 0.785 | 0.842 | 0.801 | 0.784 | 0.778 | 0.793 | 0.814 |
-| Real Bug Detec. | 79.4% | 78.4% | 72.2% | 72.5% | 84.6% | 84.0% | 81.4% | 82.6% |
-| Avg. Ochiai | 63.1% | 61.8% | 60.8% | 62.1% | 67.3% | 64.1% | 65.0% | 63.7% |
-| Coupling Rate | 52.1% | 52.2% | 47.4% | 49.6% | 53.1% | 51.9% | 52.9% | 50.7% |
-| Compilability Rate | 66.9% | 66.8% | 60.0% | 61.2% | 79.1% | 78.9% | 74.9% | 78.2% |
-| Dup. Mut. Rate | 11.4% | 13.3% | 14.9% | 12.7% | 10.9% | 9.4% | 9.8% | 12.7% |
-| Avg. Gen. Time | 1.789 | 1.782 | 1.809 | 1.884 | 4.252 | 4.254 | 4.199 | 4.397 |
-Table 12 presents a comparative evaluation of different prompts for GPT-3.5 and DeepSeek across multiple metrics.
-Among the four prompts, P1 consistently demonstrates superior performance across both models, achieving the highest Compilability Rates (66.9% for GPT-3.5 and 79.1% for DeepSeek) and the best Real Bug Detectability Rates (79.4% and 84.6%, respectively).
-P1 also yields the highest Avgerage Ochiai Coefficient across both models.
-These results overall indicate that providing LLMs with well-structured and appropriate information significantly enhances their performance (e.g., P1), whereas excessive information can lead to a decline in effectiveness (e.g., P4).
+Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of the corresponding unit tests of the target method into Context.
+Prompt 5 (P5): P1 with Whole Java File. P5 is a variant of P1, which uses the code of the entire Java file as the code context, rather than the entire method.
+| Metric | LLMut (GPT-3.5) | LLMut (DS-761b) | ||||||||
+| P1 | P2 | P3 | P4 | P5 | P1 | P2 | P3 | P4 | P5 | |
+| Mut. # | 3681 | 3688 | 3350 | 3210 | 3496 | 3851 | 3847 | 3834 | 3844 | 3878 |
+| Mut. Score | 0.773 | 0.785 | 0.842 | 0.801 | 0.694 | 0.765 | 0.750 | 0.760 | 0.753 | 0.776 |
+| Real Bug Detec. | 79.5% | 78.4% | 72.2% | 72.5% | 79.5% | 85.5% | 85.8% | 83.9% | 84.7% | 83.1% |
+| Avg. Ochiai | 63.1% | 61.8% | 60.8% | 62.1% | 62.3% | 67.5% | 65.6% | 66.4% | 65.8% | 66.4% |
+| Coupling Rate | 52.1% | 52.2% | 47.4% | 49.6% | 47.3% | 53.2% | 51.7% | 50.9% | 50.0% | 49.1% |
+| Compling Rate | 66.9% | 66.8% | 60.0% | 61.2% | 70.3% | 82.8% | 83.5% | 78.0% | 82.0% | 79.5% |
+| Dup. Mut. Rate | 11.4% | 13.3% | 14.9% | 12.7% | 16.9% | 7.3% | 6.4% | 8.5% | 7.0% | 8.2% |
+| Avg. Gen. Time | 1.789 | 1.782 | 1.09 | 1.884 | 1.907 | 2.573 | 2.571 | 2.488 | 2.601 | 2.596 |
+| Avg. # Token | 105 | 95 | 44 | 219 | 393 | 104 | 92 | 43 | 219 | 392 |
+Note: The prompt template that achieves the best performance is highlighted in grey background color.
+Table 14 presents a comparative evaluation of different prompts for GPT-3.5 and DeepSeek-V3-761b across multiple metrics.
+Across both models, P1 delivers the best overall performance on all effectiveness metrics (i.e., real bug detectability, average Ochiai, and coupling rate), demonstrating its strength in generating mutations that are semantically closer to real bugs.
+For the validity metrics, P1 ranks second in compilation rate on GPT-3.5 and first on DeepSeek-V3-761b, while achieving the lowest duplicate mutation rate on GPT-3.5 and third-lowest on DeepSeek-V3-761b.
+Notably, compared with P5, which provides the entire file as context, P1 relies only on the full method yet achieves higher semantic similarity to real bugs, with reduced generation time and token usage, while maintaining comparable compilation and duplicate mutation rates.
+Across P1 to P5, we observe a clear context–performance trade-off.
+Removing relevant context (P3) substantially degrades effectiveness, as seen in lower real bug detectability, average Ochiai, and coupling rates.
+Providing a focused, method-level context (P1 and P2) delivers the best overall balance of effectiveness and validity.
+However, adding excessive context beyond the focal method (P4 and P5) does not yield further improvements and often increases token usage and generation time, while in some cases even reducing key effectiveness metrics.
+This suggests that moderate, relevant context benefits LLM mutation generation, whereas both insufficient and excessive context hurt efficiency and quality.
 We further explore the performance similarity of different prompts.
@@
 We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the -values are all below 0.05.
+Although P3 achieves the most balanced overall performance, the differences among prompts are not statistically significant.
 4.1.2. Experiment Setup of the Context Length
@@
 | Dup. Mut. Rate | 10.8% | 12.5% | 11.9% |
-| Settings | Spea. Coef. | -value | Pear. Coef. | -value |
+| Settings | Spea. Coef | -value | Pear. Coef | -value |
 | GPT-3.5 P1 v.s. GPT-3.5 P2 | 0.9762 | 3.31E-05 | 0.9999 | 1.35E-31 |
 | GPT-3.5 P1 v.s. GPT-3.5 P3 | 0.9040 | 2.00E-03 | 0.9999 | 8.66E-23 |
-| GPT-3.5 P1 v.s. GPT-3.5 P4 | 0.9520 | 2.60E-04 | 0.9999 | 1.45E-25 |
-| DS-236b P1 v.s. DS-236b P2 | 1.0000 | 0.00E+00 | 0.9999 | 1.94E-31 |
-| DS-236b P1 v.s. DS-236b P3 | 0.9999 | 3.31E-05 | 0.9999 | 3.71E-30 |
-| DS-236b P1 v.s. DS-236b P4 | 0.9762 | 3.31E-05 | 0.9999 | 1.83E-27 |
-| 1-Line v.s. 2-Line | 0.9762 | 3.31E-05 | 0.9980 | 2.12E-08 |
-| 2-Line v.s. 3-Line | 0.9762 | 3.31E-05 | 0.9997 | 1.04E-10 |
-| 1-Line v.s. 3-Line | 0.9524 | 2.60E-04 | 0.9978 | 2.69E-08 |
-| QB-6a (Default) v.s. QB-3 | 0.9701 | 6.55E-05 | 0.9996 | 1.70E-10 |
-| QB-6a (Default) v.s. QB-3b | 0.9762 | 3.31E-05 | 0.9984 | 1.00E-08 |
-| QB-6a (Default) v.s QB-9 | 0.9762 | 3.31E-05 | 0.9985 | 7.98E-09 |
+| GPT-3.5 P1 v.s. GPT-3.5 P4 | 0.9520 | 2.60E-04 | 0.9999 | 1.94E-31 |
+| GPT-3.5 P1 v.s. GPT-3.5 P5 | 0.9762 | 3.31E-05 | 0.9999 | 1.57E-26 |
+| DS-671b P1 v.s. DS-671b P2 | 1.0000 | 0.00E+00 | 0.9999 | 1.08E-31 |
+| DS-671b P1 v.s. DS-671b P3 | 1.0000 | 0.00E+00 | 0.9999 | 2.93E-29 |
+| DS-671b P1 v.s. DS-671b P4 | 1.0000 | 0.00E+00 | 0.9999 | 4.95E-31 |
+| DS-671b P1 v.s. DS-671b P5 | 1.0000 | 0.00E+00 | 0.9999 | 1.51E-30 |
+| 1-Line vs. 2-Line | 0.9762 | 3.31E-05 | 0.9980 | 2.12E-08 |
+| 2-Line vs. 3-Line | 0.9762 | 3.31E-05 | 0.9997 | 1.04E-10 |
+| 1-Line vs. 3-Line | 0.9524 | 2.60E-04 | 0.9978 | 2.69E-08 |
+| QB-6a (Default) v.s. QB-6b | 0.9429 | 4.81E-03 | 0.9984 | 1.00E-08 |
+| QB-6a (Default) v.s. QB-3 | 0.9856 | 3.09E-05 | 0.9999 | 1.62E-13 |
+| QB-6a (Default) v.s. QB-4 | 0.9643 | 4.54E-04 | 0.9999 | 7.28E-23 |
+| QB-6a (Default) v.s. QB-5 | 0.9643 | 4.54E-04 | 0.9999 | 9.62E-24 |
+| QB-6a (Default) v.s. QB-7 | 1.0000 | 0.0E+00 | 0.9999 | 9.39E-16 |
+| QB-6a (Default) v.s. QB-8 | 1.0000 | 0.0E+00 | 0.9999 | 3.34E-15 |
+| QB-6a (Default) v.s. QB-9 | 1.0000 | 0.00E+00 | 0.9999 | 5.01E-14 |
+| QB-6a (Default) v.s. QB-10 | 0.9856 | 3.09E-04 | 0.9999 | 2.89E-13 |
 | QB-6a (Default) v.s D4J-6 | 0.9524 | 2.60E-04 | 0.9980 | 2.13E-08 |
-QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples. D4J-6 involves random 6 examples from Defects4J.
+QB-N: Randomly samples N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b randomly samples 6 other examples. D4J-6 randomly samples 6 examples from Defects4J.
 To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches.
@@
 Similarly, the middle part of Table 16 shows the performance similarity among different contexts.
-We can see that their mutual similarity exceeds 0.95 and the -values are all below 0.05.
+We can see that their mutual similarity exceeds 0.95, and the -values are all below 0.05.
 Different context lengths perform similarly, thus validating the setup of our experiment.
 4.1.3. Impact of Different Few-Shot Examples
-| Metric | QB-6a (Default) | QB-3 | QB-6b | QB-9 | D4J-6 |
-| Real Bug Det. | 0.917 | 0.933 | 0.917 | 0.917 | 0.900 |
-| Avg. Ochiai | 0.467 | 0.450 | 0.434 | 0.434 | 0.367 |
-| Coupling Rate | 0.136 | 0.128 | 0.132 | 0.121 | 0.139 |
-| Comp. Rate | 65.7% | 62.7% | 63.9% | 61.8% | 62.1% |
-| Dup Mut. Rate | 13.7% | 12.8% | 12.8% | 13.8% | 12.5% |
-| Equivalent Mutation Rate | 1.3% | 1.6% | 1.3% | 1.1% | 1.6% |
-QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples. D4J-6 involves random 6 examples from Defects4J.
+| Metric | QB-6a | QB-3 | QB-4 | QB-5 | QB-6b | QB-7 | QB-8 | QB-9 | QB-10 | D4J-6 |
+| Real Bug Detec | 0.917 | 0.933 | 0.883 | 0.883 | 0.917 | 0.900 | 0.900 | 0.917 | 0.867 | 0.900 |
+| Avg. Ochiai | 0.467 | 0.450 | 0.400 | 0.433 | 0.433 | 0.467 | 0.433 | 0.433 | 0.400 | 0.367 |
+| Coupling Rate | 0.136 | 0.128 | 0.126 | 0.127 | 0.132 | 0.112 | 0.113 | 0.121 | 0.136 | 0.139 |
+| Comp. Rate | 67.5% | 62.7% | 65.6% | 66.8% | 63.9% | 65.1% | 66.0% | 61.8% | 64.9% | 62.1% |
+| Dup.Mut.Rate | 13.7% | 12.8% | 12.6% | 9.4% | 12.8% | 11.4% | 13.2% | 13.8% | 13.6% | 13.9% |
+| Avg. # Token | 106 | 98 | 102 | 103 | 106 | 108 | 109 | 110 | 113 | 107 |
+QB-N: Randomly samples N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b randomly samples 6 other examples. D4J-6 randomly samples 6 examples from Defects4J.
 To measure the impact of different few-shot examples on the results, we conduct comparative experiments with various examples.
-First, we randomly select 3, 6, and 9 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples.
+First, we randomly select 3, 4, 5, 6, 7, 8, 9, and 10 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples.
 Additionally, despite the risk of data leakage, we select 6 examples from the Defects4J dataset.
@@
 We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the -values are all below 0.05.
-Therefore, we can conclude that although different few-shot examples have perturbation, their performance is very similar.
+Therefore, despite variations in few-shot examples, there is no statistically significant evidence indicating superior performance among the prompts.
 4.1.4. Performance of Using the Same Number of Mutations
 In Section 3, we did not restrict the number of mutations generated by each approach.
-To study the performance of each approach under fixed quantity conditions, we followed the setting of existing study (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by Bert (Degiovanni and Papadakis, 2022), which is 11397.
+To study the performance of each approach under fixed quantity conditions, we followed the setting of the existing studies (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by BERT (Degiovanni and Papadakis, 2022), which is 11397.
+This ensures a fair comparison across approaches by controlling for differences in mutation counts, thereby eliminating the potential bias that methods generating more mutations might appear to perform better simply due to quantity rather than quality.
 For counts exceeding this number, we randomly selected the specified number of mutations for analysis.
@@
 Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
-| Metric | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
-| Real Bug Det. | 31.6% | 68.8% | 64.6% | 66.0% | 82.2% | 87.4% | 88.6% | 87.0% | 33.9% | 68.5% |
-| Coupling Rate | 17.4% | 45.4% | 37.2% | 43.6% | 45.6% | 48.8% | 47.9% | 47.9% | 27.1% | 43.6% |
-| Ochiai Coefficient | 12.6% | 22.2% | 14.1% | 19.0% | 26.5% | 36.0% | 31.8% | 34.1% | 15.4% | 16.9% |
-| Comp. Rate | — | 96.9% | 54.1% | 26.1% | 61.2% | 75.5% | 74.3% | 74.6% | 12.4% | 69.5% |
-| Dup Mut. Rate | — | 0.0% | 2.6% | 1.9% | 9.9% | 7.7% | 7.1% | 7.6% | 9.6% | 38.3% |
+| Name | R. B. Detec. | Ochiai Coeffi. | Coup. Rate | Comp. Rate | Dup. Rate |
+| PIT | 31.6% | 12.6% | 17.4% | — | — |
+| Major | 68.8% | 22.2% | 45.4% | 96.9% | 0.0% |
+| LEAM | 64.6% | 14.1% | 37.2% | 54.1% | 2.6% |
+| BERT | 66.0% | 19.0% | 43.6% | 26.1 | 1.9% |
+| BugFarm (GPT-4o) | 58.6% | 17.7% | 47.9% | 74.2% | 7.5% |
+| LLMorpheus (GPT-4o) | 67.5% | 23.6% | 50.2% | 68.2% | 7.3% |
+| LLMut (GPT-3.5) | 82.2% | 26.5% | 45.6% | 61.2% | 9.9% |
+| LLMut (GPT-4o) | 87.4% | 36.0% | 48.8% | 75.5% | 7.7% |
+| LLMut (GPT-4o-M) | 88.6% | 31.8% | 47.9% | 74.3% | 7.1% |
+| LLMut (DS-761b) | 89.2% | 39.4% | 49.4% | 76.3% | 7.7% |
+| LLMut (DS-236b) | 87.0% | 34.1% | 47.9% | 74.6% | 7.6% |
+| LLMut (SC-16b) | 33.9% | 15.4% | 27.1% | 12.4% | 9.6% |
+| LLMut (CL-13b) | 68.5% | 16.9% | 43.6% | 69.5% | 38.3% |
+| Metric | LLMut (GPT-3.5) | LLMut (DS-761b) | ||||
+| Default | More Detailed | Default + 10 Err. | Default | More Detailed | Default + 10 Err. | |
+| Real Bug Detec | 0.917 | 0.900 | 0.900 | 0.950 | 0.950 | 0.933 |
+| Avg. Ochiai | 0.467 | 0.483 | 0.467 | 0.500 | 0.500 | 0.517 |
+| Coupling Rate | 0.136 | 0.129 | 0.117 | 16.9% | 17.8% | 16.0% |
+| Comp. Rate | 67.5% | 71.1% | 68.7% | 78.4% | 76.4% | 77.5% |
+| Dup.Mut.Rate | 13.7% | 14.1% | 10.4% | 5.1% | 4.1% | 6.2% |
+Note: The prompt achieving the best performance in terms of a metric is highlighted in grey background color.
+4.2. Impact of Prompt Turbulence on LLM-Based Mutation Generation
+In software engineering practice, even when using the same information sources, users may design prompts differently, for example, varying the level of detail in the requirements description or introducing minor language errors.
+To investigate the impact of such variations, we derived two prompt variants from the default prompt shown in Figure 1.
+More Detailed Output Specification: replacing the output specifications of the default prompt by explicitly requiring: Prohibit generating identical mutants, i.e., ensure ‘aftercode’ ‘precode’, and all mutants are unique.
+Write all output mutants in a JSON file.
+Default + 10 Errors: randomly injecting 10 English grammar mistakes into the default prompt, such as subject–verb agreement errors (e.g., it are) and noun number errors (e.g., two mutation).
+Following the same experimental setup as before, we compared these two variants using GPT-3.5 and DeepSeek-V3-761b, with the results summarized in Table 19.
+For GPT-3.5, the More Detailed Output Specification variant slightly improves the compilation rate (+3.6%) but also increases the duplicate mutation rate.
+In contrast, the Default + 10 Errors variant reduces real bug detection and coupling rate compared to the default prompt.
+Interestingly, the Default + 10 Errors variant, presumably the lowest in prompt quality, achieves the best performance on duplicate mutation rate.
+For DeepSeek-V3-761b, the differences across variants are minimal.
+For instance, the default and the More Detailed Output Specification prompts achieve identical real bug detection and average Ochiai, and the More Detailed Output Specification variant even shows a small increase in average Ochiai.
+Overall, these results suggest that prompt wording has a certain degree of randomness in its effect on mutation generation performance.
+To further verify this, we conducted similarity analysis across the three prompts for each metric and found no statistically significant superiority of any one prompt over the other two.
+| Metric | LLMut (GPT-3.5) | LLMut (GPT-4o) | LLMut (DS-671b) | |||
+| FOM | HOM | FOM | HOM | FOM | HOM | |
+| Real Bug Detec. | 0.917 | 0.900 | 0.933 | 0.933 | 0.950 | 0.933 |
+| Avg. Ochiai | 0.467 | 0.483 | 0.467 | 0.483 | 0.500 | 0.483 |
+| Coupling Rate | 0.136 | 0.249 | 0.151 | 0.307 | 0.169 | 0.306 |
+| Compilability Rate | 67.5% | 63.4% | 76.3% | 70.2% | 78.4% | 73.3% |
+| Dup. Mut. Rate | 13.7% | 5.2% | 5.3% | 3.5% | 5.1% | 3.1% |
+Note: The order achieving the best performance in terms of a metric is highlighted in grey background color.
+4.3. Effectiveness on High-Order Mutation Generation
+Previously, we only evaluated LLMs in generating first-order mutations (FOM).
+To explore the effectiveness of LLMs in generating high-order mutations (HOMs) (Jia and Harman, 2009; Wong et al., 2020), we conducted a preliminary experiment by evaluating the recent mainstream models, including GPT-3.5, GPT-4o, and DeepSeek-V3-761b.
+We modified the original prompt in Figure 1 to explicitly request HOM generation.
+Table 20 shows the results.
+For the effectiveness metrics, HOMs generally achieve average Ochiai scores and higher coupling rates across these models, indicating closer semantic alignment with real bugs.
+However, FOMs achieve higher real bug detectability on all models.
+For the validity metrics, HOMs exhibit lower duplicate mutation rates, suggesting improved mutation diversity.
+However, FOMs exhibit significantly higher compilability rates, which aligns with the expectation that larger code changes are more prone to introducing compilation errors.
+Our preliminary experiments suggest that strengthening LLM-based HOM generation is promising, as HOMs are semantically closer to real bugs and face more challenges in maintaining syntactic correctness.
 4.4. Implications
-Based on our findings, we discuss the pros and cons of traditional and LLM-based mutation generation approaches, and suggest potential improvements.
+Based on our findings, we discuss the pros and cons of traditional and LLM-based mutation generation approaches and suggest potential improvements.
 4.4.1. Advantages of LLM-based Approaches
@@
 In Mutation-2, the LLM transforms an else branch into an else if condition by synthesizing a logical condition using the integer variable seriesCount from the context, demonstrating an ability to introduce structurally meaningful changes that traditional approaches cannot achieve.
-Additionally, our study explores how different prompt designs, model choices, and contextual information influence mutation generation performance.
+Additionally, our study explores how different prompt designs, prompt turbulence, model choices, and contextual information influence mutation generation performance.
 We find that selecting models with stronger code generation capabilities (e.g., GPT-4o) and designing prompts carefully are critical to achieving optimal results.
+Moreover, our study demonstrates that LLMs have clear potential in generating HOMs.
+Compared with first-order mutations (FOMs), HOMs generally achieve higher coupling rates and average Ochiai scores, suggesting that they are semantically closer to real bugs, while suffering from lower compilability rates and reduced real bug detectability.
+The results highlight that strengthening LLM-based HOM generation is therefore a promising research direction.
 4.4.2. Advantages of Traditional Approaches
@@
 4.4.3. Challenges and Future Directions
-Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate substantial room for improvement.
+Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate that there is still substantial room for improvement.
 One major limitation is the high rate of non-compilable mutations, which limits their practical application in mutation testing.
 Our analysis reveals that common causes include syntactic errors in generated expressions, incorrect method invocations, and type mismatches.
-To address these challenges, future research should explore integrating LLM-based approaches with program analysis and repair techniques to refine mutation generation.
+To address these challenges, future research should explore the integration of LLM-based approaches with program analysis and repair techniques to refine mutation generation.
 Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutations before execution can enhance their reliability.
+Moreover, exploring the applications of LLM-generated mutation in the downstream tasks, such as mutation-based test case prioritization (MBTCP) and mutation-based fault localization (MBFL).
+Our study has shown that LLM-generated mutations often have a higher similarity with real bugs (e.g., have larger Ochiai coefficients), highlighting the potential of adopting them for these tasks.
+Another important direction is cross-language generalization.
+While our study primarily focuses on Java, mutation testing is equally critical in languages such as C/C++, Python, JavaScript, and Rust.
+Traditional mutation testing tools usually require language-specific implementations, leading to duplicated engineering efforts and limited portability.
+In contrast, modern LLMs possess inherent multilingual capabilities, as they are pre-trained on large-scale, cross-language code corpora.
+This enables them to naturally transfer mutation strategies across programming languages without retraining from scratch.
+Future work should investigate how well LLMs can exploit this cross-language knowledge, and how prompts, training data, or filtering strategies can be adapted to different language paradigms (e.g., strongly typed vs.
+dynamically typed).
+Advancing towards multilingual mutation generation would not only broaden the applicability of LLMs in software testing but also provide stronger evidence for their robustness in heterogeneous, real-world development environments.
+4.5. Limitations of Our Study
+Our study has the following limitations that are left for further exploration.
+First, while we evaluated a comprehensive set of static and behavioral metrics, we did not investigate more advanced dynamic indicators.
+For instance, the phenomenon of subsuming mutations (Garg et al., 2022; Diniz et al., 2021), where one mutation is always killed by any test that kills another, has been studied in the mutation testing literature as a way to assess better redundancy and the true fault-detection power of test suites.
+Incorporating such dynamic analyses would provide deeper insights into the effectiveness of LLM-generated mutations beyond the metrics reported in this paper.
+Additionally, while in the context of test suite enhancement, two mutants with identical failing test sets may be considered redundant, in mutation-based fault localization (Wang et al., 2025), the number of mutants per statement directly influences suspiciousness scores, since these are typically computed as averages across all associated mutants.
+Second, our experiments are limited to the Java programming language.
+Although Java is the most widely studied language in mutation testing, it does not cover the diversity of programming paradigms.
+Mutation behaviors may differ significantly in languages such as C/C++, Python, JavaScript, and Rust, due to differences in type systems, memory models, and runtime environments.
+Future research should examine the generalizability of LLM-based mutation generation across multiple languages.
+Notably, modern LLMs are inherently multilingual, trained on large-scale cross-language corpora, which makes cross-language mutation generation a promising direction for broadening the applicability of our findings.
+Third, our study relies on fixed few-shot examples, which restricts LLMs to producing context-specific mutations for the given code snippets.
+In addition, we impose no restrictions on the mutation operators applied by LLMs, whereas rule-based approaches rely on predefined and relatively simple operators.
+Consequently, it is still an open question how LLMs would perform if constrained to the same operator set as rule-based approaches.
 5. Threats to Validity
@@
 To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek).
-Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014a) may be covered in the training set of the studied LLMs.
-To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024) which includes programs and faults that were made after the release time of the LLMs we use and thus have limited data leakage risk.
+Moreover, our study further adopts two state-of-the-art LLM-based mutation generation approaches (i.e., BugFarm and LLMorpheus) as baselines.
+The metrics we used may threaten the conclusion of our study.
+Each metric has inherent limitations.
+For instance, Real Bug Detection Rate and Coupling Rate capture only partial aspects of similarity to real bugs: a method may detect many bugs but with weak coupling on each bug, or achieve strong coupling concentrated on a few bugs.
+Likewise, AST Edit Distance reflects structural differences, but a higher syntactic distance does not necessarily imply larger semantic changes.
+Additionally, the measurement of Average Generation Time may be affected by external factors such as GPU type, service provider implementation, and network fluctuations, introducing unavoidable noise.
+To mitigate the bias of individual metrics, we report a comprehensive set of complementary measures.
+Particularly, besides time cost, we also report token consumption, which provides a more stable and model-independent estimate of computational and financial cost.
+Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014b) may be covered in the training set of the studied LLMs.
+To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024), which includes programs and faults that were made after the release time of the most LLMs we use (expect DeepSeek-V3-761b only) and thus have limited data leakage risk.
 Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults.
-We hypothesize that in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate.
-On Defects4J, the results are GPT-3.5 (133), GPT-4o (63), GPT-4o-Mini (138), DeepSeek (66), StarChat (13), CodeLlama (28), LEAM (287), Bert (12), and Major (22), respectively.
-On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek (32), StarChat (35), CodeLlama (56), LEAM (56), Bert (5), and Major (92), respectively.
+We hypothesize that, in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate.
+On Defects4J, the results are GPT-3.5 (133), GPT-4o (63), GPT-4o-Mini (138), DeepSeek-V3-761b (72), DeepSeek-V2-236b (66), BugFarm (59), LLMorpheus (26), StarChat (13), CodeLlama (28), LEAM (287), Bert (12), and Major (22), respectively.
+On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek-V3-761b (35), DeepSeek-V2-236b (32), BugFarm (22), LLMorpheus (19), StarChat (35), CodeLlama (56), LEAM (56), Bert (5), and Major (92), respectively.
 The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches.
 Interestingly, Major produces a similar number of exact matches with Codellama.
-Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except from GPTs and LEAM) due to exact matches (in the case of Defects4J).
+Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except for GPTs and LEAM) due to exact matches (in the case of Defects4J).
 Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of exact matches, indicating a minor influence of any data leakage on the reported results.
@@
 The different experimental settings may also threaten the validity of our results.
-To address the threat, we explore the impacts of prompts, context length, few-shot examples, and mutation numbers on the performance of LLMs.
+To address the threat, we explore the impacts of prompt design, prompt turbulence, context length, few-shot examples, and mutation numbers on the performance of LLMs.
 The results show that different settings are highly similar.
@@
 In other words, our study operates under the clean program assumption, which may not fully capture the dynamics of bug detection in practical scenarios.
-However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs,
-The subjective nature of human decisions in labeling for equivalent mutations and non-compilation errors is another potential threat.
+However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs.
+The subjective nature of human decisions in labeling for equivalent mutations, non-compilation errors, and alive mutations is another potential threat.
 To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutation.
@@
 It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test.
-By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978) one can identify weaknesses of their test suites, i.e., mutants that are not killed.
+By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978), one can identify weaknesses of their test suites, i.e., mutants that are not killed.
 Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017).
-Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014b; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
+Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014c; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
 All these tasks require high-utility mutations.
@@
 Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019).
-Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to couple with real faults/bugs to allow these techniques to be effective.
-Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to rule-based.
+Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to be coupled with real faults/bugs to allow these techniques to be effective.
+Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to as rule-based.
 For example, one widely used mutation operator is to change one arithmetic binary operator with another one (e.g., ).
@@
 For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion.
-However, these approaches suffer from the challenges in generating syntactic correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations.
-Perhaps the closest to our work is that of Degiovanni et al.
-who proposed Bert (Degiovanni and Papadakis, 2022).
+LEAM++, the extension of LEAM, further explores learning-based mutation selection (Tian et al., [n.
+d.]).
+However, these approaches suffer from the challenges in generating syntactically correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations.
+Perhaps the closest to our work is that of Degiovanni et al., who proposed BERT (Degiovanni and Papadakis, 2022).
 BERT generates mutation by masking the code elements to be mutated and then predicting the mutation via BERT (Devlin et al., 2018).
@@
 Tip et al.
-concurrently worked on using LLMs for mutation generation for JavaScript, but their approach neither included a comprehensive comparison with existing techniques nor did they utilize only three open-source LLMs (Tip et al., 2024).
+concurrently investigated the use of LLMs for mutation generation in JavaScript (Tip et al., 2025), primarily focusing on the feasibility of leveraging LLMs to produce mutations.
+In contrast, our study conducts a broader comparison by evaluating LLMs against a diverse set of baselines, including both rule-based and learning-based approaches.
+Furthermore, we explore different prompt strategies, such as incorporating few-shot examples and enforcing structured JSON output, and we systematically examine multiple prompt templates in the context of mutation testing.
 Similarly, Endres et al.
@@
 In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches.
-Dakhel et al.(Dakhel et al., 2024) and Ibrahimzada et al.(Ibrahimzada et al., 2023) also proposed generating artificial bugs using LLMs.
+Dakhel et al.
+(Dakhel et al., 2024) and Ibrahimzada et al.
+(Ibrahimzada et al., 2023) also proposed generating artificial bugs using LLMs.
 These approaches mainly evaluate from a certain perspective, such as mutation score or program repair.
+BugFarm (Ibrahimzada et al., 2023) intends to generate artificial bugs that are hard to kill, by first choosing code locations based on a trained small model and then mutating the code via an LLM.
 Compared with these approaches, our study is more comprehensive and evaluates existing mutation generation approaches from many aspects.
@@
 6.2. LLMs for Code
-Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023).
+Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023b).
 Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes.
@@
 Our study conducts extensive and detailed comparative experiments, covering both widely studied rule-based mutation generation approaches and LLM-based methods.
+Particularly, we evaluate not existing LLM-based approaches but also our new prompt for mutation generation.
 Our evaluation results highlight that LLMs significantly outperform traditional approaches in generating diverse mutations that closely mimic real bugs.
 However, the results also expose key limitations of LLMs, particularly their tendency to generate a high proportion of non-compilable mutations.
-Therefore, we analyze the error types of the LLM-generated non-compilable mutations and identify member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
+Therefore, we analyze the error types of the LLM-generated non-compilable mutations and identify that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
 Based on our findings, we emphasize the need for further research to enhance LLM-based mutation testing, addressing its current shortcomings while leveraging its strengths.

### v5

@@
 However, there has been no comprehensive empirical study of their performance for this increasingly important LLM-based Software Engineering application.
-To address this, we conduct a comprehensive empirical study evaluating BugFarm and LLMorpheus (the two state-of-the-art LLM-based approaches), alongside seven LLMs using our newly designed prompt, including both leading open- and closed-source models, on 851 real bugs from two Java real-world bug benchmarks.
-Our results reveal that, compared to existing rule-based approaches, LLMs generate more diverse mutants, that are behaviorally closer to real bugs and, most importantly, with higher fault detection.
-That is, 87.98% (for LLMs) vs.
-41.64% (for rule-based); an increase of percentage points.
+To address this, we conduct a comprehensive empirical study evaluating both a wide range of traditional approaches and LLM-based approaches.
+Particularly, we evaluate BugFarm, LLMorpheus, and our newly designed prompt for mutant generation.
+The experiments cover both leading open- and closed-source LLMs, on 851 real bugs from two Java real-world bug benchmarks (i.e., 605 bugs from 12 projects of Defects4J 2.0 and 246 bugs of ConDefects).
+Our results reveal that, compared to existing traditional approaches, LLMs generate more diverse mutants that are behaviorally closer to real bugs and, most importantly, achieve a 1.75 improvement in real bug detection.
+Specifically, LLM-based approaches reach a detection rate of 76.47%, compared to 44.15% for rule-based techniques, representing an absolute gain of percentage points.
 Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by , , and percentage points, respectively.
-These findings are immediately actionable for both research and practice.
+These findings provide actionable insights for both research and practice.
 They allow practitioners to have greater confidence in deploying LLM-based mutation, while researchers now have a baseline for the state-of-the-art, with which they can research techniques to further improve effectiveness and reduce cost.
 1. Introduction
-For many years, mutation testing has been the topic of scientific research (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019).
+For many years, mutation testing has been the topic of scientific research (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019; Potter et al., 2025).
 Nevertheless, until the last five years, it had made little penetration into the industrial sector.
@@
 In particular, results recently reported by Meta (Harman et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
-This recent trend towards LLM-based mutant generation traces back to work on earlier learning-based approaches, which leverage deep learning models to generate mutations by learning patterns from historical bug fixes (Beller et al., 2021; Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., 2025).
-These methods aim to introduce mutations that better resemble real faults, potentially increasing their relevance for testing.
-However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, to date, no large-scale comprehensive empirical study.
-Without such a study, we have no basis on which to deploy an industrial scale with confidence.
-Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutation generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
+This recent trend towards LLM-based mutant generation traces back to work on earlier learning-based approaches, which leverage deep learning models to generate mutants by learning patterns from historical bug fixes (Beller et al., 2021; Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., 2025).
+These methods aim to introduce mutants that better resemble real faults, potentially increasing their relevance for testing.
+However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, to date, no large-scale comprehensive empirical study exists.
+Without such a study, there is no reliable foundation for confidently deploying the technique at an industrial scale.
+Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutant generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
 The generation of mutants, using any technique, is challenging because of the following three important issues:
-Generating effective mutations: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019; Petrović et al., 2021a).
+Generating effective mutants: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019; Petrović et al., 2021a).
 While existing methods have shown promise, there is still significant room for improvement in enhancing their fault-revealing potential (Degiovanni and Papadakis, 2022).
-Generating valid mutations.: Mutants must be executable against test cases, which allows tests to assess the program’s behavior in the presence of mutations, to judge whether the mutants are killed.
-Ensuring executable mutations is challenging because mutations must remain syntactically and semantically valid, respect type constraints, preserve control and data flow, avoid breaking external dependencies, prevent unintended side effects, and align with test case assumptions to ensure meaningful evaluation (Jia and Harman, 2010).
-Generating diverse and killable mutations: Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021).
-At the same time, mutations need to be diverse and cover the entire spectrum of code and program behaviors.
-This paper examines how well Large-Language Models (LLMs) address these challenges, in comparison to more traditional rule-based approaches.
-We conducted a comprehensive empirical study which studies BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025) (i.e., the state-of-the-art LLM-based mutation generation approaches), alongside seven LLMs adopting our newly designed prompt, covering both leading open- and closed-source models, and four existing mutation testing techniques (i.e., Major (Just et al., 2011), PIT (Coles et al., 2016), LEAM (Tian et al., 2022), and BERT (Degiovanni and Papadakis, 2022)) with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014b) and 246 bugs from ConDefects (Wu et al., 2024)).
-Overall, our primary experiments required the generation of 666,148 mutations.
-Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutation rate, equivalent mutation rate, and efficiency).
-Additionally, we study how different prompt engineering strategies, turbulence on prompt, and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), DeepSeek-V3 (Liu et al., 2024), StarChat--16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task.
-We also explore the potential of using LLMs for generating high-order mutations.
+Generating valid mutants.: Mutants must be executable against test cases, which allows tests to assess the program’s behavior in the presence of mutants, to judge whether the mutants are killed.
+Ensuring executable mutants is challenging because mutants must remain syntactically and semantically valid, respect type constraints, preserve control and data flow, avoid breaking external dependencies, prevent unintended side effects, and align with test case assumptions to ensure meaningful evaluation (Jia and Harman, 2010).
+Generating diverse and killable mutants: Determining whether a mutant is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021).
+At the same time, mutants need to be diverse and cover the entire spectrum of code and program behaviors.
+This paper examines how well Large Language Models (LLMs) address these challenges compared to more traditional rule-based approaches.
+We conducted a comprehensive empirical study that studies BugFarm (Ibrahimzada et al., 2025) and LLMorpheus (Tip et al., 2025) (i.e., the state-of-the-art LLM-based mutant generation approaches).
+Moreover, we propose a novel prompt template for mutant generation, which is equipped with few-shot examples (referred to as LLMut for ease of presentation).
+All three methods are evaluated on two representative LLMs, GPT-4o and DeepSeek-V3-671b (Liu et al., 2024) (DS-671b), denoted as BugFarm (GPT-4o), BugFarm (DS-671b), LLMorpheus (GPT-4o), LLMorpheus (DS-671b), LLMut (GPT-4o), and LLMut (DS-671b), respectively.
+Furthermore, LLMut is additionally evaluated on five other widely used models: GPT-3.5-Turbo (Achiam et al., 2023) (GPT-3.5), GPT-4o-Mini (GPT-4o-M), DeepSeek-Coder-V2-236b (Zhu et al., 2024) (DS-236b), StarChat--16b (Li et al., 2023a) (SC-16b), and CodeLlama-Instruct-13b (Roziere et al., 2023) (CL-13b), to provide a more comprehensive analysis across both open- and closed-source LLMs.
+Other variants of LLMut follow the same notation, e.g., LLMut (GPT-3.5) and LLMut (CL-13b).
+We compare these LLM-based mutant generation approaches with four widely-studied traditional approaches (i.e., Major (Just et al., 2011), PIT (Coles et al., 2016), LEAM (Tian et al., 2022), and BERT (Degiovanni and Papadakis, 2022)), with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014b) and 246 bugs from ConDefects (Wu et al., 2024)).
+Overall, our primary experiments required the generation of more than 701,400 mutants.
+Our evaluation includes both quantitative and qualitative analysis of the mutants generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutant rate, equivalent mutant rate, and efficiency).
+Additionally, we study how different prompt engineering strategies affect the effectiveness of the task.
+We also explore the potential of using LLMs for generating high-order mutants.
 Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
-Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs.
+Our results reveal that, compared to existing approaches, LLMs generate more diverse mutants that are behaviorally closer to real bugs.
 This results in a higher fault detection rate compared to traditional approaches.
-For instance, the mutants generated by DeepSeek-V3-761b and GPT-4o-Mini achieve fault detection rates of 91.1 % and 90.8%, respectively, whereas those generated by LEAM, PIT, and Major yield substantially lower rates of 63.9%, 40.1%, and 66.9%, respectively.
-Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutation rate, and equivalent mutation rate.
-In particular, Major excels with a 97.6% compilability rate, 0% duplicate mutations, and 2.0% equivalent mutants, while GPT-4o exhibits a compilability rate of 76.4%, a duplicate mutation rate of 7.8%, and an equivalent mutation rate of 3.1%, respectively.
+For instance, LLM-based approaches achieve 1.75 improvement in real bug detection rate, with an increase of 32.99 percentage points.
+Specifically, the mutants generated by LLMut (DS-671b) achieve fault detection rates of 91.1%, whereas those generated by LEAM, PIT, and Major yield substantially lower rates of 63.9%, 40.1%, and 66.9%, respectively.
+In terms of coupling rate, LLMorpheus (DS-671b) performs best, with 52% of its mutants coupled with real bugs.
+Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutant rate, and equivalent mutant rate.
+In particular, Major excels with a 97.6% compilability rate, 0% duplicate mutants, and 2.0% equivalent mutants, while GPT-4o exhibits a compilability rate of 76.4%, a duplicate mutant rate of 7.8%, and an equivalent mutant rate of 3.1%, respectively.
 Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types).
-In particular, our results show that the DeepSeek models exhibit the greatest diversity, newly introducing 49 different AST node types.
+In particular, our results show that LLMut (DS-236b) and LLMut (DS-671b) exhibit the greatest diversity, newly introducing 49 different AST node types.
 In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 21.
-The existing LLM-based approach, BugFarm, introduces 31 node types, higher than the traditional approaches as well.
-The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations.
-Moreover, the distributions of AST edit distance of mutations compared to the original code also show that LLMs generate more balanced and expressive mutations.
-For instance, almost 70% of the mutations of Major have an AST edit distance of one, whereas for GPT-4o, only 50% of its mutations have the same distance.
-We also categorise the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall into 9 compilation error types, with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones.
-Moreover, we further analyze the code context of the non-compilable mutations generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features.
-This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, thereby enhancing the effectiveness of mutation testing using AI models.
-We analyze the alive (i.e., survive) mutations of each approach, which are important to strengthen the test case.
-For one alive non-equivalent mutation, it can be (1) not covered, (2) covered but no state infection, or (3) state infection without detection.
-We analyze the distribution of alive mutations across different approaches and find that LLMs based on our prompt are more likely to generate uncovered mutations, which consequently remain alive, providing more valuable guidance for enhancing unit tests.
-Our implementation and the experimental data are available at our repo: https://github.com/da-123-snake/Mutantion-Generation-Study.
+The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutants.
+Moreover, the distributions of AST edit distance of mutants compared to the original code also show that LLMs generate more balanced and expressive mutants.
+For instance, almost 70% of the mutants of Major have an AST edit distance of one, whereas for the LLM-based approach, BugFarm (DS-671b), only 38% of its mutants have the same distance, indicating that LLM-based approaches tend to conduct more complex code transformations.
+We also categorise the mutants generated by GPT models that fail to compile, the most common issue with learning-based approaches generated mutants, and found that they fall into 9 compilation error types, with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones.
+Moreover, we further analyze the code context of the non-compilable mutants generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features.
+This highlights the need for improved strategies to guide model-generated mutants toward more syntactically valid and semantically meaningful variations, thereby enhancing the effectiveness of mutation testing using AI models.
+We analyze the surviving mutants of each approach, which are important to strengthen the test case.
+For one surviving non-equivalent mutant, it can be (1) not covered, (2) covered but no state infection, or (3) state infection without detection.
+We analyze the distribution of surviving mutants across different approaches and find that LLMs based on our prompt are more likely to generate uncovered mutants, which consequently remain surviving, providing more valuable guidance for enhancing unit tests.
 In summary, our paper makes the following main contributions:
 We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods.
-Our study not only covers other existing LLM-based mutation generation approaches, but also introduces a new effective prompt template.
-Our findings indicate that LLMs excel in generating diverse mutations that closely mimic real bugs.
-We analyze the error types of non-compilable mutations and determine that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
-We construct a high-quality dataset of Java mutations, comprehensively annotated with detailed quality and behavior metrics, which can also serve as a valuable resource for fine-tuning in future research.
+Our study not only covers other existing LLM-based mutant generation approaches, but also introduces a novel, effective prompt template.
+Our findings indicate that LLMs excel in generating diverse mutants that closely mimic real bugs.
+We analyze the error types of non-compilable mutants and determine that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutants.
+We construct a high-quality dataset of Java mutants, comprehensively annotated with detailed quality and behavior metrics, which can also serve as a valuable resource for fine-tuning in future research.
 The remainder of this paper is organized as follows.
@@
 2. Study Design
-In this section, we introduce how our study is designed and conducted.
+In this section, we introduce the design and conduct of our study.
 2.1. Overview and Research Questions
-This paper intends to evaluate existing mutation generation approaches by answering the following research questions.
-RQ1 (Effectiveness) How well do mutations generated by LLMs mimic real bugs compared to existing methods?
-This research question investigates the effectiveness of LLM-generated mutations in representing real bugs.
-Specifically, it evaluates whether these mutations exhibit similar characteristics to real faults in terms of detection capability, behavioral similarity, and diversity.
+This paper aims to evaluate existing mutant generation approaches by answering the following research questions.
+RQ1 (Effectiveness) How well do mutants generated by LLMs mimic real bugs compared to existing methods?
+This research question investigates the effectiveness of LLM-generated mutants in representing real bugs.
+Specifically, it evaluates whether these mutants exhibit similar characteristics to real faults in terms of detection capability, behavioral similarity, and diversity.
 This RQ is divided into the following sub-RQs:
-RQ1.1: (Detectability) To what extent do bug-triggering tests successfully detect LLM-generated mutations?
-RQ1.2 (Coupling) To what extent do LLM-generated mutations couple with real bugs?
-RQ1.3 (Diversity) How diverse are LLM-generated mutations compared to traditional approaches?
+RQ1.1: (Detectability) To what extent do bug-triggering tests successfully detect LLM-generated mutants?
+RQ1.2 (Coupling) To what extent do LLM-generated mutants couple with real bugs?
+RQ1.3 (Diversity) How diverse are LLM-generated mutants compared to traditional approaches?
 RQ2 (Validity) How do LLMs compare to existing approaches in terms of validity?
-This research question explores the validity of mutations generated by LLMs compared to traditional approaches.
-Validity is assessed based on key factors such as compilability, duplication, and equivalence of generated mutations.
-By analyzing these aspects, we aim to determine whether LLMs produce valid and meaningful mutations that contribute to practical mutation testing.
+This research question explores the validity of LLM-generated mutants compared to traditional approaches.
+Validity is assessed based on key factors such as compilability, duplication, and equivalence of generated mutants.
+By analyzing these aspects, we aim to determine whether LLMs produce valid and meaningful mutants that contribute to practical mutation testing.
 This RQ is divided into the following sub-RQs:
-RQ2.1 (Compilability) What percentage of generated mutations are compilable across different approaches?
-RQ2.2 (Duplication) What percentage of the generated mutations are duplicates of the original code across different approaches?
-RQ2.3 (Equivalence) What percentage of generated mutations are equivalent across different approaches?
-RQ3 (Efficiency) What are the general costs of different mutation generation approaches?
-This research question investigates the general costs of different mutation generation approaches.
-We consider time and token costs in this RQ, which are associated with the efficiency and practicality of mutation generation approaches.
-RQ4 (Compilation Error) What are the types of compilation errors of the non-compilable mutations of different approaches?
-This RQ aims to categorize and analyze the types of compilation errors that arise from non-compilable mutations generated by different approaches.
-Understanding these errors can guide mutation generation approaches to enhance their practical usability.
-RQ5 (Alive Mutations) What are the distributions of alive mutations across different approaches?
-This RQ aims to investigate the underlying reasons why mutations remain alive, by categorizing them into three cases following the existing study (Just et al., 2014a): (1) uncovered by tests, (2) covered but without infecting the program states, and (3) injecting the states but remaining undetected by tests.
-By examining how these cases are distributed across different approaches, we can better understand the effectiveness of generated mutations and their potential value for guiding test suite improvement.
-2.2. Studied Mutation Generation Approaches
-Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking111https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutation generation approaches are for Java.
+RQ2.1 (Compilability) What percentage of generated mutants are compilable across different approaches?
+RQ2.2 (Duplication) What percentage of the generated mutants are duplicates of the original code across different approaches?
+RQ2.3 (Equivalence) What percentage of generated mutants are equivalent across different approaches?
+RQ3 (Efficiency) What are the general costs of different mutant generation approaches?
+This research question investigates the general costs of different mutant generation approaches.
+We consider time and token costs in this RQ, which are associated with the efficiency and practicality of mutant generation approaches.
+RQ4 (Compilation Error) What are the types of compilation errors of the non-compilable mutants of different approaches?
+This RQ aims to categorize and analyze the types of compilation errors that arise from non-compilable mutants generated by different approaches.
+Understanding these errors can guide mutant generation approaches to enhance their practical usability.
+RQ5 (Surviving Mutants) What are the distributions of surviving mutants across different approaches?
+This RQ aims to investigate the underlying reasons why mutants remain surviving, by categorizing them into three cases following the existing study (Just et al., 2014a): (1) uncovered by tests, (2) covered but without infecting the program states, and (3) injecting the states but remaining undetected by tests.
+By examining how these cases are distributed across different approaches, we can better understand the effectiveness of generated mutants and their potential value for guiding test suite improvement.
+2.2. Studied Mutant Generation Approaches
+Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking111https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutant generation approaches are for Java.
 We categorize these approaches into two groups: traditional approaches (i.e., those developed before the emergence of LLMs) and LLM-based approaches (i.e., approaches that leverage LLMs and are designed for interaction in human-like conversations), as shown below.
 The details of the settings for each generation approach are illustrated in Section 2.5.
-Table 1 summarizes the mutation generation approaches covered by this study.
+Table 1 summarizes the mutant generation approaches covered by this study.
 | Type | Name | Description |
@@
 | LEAM | Custom-trained models learning from a large corpus of real bugs and patches. | |
-| BERT | Mutation generation with the pre-trained model CodeBERT. | |
+| BERT | Mutant generation with the pre-trained model CodeBERT. | |
 | LLM-based | BF (GPT-4o) | The existing LLM-based approach BugFarm with GPT-4o. |
-| LLMph (GPT-4o) | The existing LLM-based mutation generation approach LLMorpheus with GPT-4o. | |
+| BF (DS-671b) | The existing LLM-based approach BugFarm with DeepSeek-V3-671b. | |
+| LLMph (GPT-4o) | The existing LLM-based approach LLMorpheus with GPT-4o. | |
+| LLMph (DS-671b) | The existing LLM-based approach LLMorpheus with DeepSeek-V3-671b. | |
 | LLMut (GPT-3.5) | Our prompt with GPT-3.5-Turbo, trained on data before 2021/09, released in 2021/03. | |
@@
 | LLMut (DS-236b) | Our prompt with DeepSeek-Coder-V2-236b, whose base model is DeepSeek, released in 2024/07. | |
-| LLMut (DS-671b) | Our approach with DeepSeek-V3-761b, whose base model is DeepSeek, released in 2024/12. |
+| LLMut (DS-671b) | Our approach with DeepSeek-V3-671b, whose base model is DeepSeek, released in 2024/12. |
 2.2.1. Traditional Approaches
@@
 PIT: PIT (Coles et al., 2016) is a widely used mutation testing tool on the Java bytecode level, which supports 29 well-designed mutation operators and four groups of active mutation operator sets.
-In our study, we employ PIT 1.14.4 and leverage the ALL group which activates all operators.
-Major: Major (Just et al., 2011) is another high-impact mutation testing framework, that supports nine operators. In our study, we use Major 1.3.5 and involve all mutation operators.
+In our study, we employ PIT 1.14.4 and leverage the ALL group, which activates all operators.
+Major: Major (Just et al., 2011) is another high-impact mutation testing framework that supports nine operators. In our study, we use Major 1.3.5 and involve all mutation operators.
 LEAM: LEAM (Tian et al., 2022, [n.
-d.]) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutations.
-Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutations.
-BERT: BERT (Degiovanni and Papadakis, 2022) is the state-of-the-art mutation generation approaches based on the small-scale pre-trained model, BERT (Devlin et al., 2018).
-Equipped with the pre-trained model, BERT transforms the task of mutation generation to code completion under the surrounding context.
+d.]) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutants.
+Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutants.
+BERT: BERT (Degiovanni and Papadakis, 2022) is the state-of-the-art mutant generation approaches based on the small-scale pre-trained model, BERT (Devlin et al., 2018).
+Equipped with the pre-trained model, BERT transforms the task of mutant generation to code completion under the surrounding context.
 2.2.2. LLM-based Approaches
-Our study incorporates two existing LLM-based mutation generation approaches (i.e., BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025)), and introduces a new prompt design for mutation generation, which will be presented later.
-Existing LLM-based Approaches: We adopt two the state-of-the-art LLM-based mutation generation approaches, BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025).
-BugFarm trains a small-scale model for selecting lines for mutation, and then utilizes LLMs to mutate the target line.
-LLMorpheus performs mutation given the whole method as context.
-Note that LLMorpheus was originally designed for JavaScript.
-Upon examining its prompt, we found it to be language-agnostic.
-Therefore, we reimplemented the code-extraction component for Java (such as parsing ASTs and extracting the code context for mutation) while retaining the original prompt.
-For both approaches, we employ GPT-4o (Achiam et al., 2023), the current most widely used model, as their backend LLM.
-LLMs with Our Prompt: Different from the existing LLM-based approaches, we propose a new prompt engineering for mutation generation, shown as Figure 1.
+Our study incorporates two existing LLM-based mutant generation approaches (i.e., BugFarm (Ibrahimzada et al., 2025) and LLMorpheus (Tip et al., 2025)), and introduces a new prompt design for mutant generation, which will be presented later.
+Existing LLM-based Approaches: We adopt two the state-of-the-art LLM-based mutant generation approaches, BugFarm (Ibrahimzada et al., 2025) and LLMorpheus (Tip et al., 2025).
+BugFarm (Ibrahimzada et al., 2025) trains a small-scale model for selecting lines for mutating, and then utilizes LLMs to mutate the target line.
+Based on the prompt template given in the repository of BugFarm222https://github.com/Intelligent-CAT-Lab/BugFarm, we can find that BugFarm numbers all lines of a Java method and uses a lightweight model to predict the target line numbers to mutate.
+These line numbers are then provided to the LLM.
+BugFarm then instructs the LLM to generate three different mutants for the Java method by changing the specified lines.
+Finally, BugFarm outputs each mutant as a complete mutated method, with the mutation embedded in the full method.
+LLMorpheus (Tip et al., 2025) is an LLM-based mutation testing approach for JavaScript.
+Upon examining its prompt from its repository333https://github.com/githubnext/llmorpheus, we found it to be language-agnostic.
+Therefore, we reimplement the code-extraction component for Java (such as parsing ASTs and extracting the code context for mutants) while retaining its original prompt.
+When porting LLMorpheus to Java, we retain its default configuration.
+LLMorpheus masks specific syntactic elements (e.g., the conditional expression of an if statement) with a PLACEHOLDER to mark the target mutation location, and instructs the LLM to fill in this placeholder.
+For each masked location, LLMorpheus generates three mutants.
+By default, LLMorpheus provides up to 200 lines of surrounding context and instructs the LLM to return only the code fragment that should replace the placeholder.
+In our study, both BugFarm and LLMorpheus are equipped with GPT-4o (Achiam et al., 2023) and DeepSeek-V3-671B (Liu et al., 2024), representing the current leading closed- and open-source LLMs, respectively.
+LLMs with Our New Prompt: Different from the existing LLM-based approaches, we propose a new prompt engineering for mutant generation, shown as Figure 1.
 For ease of presentation, in the following part of this paper, we call our prompt LLMut.
-To craft effective prompts, we follow the best practices, which suggest that prompts should comprise four aspects: Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024).
+To craft effective prompts, we follow the best practices which suggest that prompts should comprise four aspects: Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024).
 In the Instructions, we direct the LLMs to generate mutants for the target code element.
-The placeholder MUT_NUMBER specifies the number of mutations required, while the placeholder CODE_ELEMENT specifies the target code snippet to be mutated.
+The placeholder MUT_NUMBER specifies the number of mutants required, while the placeholder CODE_ELEMENT specifies the target code snippet to be mutated.
 In the Context, we clearly state that a mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark.
 In the Input Data, we specify the code element to be mutated and the number of mutants to generate.
-In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments.
+In the Output Indicator, we specify the JSON file format for outputs, facilitating further experiments.
 In addition, our prompts need few-shot examples, which should originate from real bugs, enabling LLMs to learn how to mutate code from real bugs.
-We further explore the different prompt settings in Section 4.1 and the impact of prompt turbulence in Section 4.2.
+We further explore the different prompt settings in Section 4.1.
 To avoid the few-shot examples leaking information, beyond the evaluation datasets (i.e., as shown later, they are Defects4J and ConDefects), we employ another benchmark, QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b).
-Referring to the study by Deng et al.
-(Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples.
+As shown later, our experimental results indicate that providing six examples results in the best overall performance, consistent with observations by Deng et al.
+(Deng et al., 2024).
 Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt.
 To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected.
-If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 2.
+If it is different, we add it to our collection and continue until we have collected all examples, shown as Table 2.
 As LLM capabilities continue to advance, the optimal number of few-shot examples may vary.
 We therefore investigate the impact of different numbers of few-shot examples in Section 4.1.3.
+Note that adopting few-shot examples may also have drawbacks, which we will discuss later.
 | Correct Version | Buggy Version |
@@
 | while(Math.abs(x-approx*approx)>epsilon) | while(Math.abs(x-approx)>epsilon) |
-To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016).
-We use PIT-generated mutations with the universal set of operators and find that PIT generates 1.15 mutations per line of code.
-Therefore, based on the given code context, we generate one mutation per line.
-We explore the influence of context length on mutation generation in Section 4.
+To determine the number of mutants generated in a prompt, we adopt the setting used in PIT (Coles et al., 2016).
+We use PIT-generated mutants with the universal set of operators and find that PIT generates 1.15 mutants per line of code.
+Therefore, based on the given code context, we generate one mutant per line.
+We explore the influence of context length on mutant generation in Section 4.
 Additionally, for the temperature, we use the default value for each LLM.
-Compared with the prompts used in BugFarm222https://github.com/Intelligent-CAT-Lab/BugFarm and LLMorpheus333https://github.com/githubnext/llmorpheus, our prompt design offers several advantages:
-it incorporates real-world bugs as few-shot examples, which have been shown to significantly enhance LLM performance in many code-related tasks;
-it determines the number of generated mutants based on the size of the target code, rather than on a fixed number;
-it specifies the mutation location using the target statement itself rather than line numbers, enabling more precise localization (Fan et al., 2023a); and
-it employs JSON output formatting, which is widely used in LLM fine-tuning and downstream tasks.
+Each LLM-based approach in our study adopts a different prompt-engineering strategy.
+For example, LLMut incorporates few-shot examples and uses a structured JSON output format, whereas LLMorpheus masks high-value code fragments and instructs the LLM to fill in the masked content to ensure the quality of the output mutants.
+Note that the different prompt-engineering strategies are complementary, with different approaches emphasizing different aspects of the mutants.
 Based on LLMut, our study comprehensively covers the most widely studied LLMs.
 Specifically, we select a representative set of LLMs with code capabilities, both from commercial closed-source and open-source models.
-Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5 (GPT-3.5-Turbo) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-M (GPT-4o-Mini) (Achiam et al., 2023)), and open-source LLMs, including SC-16b (StarChat--16b), CL-13b (CodeLlama-Instruct-13b), DS-236b (DeepSeek-Coder-V2-236b), and DS-671b (DeepSeek-V3-671b).
+Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5-Turbo (GPT-3.5) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-Mini (GPT-4o-M) (Achiam et al., 2023)), and open-source LLMs, including StarChat--16b (SC-16b), CodeLlama-Instruct-13b (CL-13b), DeepSeek-Coder-V2-236b (DS-236b), and DeepSeek-V3-671b (DS-671b).
 The details of these models are shown as Table 1.
@@
 Moreover, the selected LLMs represent a balance between general-purpose models (e.g., GPT-4o and DeepSeek-V3-671b), which are trained for a wide range of natural language and reasoning tasks, and code-specialized models (e.g., CodeLlama-Instruct-13b and DeepSeek-V2-Coder-236b), which are further fine-tuned on large-scale code corpora.
-This selection allows us to evaluate mutation generation across both broadly capable models and those explicitly optimized for programming tasks, ensuring the conclusions are representative of the current landscape of LLMs.
+This selection allows us to evaluate mutant generation across both broadly capable models and those explicitly optimized for programming tasks, ensuring the conclusions are representative of the current landscape of LLMs.
 2.3. Datasets
@@
 | Total | — | 851 | 2006/07/16 - 2024/06/30 |
-As one of the major tasks of our study is to evaluate the behavioral similarity between mutations and real-world bugs, we selected datasets that include real-world bugs along with their corresponding fault-revealing test cases.
+As one of the major tasks of our study is to evaluate the behavioral similarity between mutants and real-world bugs, we selected datasets that include real-world bugs along with their corresponding fault-revealing test cases.
 Specifically, we employ the datasets with the following properties:
 The datasets should comprise Java programs, as existing methods are primarily based on Java, and we need to compare with them.
-The bugs of the datasets should be real-world bugs so that we can compare the difference between mutations and real bugs.
+The bugs of the datasets should be real-world bugs so that we can compare the difference between mutants and real bugs.
 Every bug in datasets has the correctly fixed version provided by developers so that we can mutate the fixed version and compare them with the corresponding real bugs.
-Every bug is accompanied by at least one bug-triggering test because we need to measure whether the mutations affect the execution of bug-triggering tests.
-To this end, we employ the Defects4J 2.0 (Just et al., 2014b) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 3.
+Every bug is accompanied by at least one bug-triggering test because we need to measure whether the mutants affect the execution of bug-triggering tests.
+To this end, we employ the Defects4J 2.0 (Just et al., 2014b) and ConDefects (Wu et al., 2024) to evaluate the mutant generation approaches, shown as Table 3.
 In total, we conducted experiments on 851 bugs.
@@
 To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs.
-Among them, only the most recent model, DeepSeek-V3-731b, cannot guarantee data leakage-free, since it is released after the ConDefects dataset.
+Among them, only the most recent model, DeepSeek-V3-731b, cannot guarantee data leakage-free, since it was released after the ConDefects dataset.
 2.4. Evaluation Metrics
-In this subsection, we present the evaluation metrics used in our study to assess the quality of the generated mutations comprehensively.
-We evaluate the mutations from multiple perspectives, as mentioned before.
+In this subsection, we present the evaluation metrics used in our study to comprehensively assess the quality of the generated mutants.
+We evaluate the mutants from multiple perspectives, as mentioned before.
 Note that only relying on the mutation score is insufficient for a meaningful assessment.
-A higher mutation score does not necessarily indicate better quality mutations, as it only reflects the proportion of killed mutants rather than their effectiveness.
-However, we still report the mutation score of each mutation generation approach later in Section 3.
-To better understand the metrics, we first illustrate the categories of mutations.
-Let all generated mutations be the set , where LLMs may generate syntactically incorrect mutations that can not pass compilation.
-Thus, we denote the compilable mutations as the set .
-Note that the set contains two types of mutations that should be removed, the duplicate mutations (denoted as ) and equivalent mutations (denoted as ).
+A higher mutation score does not necessarily indicate better quality mutants, as it only reflects the proportion of killed mutants rather than their effectiveness.
+However, we still report the mutation score of each mutant generation approach later in Section 3.
+To better understand the metrics, we first illustrate the categories of mutants.
+Let all generated mutants be the set , where LLMs may generate syntactically incorrect mutants that can not pass compilation.
+Thus, we denote the compilable mutants as the set .
+Note that the set contains two types of mutants that should be removed, the duplicate mutants (denoted as ) and equivalent mutants (denoted as ).
 Therefore, we have .
-The duplicate mutations refer to those identical to the original code or other mutations.
-In mutation testing, the ideal set of mutations to employ is from the set , excluding equivalent mutations that have no contribution to discovering the weakness of tests.
-However, determining the equivalent mutations set is undecidable, so we follow the common practice and use the set as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025).
+The duplicate mutants refer to those identical to the original code or other mutants.
+Ideally, we should perform mutation testing against the mutant set of , excluding all duplicated mutants () and equivalent mutants (), which have no contribution to discovering the weakness of tests.
+Based on our definition, the duplicated mutants are syntactically identical mutants that are trivially decidable.
+However, determining the equivalent mutants set is undecidable, and is negligible compared to the universal mutant set (i.e., ).
+Therefore, we use the set of mutants as an approximation to compute the mutation score (i.e., ), which is the common practice of existing studies (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025a).
 Therefore, let denote the set of killed mutants among the compilable and non-duplicate ones (i.e., ).
 The mutation score is then computed as follows:
-Note that traditional mutation generation approaches, such as Major and PIT, produce very few duplicate mutations (i.e., is 0).
-Hence, early studies (Kintis et al., 2017) defined equivalent mutations as those semantically equivalent to the original code, which may implicitly include , which is an empty set.
+Note that traditional mutant generation approaches, such as Major and PIT, produce very few duplicate mutants (i.e., is 0).
+Hence, early studies (Kintis et al., 2017) defined equivalent mutants as those semantically equivalent to the original code, which may implicitly include , which is an empty set.
 However, LLM-based approaches tend to generate a large number of non-mutated duplicates.
-Moreover, while duplicate mutations can be easily filtered, identifying equivalent mutations remains challenging (Tian et al., 2024).
-Therefore, in this paper, we refine the definition of equivalent mutations to be those that are syntactically different but semantically equivalent to the original code.
+Moreover, while duplicate mutants can be easily filtered, identifying equivalent mutants remains challenging (Tian et al., 2024).
+Therefore, in this paper, we refine the definition of equivalent mutants to be those that are syntactically different but semantically equivalent to the original code.
 2.4.1. RQ1: Effectiveness
-Evaluating the effectiveness of mutation generation approaches requires assessing how well the generated mutations mimic real bugs.
-A key aspect of this evaluation is understanding whether LLM-generated mutations exhibit fault characteristics that align with real-world defects.
-To quantify this effectiveness, we employ multiple metrics that capture different aspects of behavioral similarity between mutations and real bugs.
-These metrics assess whether bug-triggering tests can detect LLM-generated mutations (i.e., RQ1.1), how closely mutations resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutation generation techniques (i.e., RQ1.3).
+Evaluating the effectiveness of mutant generation approaches requires assessing how well the generated mutants mimic real bugs.
+A key aspect of this evaluation is understanding whether LLM-generated mutants exhibit fault characteristics that align with real-world defects.
+To quantify this effectiveness, we employ multiple metrics that capture different aspects of behavioral similarity between mutants and real bugs.
+These metrics assess whether bug-triggering tests can detect LLM-generated mutants (i.e., RQ1.1), how closely mutants resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutant generation techniques (i.e., RQ1.3).
 The metrics of these sub-RQs are listed below.
 RQ1.1: Real Bug Detection Rate: Following the approach of Just et al.
-(Just et al., 2014c), we evaluate the effectiveness of mutation generation by measuring its ability to detect real bugs using developer-written tests.
-For a given real bug (which may span multiple lines), we apply a mutation generation approach to create a set of mutations on the fixed version of the program.
-If the test cases that kill a mutation also fail on the real bug, we consider that the mutation generation approach has successfully “detected” this real bug.
-To quantify this effectiveness, we compute the Real Bug Detection Rate, which represents the proportion of real bugs for which at least one generated mutation is detected by the same test cases that reveal the real bug.
-This metric provides insight into how well LLM-generated mutations align with actual software defects in a testing context.
-RQ1.2: Average Ochiai Coefficient: This metric estimates the semantic similarity between a mutation and its corresponding real bug based on their passing and failing test cases (Offutt and Untch, 2001).
+(Just et al., 2014c), we evaluate the effectiveness of mutant generation by measuring its ability to detect real bugs using developer-written tests.
+For a given real bug (which may span multiple lines), we use a mutant-generation approach to create a set of mutants on the fixed version of the program.
+If the test cases that kill a mutant also fail on the real bug, we consider that the mutant generation approach has successfully “detected” this real bug.
+To quantify this effectiveness, we compute the Real Bug Detection Rate, which represents the proportion of real bugs for which at least one generated mutant is detected by the same test cases that reveal the real bug.
+In other words, it measures whether the generated mutants can reproduce the failing behaviors of a set of real bugs under tests.
+This metric provides insight into how well LLM-generated mutants align with actual software defects in a testing context.
+RQ1.2: Average Ochiai Coefficient: This metric estimates the semantic similarity between a mutant and its corresponding real bug based on their passing and failing test cases (Offutt and Untch, 2001).
 Ochiai coefficient measures how a mutation resembles its corresponding real bug.
 The Ochiai coefficient was originally introduced as a general similarity measure but later became widely adopted in spectrum-based fault localization (SBFL) (Zou et al., 2019).
-In this study, we use the Ochiai coefficient in its original form to assess how closely a mutation resembles its corresponding real bug, which is consistent with the study of Ojdanic et al.
+In this study, we use the Ochiai coefficient in its original form to assess how closely a mutant resembles its corresponding real bug, which is consistent with the findings of Ojdanic et al.
 (Ojdanic et al., 2023).
-More concretely, let and be a mutation and its corresponding real bug, and let and be the set of their failing test cases, respectively. The coefficient of is computed by the following formula:
+More concretely, let and be a mutant and its corresponding real bug, and let and be the set of their failing test cases, respectively. The coefficient of is computed by the following formula:
 Therefore, if the Ochiai coefficient of is , it is semantically equal to its corresponding bug .
-Given a bug , its Ochiai coefficient is calculated as the mean Ochiai coefficient across all its corresponding de-duplicated mutations (i.e., ), shown as the following formula:
-Given a mutation generation approach and a set of bugs , the Average Ochiai Coefficient of the approach is calculated as the mean Ochiai coefficient across all bugs, shown as the following formula:
-RQ1.2: High Ochiai Bug Count: Following the study (Ojdanic et al., 2023), a bug and a set of mutations are considered to have high Ochiai if they satisfy .
-This threshold indicates a strong semantic similarity between the generated mutations and the real bug.
-For each mutation generation approach, we compute the number of bugs that meet this criterion.
+Given a bug , the Ochiai coefficient for is calculated as the mean Ochiai coefficient across all its corresponding de-duplicated mutants (i.e., ), shown as the following formula:
+Given a mutant generation approach and a set of bugs , the Average Ochiai Coefficient of the approach is calculated as the mean Ochiai coefficient across all bugs, shown as the following formula:
+RQ1.2: High Ochiai Bug Count: Following the study (Ojdanic et al., 2023), a bug and a set of mutants are considered to have high Ochiai if they satisfy .
+This threshold indicates a strong semantic similarity between the generated mutants and the real bug.
+For each mutant generation approach, we compute the number of bugs that meet this criterion.
 RQ1.2: Coupling Rate: Given a real bug, one mutant is considered coupling if the set of test cases that kill the mutant overlaps with the bug-triggering tests.
 Note that each bug we analyzed has at least one failing test to reveal the bug, called bug-triggering test.
-The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutations (i.e., ).
-Note that Real Bug Detection Rate and Coupling Rate measure effectiveness from different perspectives.
-Real Bug Detection Rate evaluates effectiveness at the bug level, i.e., the proportion of bugs that can be revealed by at least one mutant.
-In contrast, Coupling Rate evaluates effectiveness at the mutation level, i.e., the proportion of mutants whose failing tests overlap with bug-triggering tests.
-For instance, a mutation approach may generate a large number of high-quality mutants for only a few specific bugs, but few mutants for others.
-This can result in a high Coupling Rate (since the generated mutants are strongly aligned with the bug-triggering tests of those few bugs) but a low Real Bug Detection Rate (since many other bugs remain undetected).
-Conversely, another approach may generate mutants that cover a wide range of bugs, thereby achieving a high Real Bug Detection Rate.
-However, if the mutants for each bug exhibit only weak coupling with the bug-triggering tests, the overall Coupling Rate will remain low.
-Therefore, these two metrics are complementary, with each reflecting only one aspect of mutation effectiveness.
-RQ1.2: Coupling Categorization: We categorize each generated mutation into distinct coupling scales based on their similarity to real bugs, following the classification proposed by (Gay and Salahirad, 2023), and analyze their distribution across different approaches.
-Specifically, given a compilable mutation (i.e., an element from the set ) on the fixed version of a real bug (which can be uncovered by at least one triggering test), we can classify it into the following six scales based on its behavior:
-Strong Substitution: All triggering tests fail and no additional tests fail, which represents the mutation is a semantic replacement of the real bug under the given test suite.
-Strong Substitution + Additional Tests Fail: All triggering tests fail but additional non-triggering tests fail.
+The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutants (i.e., ).
+Note that Real Bug Detection Rate and Coupling Rate assess effectiveness from different granularities.
+Real Bug Detection Rate evaluates effectiveness at the bug level—the proportion of bugs for which at least one generated mutant reproduces the failing behavior.
+Coupling Rate evaluates effectiveness at the mutant level—the proportion of mutants whose failing tests overlap with the bug-triggering tests.
+These two metrics capture different phenomena.
+For example, an approach may generate many high-quality mutants for only a few bugs while producing very few mutants for the remaining bugs.
+In this case, the Coupling Rate can be high (because many mutants align well with the few covered bugs), but the Real Bug Detection Rate remains low (because many bugs never receive any useful mutants).
+Conversely, an approach might generate mutants that touch a wide range of bugs, achieving a high Real Bug Detection Rate.
+However, if per-bug mutants only weakly reflect the bug-triggering behavior, the overall Coupling Rate may remain low.
+Therefore, the two metrics are complementary: Real Bug Detection Rate measures coverage across bugs, while Coupling Rate measures behavioral precision within mutants.
+Neither metric alone can fully represent the effectiveness of mutation generation.
+RQ1.2: Coupling Categorization: We categorize each generated mutant into distinct coupling scales based on their similarity to real bugs, following the classification proposed by (Gay and Salahirad, 2023), and analyze their distribution across different approaches.
+Specifically, given a compilable mutant (i.e., an element from the set ) on the fixed version of a real bug (which can be uncovered by at least one triggering test), we can classify it into the following six scales based on its behavior:
+Strong Substitution: All triggering tests fail and no additional tests fail, which represents the mutant is a semantic replacement of the real bug under the given test suite.
+Strong Substitution + Additional Tests Fail: All triggering tests fail, but additional non-triggering tests fail.
 Partial Substitution: Some triggering tests fail and no additional tests fail.
-Partial Substitution + Additional Tests Fail: Some triggering tests fail and some additional non-triggering tests fail.
+Partial Substitution + Additional Tests Fail: Some triggering tests fail, and some additional non-triggering tests fail.
 Not Substitution: Only non-triggering tests fail.
-Not Detected: No tests fail (i.e., the mutation is alive under the given test suite).
-Later, we will classify the mutations of each approach and analyze their distributions.
-RQ1.3: AST Node Diversity: We prefer the set of mutations with diverse syntactic forms, i.e., method invocations, field accesses, etc.
+Not Detected: No tests fail (i.e., the mutant is surviving under the given test suite).
+Later, we will classify the mutants of each approach and analyze their distributions.
+RQ1.3: AST Node Diversity: We prefer the set of mutants with diverse syntactic forms, i.e., method invocations, field accesses, etc.
 Therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes.
@@
 The AST Edit Distance metric is well-suited for structured code, as it captures syntactic differences at the tree level rather than merely comparing token sequences, as in BLEU and similar metrics.
-However, it still has inherent limitations, such as its inability to fully reflect the impact of mutations on program behavior.
+However, it still has inherent limitations, such as its inability to fully reflect the impact of mutants on program behavior.
 We further discuss these limitations later in Section 5.
 2.4.2. RQ2: Validity
-We consider three key metrics to evaluate the validity of different mutation generation approaches: generation efficiency, compilability, and the proportion of duplicate and equivalent mutations, providing a basic understanding of the validity of each approach.
-RQ2.1: Compilability Rate: the proportion of mutations that successfully compile, which is computed as:
+We consider three key metrics to evaluate the validity of different mutant generation approaches: generation efficiency, compilability, and the proportion of duplicate and equivalent mutants, providing a basic understanding of the validity of each approach.
+RQ2.1: Compilability Rate: the proportion of mutants that successfully compile, which is computed as:
 Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher compilability rate.
-RQ2.2: Duplicate Mutation Rate: the proportion of duplicate mutations, computed as:
-LLMs sometimes generate duplicate mutations, i.e., mutations that are either identical to the original code or identical to each other.
-While traditional approaches rarely generate duplicate mutations, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data.
-After eliminating duplicate mutations, all remaining mutations are guaranteed to be syntactically distinct both from the original code and from each other.
-RQ2.3: Equivalent Mutation Rate: the proportion of equivalent mutation, computed as:
-Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score.
+RQ2.2: Duplicate Mutant Rate: the proportion of duplicate mutants, computed as:
+LLMs sometimes generate duplicate mutants, i.e., mutants that are either identical to the original code or identical to each other.
+While traditional approaches rarely generate duplicate mutants, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data.
+After eliminating duplicate mutants, all remaining mutants are guaranteed to be syntactically distinct both from the original code and from each other.
+RQ2.3: Equivalent Mutant Rate: the proportion of equivalent mutants, computed as:
+Equivalent mutants are semantically equivalent to the original program, which impacts the accuracy of the mutation score.
 We prefer the approach with a lower EMR.
-Note that while equivalent mutations are also considered semantically duplicate, they are not easily identifiable.
-Therefore, we distinguish them from other duplicate mutations, which can be filtered syntactically.
+Note that while equivalent mutants are also considered semantically duplicate, they are not easily identifiable.
+Therefore, we distinguish them from other duplicate mutants, which can be filtered syntactically.
 2.4.3. RQ3: Efficiency
-Efficiency describes how fast a mutation approach produces mutations.
-Mutation testing requires numerous mutations to evaluate the robustness of a test suite effectively.
-Therefore, the efficiency of mutation generation directly impacts the feasibility and scalability of mutation testing.
+Efficiency describes how fast a mutation approach produces mutants.
+Mutation testing requires numerous mutants to evaluate the robustness of a test suite effectively.
+Therefore, the efficiency of mutant generation directly impacts the feasibility and scalability of mutation testing.
 To assess this, we compare these approaches based on time costs.
-RQ3.1: Average Generation Time: measures the average time required to generate each mutation, providing a quantitative evaluation of the efficiency of each approach.
-RQ3.2: Average Token: measures the average number of consumed tokens per mutation, reflecting both the financial cost (due to token-based pricing) and the computational time cost (since longer token sequences typically require more processing time).
+RQ3.1: Average Generation Time: measures the average time required to generate each mutant, providing a quantitative evaluation of the efficiency of each approach.
+RQ3.2: Average Token: measures the average number of consumed tokens per mutant, reflecting both the financial cost (due to token-based pricing) and the computational time cost (since longer token sequences typically require more processing time).
 Since input and output token costs differ across models, we calculate them separately.
-2.5. Mutation Generation Setups
-In this subsection, we discuss the setups for mutation generation approaches, especially the setups for LLM-based generation.
-For the well-established mutation generation approaches (i.e., PIT, Major, LEAM, and BERT), we activate all their mutation operators.
-Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutations within the context of real-world bugs.
-Note that we supply the same code snippets as context across all mutation generation approaches to ensure a fair comparison.
+2.5. Mutant Generation Setups
+In this subsection, we discuss the setups for mutant generation approaches, especially the setups for LLM-based generation.
+For the well-established mutant generation approaches (i.e., PIT, Major, LEAM, and BERT), we activate all their mutation operators.
+Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutants within the context of real-world bugs.
+Note that we supply the same code snippets as context across all mutant generation approaches to ensure a fair comparison.
 For all LLM-based approaches, we use the default settings of each model (e.g., token limits and temperature).
@@
 Table 4 presents the overall performance of each generation approach.
-In terms of the number of mutations generated (Mutation Count), PIT produces the largest number of mutations at 340,728, significantly more than other approaches, while the remaining approaches range from about twenty thousand to thirty thousand.
-In terms of mutation score, the traditional rule-based approaches (PIT and Major) are over 0.5, LEAM and BERT are over 0.6, and the LLMs from the GPT family are all over 0.7.
-Note that the number of generated mutations and mutation scores are not quality metrics for evaluating these approaches.
-A large number of mutations may include redundant or equivalent mutants that do not contribute to mutation scores, while a high mutation score may be influenced by the ease with which mutants are detected rather than their resemblance to real faults.
+In terms of the number of mutants generated (Mutant Count), PIT produces the largest number of mutants at 340,728, significantly more than other approaches, while the remaining approaches range from about twenty thousand to thirty thousand.
+In terms of mutation score, the traditional rule-based approaches (PIT and Major) are over 0.5, LEAM and BERT are over 0.6, while most LLM-based approaches are over 0.7.
+Particularly, BugFarm (DS-671b) achieves the highest mutation score, which is nearly 0.79.
+Note that the number of generated mutants and mutation scores are not quality metrics for evaluating these approaches.
+A large number of mutants may include redundant or equivalent mutants that do not contribute to mutation scores, while a high mutation score may be influenced by the ease with which mutants are detected rather than their resemblance to real faults.
 | Name | #Mut. | Mut. S. | R. B. Detec. | Avg. Ochiai | Coup. Rate | Comp. Rate | Dup. Rate | Time |
@@
 | BF (GPT-4o) | 13068 | 0.762 | 58.6% | 17.7% | 47.9% | 74.1% | 7.5% | 5.35 |
+| BF (DS-671b) | 13896 | 0.787 | 58.9% | 17.9% | 48.0% | 77.5% | 4.5% | 7.29 |
 | LLMph (GPT-4o) | 20043 | 0.754 | 68.5% | 24.1% | 51.7% | 67.8% | 7.1% | 3.41 |
+| LLMph (DS-671b) | 21356 | 0.759 | 69.3% | 26.0% | 52.0% | 68.1% | 6.9% | 5.44 |
 | LLMut (GPT-3.5) | 31656 | 0.702 | 87.0% | 47.7% | 46.9% | 62.5% | 10.3% | 1.79 |
@@
 | LLMut (GPT-4o-M) | 34417 | 0.708 | 90.8% | 54.1% | 49.5% | 75.3% | 7.0% | 1.65 |
-| LLMut (DS-761b) | 33596 | 0.749 | 91.1% | 59.8% | 50.8% | 77.6% | 7.5% | 2.57 |
+| LLMut (DS-671b) | 33596 | 0.749 | 91.1% | 59.8% | 50.8% | 77.6% | 7.5% | 2.57 |
 | LLMut (DS-236b) | 34741 | 0.714 | 89.4% | 50.9% | 49.5% | 75.8% | 7.7% | 4.25 |
@@
 BF: BugFarm. LLMph: LLMorpheus. LLMut: Our new prompt.
-GPT-4o-M: GPT-4o-Mini. DS-761b: DeepSeek-V3-761b. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat--16b. CL-13b: CodeLlama-Instruct-13b.
+GPT-4o-M: GPT-4o-Mini. DS-671b: DeepSeek-V3-671b. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat--16b. CL-13b: CodeLlama-Instruct-13b.
 3.1. RQ1: Performance on Effectiveness
-To measure the effectiveness of mutation generation approaches, we evaluate them with three sub-RQs, which estimate their effectiveness via real bug detectability, coupling, and diversity, respectively.
+To measure the effectiveness of mutant generation approaches, we evaluate them with three sub-RQs, which estimate their effectiveness via real bug detectability, coupling, and diversity, respectively.
 3.1.1. RQ1.1 Detectability
 As illustrated in Section 2.4.1, we estimate real bug detectability via the following metric.
-Real Bug Detection Rate: The column Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutation generation approaches.
+Real Bug Detection Rate: The column Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutant generation approaches.
 Among the traditional approaches, PIT achieves a detection rate of 40.1%, while Major, LEAM, and BERT perform significantly better with a detection rate of 66.9%, 63.9%, and 60.5%, respectively.
 For LLM-based approaches, the results are significantly more effective than those of traditional ones.
-The existing LLM-based approaches, BugFarm and LLMorpheus, achieve a rate of 58.6% and 68.5%, respectively.
-For the LLMs based on our prompting, DeepSeek-V3-761b achieves the highest detection rate at 91.1%, closely followed by GPT-4o-Mini, GPT-4o and DeepSeek-Coder-V2-236b, with detection rates of 90.8%, 89.9%, and 89.4%, respectively.
-Meanwhile, GPT-3.5 and CodeLlama achieve moderate detection rates of 87.0% and 71.2%, respectively.
-| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Gs. | Csv | Ja. | Js. | ConD. | All |
+Across all models, LLMorpheus and LLMut constantly outperform all traditional approaches.
+For example, LLMut (DS-671b) achieves the highest detection rate at 91.1%, closely followed by LLMut (GPT-4o-M) and LLMut (GPT-4o), with detection rates of 90.8% and 89.9%, respectively.
+Moreover, we can find that all DS-671b outperforms GPT-4o in terms of bug detection rate, across all three LLM-based approaches.
+| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Csv | Gs. | Ja. | Js. | ConD. | All |
 | #Bug | 26 | 65 | 27 | 106 | 38 | 133 | 39 | 18 | 16 | 18 | 26 | 93 | 246 | 851 |
@@
 | BF (GPT-4o) | 14 | 37 | 15 | 57 | 28 | 93 | 23 | 6 | 8 | 11 | 7 | 83 | 117 | 499 |
+| BF (DS-671b) | 16 | 37 | 15 | 49 | 25 | 100 | 22 | 7 | 8 | 9 | 7 | 88 | 118 | 501 |
 | LLMph (GPT-4o) | 17 | 51 | 20 | 61 | 31 | 121 | 30 | 12 | 8 | 12 | 9 | 87 | 124 | 583 |
+| LLMph (DS-671b) | 17 | 50 | 20 | 63 | 33 | 123 | 29 | 9 | 10 | 13 | 9 | 86 | 128 | 590 |
 | LLMut (GPT-3.5) | 22 | 52 | 25 | 92 | 30 | 130 | 37 | 14 | 15 | 13 | 23 | 93 | 194 | 740 |
@@
 | LLMut (GPT-4o-M) | 23 | 56 | 26 | 96 | 36 | 131 | 39 | 17 | 16 | 16 | 21 | 93 | 203 | 773 |
-| LLMut (DS-761b) | 24 | 61 | 27 | 101 | 35 | 131 | 39 | 13 | 16 | 16 | 25 | 92 | 195 | 775 |
+| LLMut (DS-671b) | 24 | 61 | 27 | 101 | 35 | 131 | 39 | 13 | 16 | 16 | 25 | 92 | 195 | 775 |
 | LLMut (DS-236b) | 23 | 57 | 22 | 95 | 34 | 132 | 39 | 17 | 16 | 15 | 21 | 90 | 200 | 761 |
@@
 For individual projects, LLM-based approaches consistently achieve top performance across multiple projects.
-Notably, DeepSeek-V3-761b detects the largest number of bugs in most Defects4J projects and achieves the highest overall bug detection count.
-Similarly, DeepSeek-Coder-V2-236b, GPT-4o, and GPT-4o-Mini demonstrate strong performance, achieving the best results in projects such as Mockio, Csv, and ConDefects.
+Notably, LLMut (DS-671b) detects the largest number of bugs in most Defects4J projects and achieves the highest overall bug detection count.
+Similarly, LLMut (DS-236b), LLMut (GPT-4o), and LLMut (GPT-4o-M) demonstrate strong performance, achieving the best results in projects such as Mockio, Csv, and ConDefects.
 3.1.2. RQ1.2 Coupling
@@
 Average Ochiai Coefficient: The column Avg.
-Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measures the semantic distance between mutations and real bugs.
+Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measure the semantic distance between mutants and real bugs.
 Among the traditional approaches, Major achieves an Ochiai coefficient of 35.5%, higher than the others.
 LEAM and BERT achieve coefficients of 32.4% and 29.3%, respectively, indicating that the best-performing rule-based approach, Major, outperforms both of them.
-Existing other LLM-based approaches achieve an overall lower Ochiai coefficient of 17.7% and 24.1%, respectively.
-The LLM-based approaches using our prompt achieve notably higher Ochiai coefficients.
-DeepSeek-V3-761b attains the highest value at 59.8%, followed by GPT-4o at 56.2% and GPT-4o-Mini at 54.1%.
-The LLMs based on our prompt outperform other existing methods, not only traditional approaches but also existing LLM-based ones, showing their superiority in generating mutations that are semantically closer to real bugs.
+The average Ochiai coefficients of LLM-based approaches vary significantly, ranging from 17.7% to 59.8%, indicating diverse semantic distances of their mutants.
+Among them, LLMut (DS-671b) achieves the highest value at 59.8%, followed by LLMut (GPT-4o) at 56.2% and LLMut (GPT-4o-M) at 54.1%, indicating that these mutants capture fault-triggering behaviors more closely aligned with real bugs compared to traditional approaches.
 High Ochiai Bug Count: Following the study of Ojdanic et al.
-(Ojdanic et al., 2023), the mutation set with an Ochiai coefficient over 0.8 is considered to be of high semantic similarity with its corresponding real bug.
-For each project, we count the number of bugs that achieved a high Ochiai coefficient by each approach, shown as Table 6.
-The overall trend is similar to Real Bug Detectability, which indicates that the LLM DeepSeek-V3-761b, based on our prompt, achieves the highest number of bugs across most projects.
-The closed-source LLMs GPT-4o and GPT-4o-Mini also exhibit strong performance, closely following DeepSeek-V3-761b in many cases.
+(Ojdanic et al., 2023), the mutant set with an Ochiai coefficient over 0.8 is considered to be of high semantic similarity with its corresponding real bug.
+For each project, we count the number of bugs that achieved a high Ochiai coefficient using each approach, as shown in Table 6.
+The overall trend is similar to Real Bug Detectability, which indicates that the LLMut (DS-671b) achieves the highest number of bugs across most projects.
+Moreover, LLMut (GPT-4o) and LLMut (GPT-4o-M) also exhibit strong performance, closely following LLMut (DS-671b) in many cases.
 Compared with our approach, all traditional approaches exhibit significantly fewer high Ochiai bugs.
-Interestingly, among traditional approaches, Major outperforms LEAM and BERT across most projects.
-| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Gs. | Csv | Ja. | Js. | ConD. | All |
+Interestingly, among traditional approaches, the rule-based approach Major outperforms the other learning-based ones, LEAM and BERT, across most projects.
+| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Csv | Gs. | Ja. | Js. | ConD. | All |
 | #Bug | 26 | 65 | 27 | 106 | 38 | 133 | 39 | 18 | 16 | 18 | 26 | 93 | 246 | 851 |
@@
 | BF (GPT-4o) | 9 | 15 | 3 | 13 | 1 | 13 | 3 | 2 | 1 | 4 | 2 | 14 | 71 | 151 |
+| BF (DS-671b) | 6 | 10 | 3 | 11 | 1 | 13 | 2 | 2 | 0 | 3 | 2 | 15 | 83 | 152 |
 | LLMph (GPT-4o) | 6 | 25 | 7 | 24 | 6 | 22 | 9 | 2 | 3 | 3 | 5 | 12 | 81 | 205 |
+| LLMph (DS-671b) | 6 | 21 | 8 | 22 | 5 | 26 | 11 | 2 | 3 | 5 | 6 | 14 | 92 | 221 |
 | LLMut (GPT-3.5) | 10 | 32 | 8 | 46 | 14 | 38 | 18 | 2 | 4 | 2 | 11 | 46 | 175 | 406 |
@@
 | LLMut (GPT-4o-M) | 12 | 37 | 12 | 48 | 17 | 56 | 23 | 6 | 10 | 7 | 12 | 39 | 181 | 460 |
-| LLMut (DS-761b) | 13 | 40 | 15 | 62 | 21 | 79 | 27 | 8 | 10 | 12 | 17 | 59 | 146 | 509 |
+| LLMut (DS-671b) | 13 | 40 | 15 | 62 | 21 | 79 | 27 | 8 | 10 | 12 | 17 | 59 | 146 | 509 |
 | LLMut (DS-236b) | 12 | 34 | 11 | 47 | 14 | 45 | 20 | 7 | 8 | 6 | 11 | 41 | 177 | 433 |
@@
 BF: BugFarm. LLMph: LLMorpheus. LLMut: our prompt. Moc.: Mockito. Clo.: Closure. Co.: Codec. Gs.: Gson. Ja.: JacksonCore. Js.: Jsoup. ConD: ConDefects.
-Coupling Rate: The overall Coupling Rate for each mutation generation approach is shown in the last row of Table 4.
+Coupling Rate: The overall coupling rate for each mutant generation approach is shown in the column Coup.
+Rate of Table 4.
 The traditional approaches, PIT, Major, LEAM, and BERT, exhibit lower coupling rates than LLM-based approaches, with coupling rates of 23.5%, 38.9%, 38.1%, and 47.3%, respectively.
-Among the LLM-based approaches, LLMorpheus achieves the highest coupling rate, with a value of 51.7%.
-It is followed closely by DeepSeek-V3-761b and GPT-4o-Mini based on our prompt, with coupling rates of 50.8% and 50.6%, respectively.
-Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly LLMorpheus, DeepSeek-V3-761b, and GPT-4o, in generating overall highly coupled mutations that better mimic real bugs.
+Among the LLM-based approaches, LLMorpheus (DS-671b) and LLMorpheus (GPT-4o) achieve the highest coupling rates, reaching 52.0% and 51.7%, respectively, suggesting that LLMorpheus produces mutants whose failing behaviors align most closely with those of real bugs.
+LLMut (DS-671b) and LLMut (GPT-4o-M) closely follow LLMorpheus, with coupling rates of 50.8% and 50.6%, respectively.
+Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly LLMorpheus (DS-671b), in generating overall highly coupled mutants that better mimic real bugs.
 | Category | Strong | Strong Addi | Partial | Partial Addi | Not Subs. | Not Detected |
@@
 | BugFarm (GPT-4o) | 8.4% | 5.4% | 0.6% | 33.6% | 28.3% | 23.8% |
+| BugFarm (DS-671b) | 10.0% | 4.8% | 0.7% | 32.4% | 30.7% | 21.3% |
 | LLMorpheus (GPT-4o) | 9.3% | 4.9% | 2.1% | 35.4% | 23.6% | 24.6% |
+| LLMorpheus (DS-671b) | 11.8% | 4.6% | 1.8% | 33.8% | 23.9% | 24.1% |
 | LLMut (GPT-3.5) | 7.0% | 6.7% | 1.2% | 32.0% | 23.3% | 29.8% |
@@
 | LLMut (GPT-4o-M) | 6.2% | 12.1% | 0.8% | 30.4% | 21.3% | 29.2% |
-| LLMut (DS-761b) | 11.2% | 4.6% | 1.9% | 33.1% | 24.1% | 25.1% |
+| LLMut (DS-671b) | 11.2% | 4.6% | 1.9% | 33.1% | 24.1% | 25.1% |
 | LLMut (DS-236b) | 6.4% | 10.3% | 1.0% | 31.8% | 21.9% | 28.6% |
@@
 - Not Substitution.
-Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutation of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected.
-Note that the first four categories are used to calculate the Coupling Rate, indicating these mutations semantically overlap with their corresponding real bugs.
+Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutant of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected.
+Note that the first four categories are used to calculate the Coupling Rate, indicating these mutants semantically overlap with their corresponding real bugs.
 Table 7 presents the distribution of each generation approach.
 Among the traditional approaches, PIT and Major exhibit a significantly higher proportion of mutants in the Not Detected category, with rates of 45.9% and 47.9%, respectively, indicating that a substantial number of their generated mutants fail to couple with real bugs.
-However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutations.
-In contrast, LLM-based approaches such as DeepSeek-V3-761b, GPT-4o, and GPT-4o-Mini, demonstrate stronger coupling capabilities across multiple categories.
-For instance, DeepSeek-V3-761b achieves the highest proportion in the Strong category at 11.2%, while GPT-4o-Mini leads in the Strong + Additional Tests Fail category, with 12.1%.
-Overall, while the traditional rule-based approaches generate a larger proportion of strong substitution mutations, LLM-based approaches demonstrate a stronger ability to generate diverse and highly coupled mutations.
+However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutants.
+In contrast, LLM-based approaches overall demonstrate stronger coupling capabilities across multiple categories.
+For instance, LLMorpheus (DS-671b) achieves the highest proportion in the Strong category at 11.8%, closely followed by LLMut (DS-671b) with the value of 11.2%.
+BugFarm (DS-671b) also achieves a notable 10.0% of strongly coupled mutants.
+In terms of the Strong + Additional Tests Fail category, LLMut (GPT-4o-M) leads with a proportion of 12.1%.
+Additionally, BugFarm and LLMorpheus exhibit the lowest proportion of No Substitution and Not Detected, indicating their mutation strategies target to.
+Overall, LLM-based approaches generate a larger proportion of both strongly coupled and partially coupled mutants, demonstrating a stronger ability to generate diverse and highly coupled mutants.
 3.1.3. RQ1.3 Diversity
-We explore the syntactic diversity of mutations generated by each approach. Note that PIT works at the Java byte code level, we thus skip it in this RQ.
-AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element).
-For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang555https://pypi.org/project/javalang, to observe which new AST node types are introduced.
+We explore the syntactic diversity of mutants generated by each approach. Note that PIT works at the Java bytecode level, we thus skip it in this RQ.
+AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutant involves deletion (i.e., removing a code element).
+For the non-deletion mutants, we parse the code before and after mutation by the parser Javalang555https://pypi.org/project/javalang, to observe which new AST node types are introduced.
 A larger number of newly introduced AST node types indicates richer diversity.
 Figure 2 illustrates the number of introduced AST node types, with traditional approaches highlighted in green and LLM-based approaches in blue.
-The figure reveals that LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches.
-The best LLM-based approaches, DeepSeek-V3-761b and DeepSeek-Coder-V2-236b, exhibit the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types.
-Additionally, all six LLMs based on our prompt are significantly higher than all other approaches in AST node types.
-The existing LLM-based approaches, BugFarm and LLMorpheus, introduce 31 and 21 new AST node types, respectively.
-While BugFarm outperforms traditional approaches, it still lags behind ours.
+The figure reveals that LLM-based approaches present more types of new AST nodes than traditional approaches.
+The best LLM-based approaches, LLMut (DS-671b) and LLMut(DS-236b), exhibit the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types.
+Additionally, all models with LLMut introduce significantly more AST node types than traditional approaches.
+Meanwhile, BugFarm (DS-671b) and BugFarm (GPT-4o) present 33 and 31 new AST node types, respectively, which also outperform other traditional approaches.
 We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 2), as shown by the “Examples” in Figure 2.
 Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches.
-This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
+This indicates that LLMs are creative in generating mutants, producing significantly more diverse mutants than the examples provided to them.
 | Approach | Deletion | 1st | 2nd | 3rd |
@@
 | BugFarm (GPT-4o) | 5.5% | LT (41.9%) | BO (12.2%) | MI (9.0%) |
+| BugFarm (DS-671b) | 2.1% | LT(45.8%) | BO(14.7%) | MR(8.8%) |
 | LLMorpheus (GPT-4o) | 0.3% | LT (24.0%) | BO (23.7%) | MR (15.8%) |
+| LLMorpheus (DS-671b) | 0.3% | LT(29.5%) | BO(25.0%) | MR(15.2%) |
 | LLMut (GPT-3.5) | 5.3% | LT (39.2%) | BO (17.3%) | MI (9.9%) |
@@
 | LLMut (GPT-4o-M) | 0.5% | LT (43.3%) | BO (21.6%) | CC (7.6%) |
-| LLMut (DS-761b) | 3.8% | LT (44.2%) | BO (15.6%) | CC (8.6%) |
+| LLMut (DS-671b) | 3.8% | LT (44.2%) | BO (15.6%) | CC (8.6%) |
 | LLMut (DS-236b) | 1.0% | LT (38.8%) | BO (19.1%) | MI (10.1%) |
@@
 Deletion: The proportion of deletion mutations. BO: BinaryOperation. CC: ClassCreator. CT: Cast. LT: Literal. MI: MethodInvocation. MR: MemberReference. SE: StatementExpression. ST: Statement.
-We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the proportion of the mutations deleting the target code element and the top three most frequent AST node types.
-For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutation ratio, with 48.9%, followed by Major at 15.3%.
-In contrast, all LLM-based approaches generate deletion mutations at a rate below 6%, with three models producing less than 1.0%.
-This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations.
+We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the proportion of the mutants that delete the target code element and the top three most frequent AST node types.
+For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutant ratio, with 48.9%, followed by Major at 15.3%.
+In contrast, all LLM-based approaches generate deletion mutants at a rate below 6%, with three models producing less than 1.0%.
+This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutants.
 As shown later, we find that deletions are one of the key factors contributing to the lower compilation success rate.
-Regarding AST node types, the most frequent new AST node type of all the LLM-based approaches is Literal (e.g., ).
+Regarding AST node types, the most frequent new AST node type of all the LLM-based approaches is Literal (LT) (e.g., ).
 However, for the traditional approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement).
-Note that although these mutations are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion.
-AST Edit Distance: To quantify the extent of syntactic modifications introduced by each mutation generation approach, we compute the AST edit distance for each mutation using Gumtree (Falleri et al., 2014).
-The distribution of edit distances for each approach is presented in Table 9, providing insights into the complexity and diversity of the generated mutations.
+Note that although these mutants are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion.
+For the second most frequent new node types, BinaryOperation (BO) appears most often among LLM-based approaches, indicating that LLMs frequently modify or insert binary expressions such as arithmetic or comparison operators.
+For the third most frequent category, LLM-based approaches introduce a variety of AST nodes, including MethodInvocation (MI), MemberReference (MR), and ClassCreator (CC), reflecting their broader syntactic diversity in mutant generation.
+AST Edit Distance: To quantify the extent of syntactic modifications introduced by each mutant generation approach, we compute the AST edit distance for each mutant using Gumtree (Falleri et al., 2014).
+The distribution of edit distances for each approach is presented in Table 9, providing insights into the complexity and diversity of the generated mutants.
 | Distance | 1 | 2 | 3 | 4 | 5 | 6 |
@@
 | BugFarm (GPT-4o) | 44.7% | 43.5% | 8.3% | 1.6% | 0.6% | 1.4% |
+| BugFarm (DS-671b) | 38.3% | 52.1% | 6.7% | 1.5% | 0.7 | 0.7 |
 | LLMorpheus (GPT-4o) | 54.2% | 25.7% | 14.6% | 3.7% | 0.9% | 0.8% |
+| LLMorpheus (DS-671b) | 53.9% | 31.8% | 10.6% | 2.5% | 0.4% | 0.8% |
 | LLMut (GPT-3.5) | 52.4% | 36.4% | 6.8% | 2.4% | 0.9% | 1.0% |
@@
 | LLMut (GPT-4o-M) | 47.3% | 41.9% | 6.3% | 2.5% | 0.8% | 1.2% |
-| LLMut (DS-761b) | 46.1% | 44.2% | 6.8% | 1.5% | 0.4% | 0.9% |
+| LLMut (DS-671b) | 46.1% | 44.2% | 6.8% | 1.5% | 0.4% | 0.9% |
 | LLMut (DS-236b) | 45.1% | 44.2% | 6.5% | 2.6% | 0.8% | 0.8% |
@@
 From the table, we can find that all traditional approaches predominantly generate small-scale changes.
-Specifically, the proportion of mutations with an edit distance of 1 for Major, LEAM, and BERT is 69.2%, 58.5%, and 73.8%, respectively, indicating they involve minimal syntactic changes to the original code in the majority of cases.
+Specifically, the proportion of mutants with an edit distance of 1 for Major, LEAM, and BERT is 69.2%, 58.5%, and 73.8%, respectively, indicating they involve minimal syntactic changes to the original code in the majority of cases.
 In contrast, LLM-based approaches exhibit a broader range of edit distances, reflecting their capability to introduce more complex transformations.
-For example, DeepSeek-V3-761b exhibits a more balanced distribution, with 44.2% of its mutations having an edit distance of 2, 7.2% with a distance of 3, and 2.2% with a distance of 4, indicating it generates both minor and moderate code changes.
+For example, only 38.3% of the mutants of BugFarm (DS-671b) have an edit distance of 1, and LLMut (DS-671b) generates 45.1% of mutants with an edit distance of 1 and 44.2% with a distance of 2, with an additional 6.9% corresponding to distances of 3.
+This pattern suggests that LLM-based methods, unlike traditional approaches, are capable of performing multi-node or compositional edits, leading to a richer diversity of syntactic structures.
 3.2. RQ2: Performance on Validity
-To evaluate the validity of mutation generation approaches, we consider the aspects that may impact their validity, i.e., compilability rate, duplicate mutation rate, and equivalent mutation rate.
+To evaluate the validity of mutant generation approaches, we consider the aspects that may impact their validity, i.e., compilability rate, duplicate mutant rate, and equivalent mutant rate.
 This RQ is divided into the following sub-RQs.
@@
 For example, the Java compiler’s analyzer rejects constructs like while(true).
-PIT uses the mutation operators more restrictively and previously filters out non-compilable mutations.
+PIT uses the mutation operators more restrictively and previously filters out non-compilable mutants.
 In contrast, LEAM and BERT, have rates at 35.0% and 22.5%, respectively.
 For LLM-based approaches, there is a considerable variation in compilability.
-The existing LLM-based approaches, BugFarm and LLMorpheus, achieve relatively high rates of 74.1% and 67.8%, respectively.
-Among the LLMs using our prompt, DeepSeek-V3-761b attains the highest at 77.6%, followed closely by GPT-4o (76.4%), DeepSeek-Coder-V2-236b (75.8%), and GPT-4o-Mini (75.3%).
-Compared to small models (i.e., LEAM and BERT), GPTs and DeepSeek models show significantly higher compilation success rates.
-For example, the compilability rate of GPT-4o is over 30% higher than that of LEAM and BERT.
-3.2.2. RQ2.2: Duplicate Mutation Rate
-As aforementioned, duplicate mutations refer to those that are redundant or identical to the original code, providing no meaningful contribution to the mutation testing process.
-Therefore, a lower Duplicate Mutation Rate () is desirable, as it indicates a more efficient mutation generation process with minimal redundancy.
+Among all LLM-based approaches, LLMut (DS-671b) and BugFarm (DS-671b) attain the highest of 77.6% and 77.5%, respectively.
+LLMut (GPT-4o), LLMut (DS-236b), and LLMut (GPT-4o-M) closely follow by achieving 76.4%, 75.8%, 75.3%, respectively.
+Most LLM-based approaches show significantly higher compilation success rates compared to the traditional learning-based approaches, such as LEAM and BERT.
+For example, the s of BugFarm, LLMorpheus, and LLMut with GPT and DeepSeek models are over 30% higher than those of LEAM and BERT.
+3.2.2. RQ2.2: Duplicate Mutant Rate
+As aforementioned, duplicate mutants refer to those that are redundant or identical to the original code, providing no meaningful contribution to the mutant testing process.
+Therefore, a lower Duplicate Mutant Rate () is desirable, as it indicates a more efficient mutant generation process with minimal redundancy.
 The column Dup.
@@
 Other traditional approaches, i.e., LEAM and BERT, show relatively low s of 2.3% and 1.7%, respectively.
-In contrast, LLMs generate a significant portion of duplicate mutations, exhibiting a higher .
-For the LLMs from OpenAI, GPT-3.5, GPT-4o, and GPT-4o-Mini exhibit moderate s of 10.3%, 7.8%, and 7.0%, respectively.
-For the existing LLM-based approaches, BugFarm and LLMorpheus, exhibit similar s of 7.5% and 7.1%, respectively.
-The open-source LLMs such as StarCoder and DeepSeek-Coder-V2-236b demonstrate higher s of 12.8% and 7.7%, respectively, while CodeLlama shows the highest at 37.1%.
-3.2.3. RQ2.3: Equivalent Mutations
-Given the undecidable nature of identifying equivalent mutations, we manually evaluate each mutation’s equivalence, following the methodology of the existing study (Kushigian et al., 2024).
-For each mutation generation approach, we randomly sample a subset of compilable mutations, ensuring a 95% confidence level with a 10% margin of error.
-Two authors independently evaluate the equivalence of the selected mutations to minimize subjectivity.
-The procedure for checking a mutation involves first transplanting it into the original project and executing all tests.
-If it passes all tests, the author then manually verifies whether the mutation is semantically equivalent.
-Note that PIT does not generate source-level mutations, so its mutations must be manually extracted from Java bytecode and transformed back into source-level representations, which is significantly more time-consuming.
-In total, 1247 mutations were labeled, 709 of which were filtered by tests (i.e., killed), and the remaining 538 were manually checked.
-The entire process took approximately twelve hours per person.
+In contrast, LLMs generate a significant portion of duplicate mutants, exhibiting a higher .
+For the LLM-based approaches evaluated with the best-performing model, DS-671b, the duplicate mutation rates (s) are as follows: BugFarm (DS-671b): 7.5%, LLMorpheus (DS-671b): 6.9%, and LLMut (DS-671b): 7.5%.
+The smaller models on LLMut exhibit much worse performance, such as LLMut (SC-16b) and LLMut (CL-13b).
+These results indicate that even equipped with the state-of-the-art LLM, these approaches still generate a non-negligible proportion of duplicate mutants.
+3.2.3. RQ2.3: Equivalent Mutants
+Given the undecidable nature of identifying equivalent mutants, we manually evaluate each mutant’s equivalence, following the methodology of the existing study (Kushigian et al., 2024).
+For each mutant generation approach, we randomly sample a subset of compilable mutants, ensuring a 95% confidence level with a 10% margin of error.
+Two authors independently evaluate the equivalence of the selected mutants to minimize subjectivity.
+The procedure for checking a mutant involves first transplanting it into the original project and executing all tests.
+If it passes all tests, the author then manually verifies whether the mutant is semantically equivalent.
+Note that PIT does not generate source-level mutants, so its mutants must be manually extracted from Java bytecode and transformed back into source-level representations, which is significantly more time-consuming.
+In total, 1437 mutants were labeled, 860 of which were filtered by tests (i.e., killed), and the remaining 577 were manually checked.
+The entire process took approximately thirteen hours per person.
 The manual verification requires checking the code context, which may involve multiple classes and methods.
-For example, to check a mutation that replaces an integer with a method invocation, where the method is invoked by the this object but defined in the parent class, we should review the parent class file and verify whether its return value is consistently equal to the integer.
+For example, to check a mutant that replaces an integer with a method invocation, where the method is invoked by the this object but defined in the parent class, we should review the parent class file and verify whether its return value is consistently equal to the integer.
 We measure the consistency between the two authors by Cohen’s kappa coefficient () (Viera et al., 2005), achieving an overall value of 0.8, which indicates a high level of agreement.
@@
 | BugFarm (GPT-4o) | 96 | 67 | 3 | 3.1% |
+| BugFarm (DS-671b) | 96 | 75 | 5 | 5.2% |
 | LLMorpheus (GPT-4o) | 96 | 69 | 4 | 4.2% |
+| LLMorpheus (DS-671b) | 96 | 65 | 3 | 3.1% |
 | LLMut (GPT-3.5) | 96 | 60 | 4 | 4.2% |
@@
 | LLMut (GPT-4o-M) | 96 | 59 | 3 | 3.1% |
-| LLMut (DS-761b) | 96 | 65 | 6 | 6.3% |
+| LLMut (DS-671b) | 96 | 65 | 6 | 6.3% |
 | LLMut (DS-236b) | 96 | 63 | 5 | 5.2% |
@@
 | LLMut (CL-13b) | 96 | 51 | 6 | 6.3% |
-Table 10 presents the estimated Equivalent Mutation Rate (), reporting the number of sampled mutations (#Sampled), the number of equivalent mutations we identified (#Eq.
+Table 10 presents the estimated Equivalent Mutant Rate (), reporting the number of sampled mutants (#Sampled), the number of equivalent mutants we identified (#Eq.
 Mut), and the for each approach.
-Among the traditional approaches, PIT achieves the lowest of 1.0%, followed by Major with an of 2.1%, indicating that rule-based techniques produce fewer equivalent mutations.
+Among the traditional approaches, PIT achieves the lowest of 1.0%, followed by Major with an of 2.1%, indicating that rule-based techniques produce fewer equivalent mutants.
 In contrast, the LLM-based approaches exhibit higher s, with variations across different models.
-StarChat has the highest s of 10.6%, while GPT-4o, GPT-4o-Mini, and DeepSeek-Coder-V2-236b show relatively lower s, ranging from 3.1% to 5.2%.
-Overall, although the OpenAI models such as GPT-4o and GPT-4o-Mini tend to have slightly higher s than PIT and Major, the increase remains modest and within an acceptable range.
+LLMut (SC-16) performs the worst by attaining the highest s of 10.6%, while LLMut (GPT-4o), LLMut (GPT-4o-M), and LLMut (DS-236b) show relatively lower s, ranging from 3.1% to 5.2%.
+BugFarm and LLMorpheus exhibit s ranging from 3.1% to 5.2%.
+Overall, the LLM-based approaches with advanced models (e.g., GPT-4o or DS-671b) tend to have slightly higher s than PIT and Major, the increase remains modest and within an acceptable range.
 Note that the estimated of Major given by the recent study of Kushigian et al.
-(Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutations for manual judgment.
-We adopt the same experimental process and the result of Major is 2.0%, which is consistent with theirs, considering the randomness introduced by sampling.
-Moreover, the definition of equivalent mutations in this paper is the set of syntactically different but semantically equivalent mutations, which may differ from existing studies (Tip et al., 2024).
+(Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutants for manual judgment.
+We adopt the same experimental process, and the result of Major is 2.0%, which is consistent with theirs, considering the randomness introduced by sampling.
+Moreover, the definition of equivalent mutants in this paper is the set of syntactically different but semantically equivalent mutants, which may differ from existing studies (Tip et al., 2025).
 3.3. RQ3: Performance on Efficiency
@@
 Average Generation Time: The column Time of Table 4 presents the results of Average Generation Time in seconds.
-Among the traditional appraoches, PIT demonstrates the fastest generation time at 0.02 seconds per mutation, followed by Major, which takes 0.08 seconds, making them the most efficient approaches.
+Among the traditional approaches, PIT demonstrates the fastest generation time at 0.02 seconds per mutant, followed by Major, which takes 0.08 seconds, making them the most efficient approaches.
 In contrast, LEAM and BERT take 3.06 seconds and 2.34 seconds, respectively.
-The existing LLM-based approaches, Among the LLMs, the LLMs of OpenAI (i.e., GPT-3.5, GPT-4o, and GPT-4o-Mini) show relatively faster generation, ranging from 1.31 to 1.79 seconds.
-The open-source LLMs StarChat, CodeLlama, and DeepSeek require 7.53, 9.06, and 4.25 seconds, respectively.
-The results suggest that the rule-based approaches are the most time-efficient for generating mutations, which are tens to hundreds of times faster than other approaches.
+The LLM-based approaches present much slower generation speed, ranging from 1.31 to 9.06 seconds.
+In particular, all LLM-based approaches require more time when adopting open-source models.
+For example, LLMut (SC-16b), LLMut (CL-13b), and LLMut (DS-236b) require 7.53, 9.06, and 4.25 seconds, respectively.
+The results suggest that the rule-based approaches are the most time-efficient for generating mutants, which are tens to hundreds of times faster than other approaches.
 Note that the time costs of self-deployed open-source models typically depend on the types of GPUs used for execution.
 Additionally, the response time of models accessed via network APIs may vary across service providers and can be further affected by network fluctuations.
-We further discuss these potential impacts in the Section 5.
-| Approach | Avg. Input Token | Avg. Output Token |
-| BugFarm (GPT-4o) | 200 | 368 |
-| LLMorpheus (GPT-4o) | 370 | 74 |
-| LLMut (GPT-3.5) | 126 | 59 |
-| LLMut (GPT-4o) | 126 | 59 |
-| LLMut (GPT-4o-M) | 126 | 59 |
-| LLMut (DS-761b) | 126 | 58 |
-| LLMut (DS-236b) | 126 | 59 |
-| LLMut (SC-16b) | 126 | 63 |
-| LLMut (CL-13b) | 126 | 58 |
-Average Token: Table 11 presents the average number of input and output tokens used by each LLM-based approach to generate one mutation.
-The existing LLM-based approaches, BugFarm and LLMorpheus, consume 200 and 370 input tokens, and 368 and 74 output tokens per mutation, respectively.
-LLMs using our prompt consume substantially fewer tokens, averaging 126 input tokens and 59–63 output tokens per mutation.
-As all LLMs share identical inputs, their input token counts remain the same.
-3.4. RQ4: Categorizations of Non-Compilable Mutations
-To answer this RQ, we first identify the error types of non-compilable mutations and then analyze the surrounding code context types that are error-prone to generate them.
+We further discuss these potential impacts in Section 5.
+| Approach | Avg. Input Token | Avg. Output Token | All Input Token | All Output Token |
+| BugFarm (GPT-4o) | 200 | 368 | 2,780,213 | 4,807,979 |
+| BugFarm (DS-671b) | 200 | 366 | 2,780,213 | 5,085,936 |
+| LLMorpheus (GPT-4o) | 370 | 74 | 7,921,356 | 1,476,568 |
+| LLMorpheus (DS-671b) | 370 | 73 | 7,921,356 | 1,558,988 |
+| LLMut (GPT-3.5) | 126 | 59 | 4,353,426 | 1,855,042 |
+| LLMut (GPT-4o) | 126 | 59 | 4,353,426 | 1,875,810 |
+| LLMut (GPT-4o-M) | 126 | 59 | 4,353,426 | 2,027,850 |
+| LLMut (DS-671b) | 126 | 58 | 4,353,426 | 1,947,224 |
+| LLMut (DS-236b) | 126 | 59 | 4,353,426 | 2,064,658 |
+| LLMut (SC-16b) | 126 | 63 | 4,353,426 | 1,766,537 |
+| LLMut (CL-13b) | 126 | 58 | 4,353,426 | 1,599,127 |
+Average Token: Table 11 presents the average number of input and output tokens used by each LLM-based approach to generate a single mutant.
+All the token usage is computed as the official library of OpenAI, Tiktoken666https://github.com/openai/tiktoken, which is the standard practice for estimating model usage.
+From the table, we observe substantial differences across prompting strategies.
+BugFarm uses relatively short prompts (200 input tokens on average), but its output token usage is high (over 360 tokens) because it returns the code of the entire mutated method.
+Consequently, its overall token consumption needs over 2.7M input tokens and about 5M output tokens.
+In contrast, LLMorpheus requires the largest number of input tokens (370 on average, 7.9M in total), as its prompt includes up to 200 lines of surrounding context, yet it produces moderate output (around 74 tokens on average, 1.5M in total).
+Moreover, LLMut requires a moderate amount of input and output tokens, which cost 126 input tokens and 58-63 output tokens to generate one mutant.
+In total, LLMut needs 4.3M input tokens and 1.6M-2.1M output tokens, which are also moderate.
+Since LLMut asks the models to generate more mutants within a single conversation and wraps the output mutants directly with JSON tags, the average cost is significantly lower than BugFarm and LLMorpheus.
+These results demonstrate that various prompt engineering strategies significantly impact token costs.
+3.4. RQ4: Categorizations of Non-Compilable Mutants
+To answer this RQ, we first identify the error types of non-compilable mutants and then analyze the surrounding code contexts that are error-prone in generating them.
 3.4.1. Categorizations of compilation error types
-Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources.
-As mentioned in Section 3.2, LLMs generate many non-compilable mutations.
-This RQ analyzes the types of errors and potential root causes of these non-compilable mutations.
-Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT models, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015).
-From the manual analysis of these non-compilable mutations, we identified nine distinct error types, shown as follows.
+Non-compilable mutants require a compilation step to filter out, which results in wasted computational resources.
+As mentioned in Section 3.2, LLMs generate many non-compilable mutants.
+This RQ analyzes the types of errors and potential root causes of these non-compilable mutants.
+Following the setting of the previous steps, we first sample 384 non-compilation mutants from the outputs of GPT models, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015).
+From the manual analysis of these non-compilable mutants, we identified nine distinct error types, shown as follows.
 Usage of Unknown Methods: This type of error is due to invoking unknown methods under the code context.
@@
 Incorrect Initialization: This type of error is due to using illegal forms of constructors or deleting necessary initialization statements.
-Incorrect Location: This type of error is due to a mutation being applied to an incorrect code location.
-Incorrect Exceptions: This type of error is due to throwing illegal exceptions.
+Incorrect Location: This type of error is due to a mutant being applied to an incorrect code location.
+Incorrect Exceptions: This type of error is due to throwing incorrect exceptions.
 | ID | Error Type | % | ID | Error Type | % |
@@
 These findings underscore the need for further improvements in LLMs to enhance their reliability and accuracy in code generation.
-3.4.2. Categorizations of code context of non-compilable mutations
-To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by all nine mutation generation approaches in Table 4, as shown in Figure 3.
+3.4.2. Categorizations of code context of non-compilable mutants
+To analyze which types of code are prone to causing LLMs to generate non-compilable mutants, we examined the code locations of all non-compilable mutants generated by all nine mutant generation approaches in Table 4, as shown in Figure 3.
 For most approaches (except Major and BugFarm), the code locations with MemberReference and MethodInvocation are the top-2 most-prevent AST node types.
-For example, for GPT and DeepSeek models, about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation.
+For example, for GPT and DeepSeek models, about 40% of non-compilable mutants occur at the location with MemberReference or MethodInvocation.
 This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references.
-If any required method or member is missing or incorrectly specified, it can easily lead to non-compileable mutations.
-The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure.
-Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 3.4% (DeepSeek-V3-761b), 1.11% (DeepSeek-Coder-V2-236b), 3.4% (BugFarm), 0.8% (LLMorpheus), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% (BERT), and 16.61% (Major) of all their non-compilable mutations, respectively.
+If any required method or member is missing or incorrectly specified, it can easily lead to non-compileable mutants.
+The errors highlight a need for better context-aware mutant generation, ensuring that method calls and member references align with the intended program structure.
+Additionally, we inspect the deletion mutants rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 3.4% (DeepSeek-V3-671b), 1.11% (DeepSeek-Coder-V2-236b), 3.4% (BugFarm), 0.8% (LLMorpheus), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% (BERT), and 16.61% (Major) of all their non-compilable mutants, respectively.
 Thus, for LLMs, deletion is not the major reason for non-compilation.
-3.5. RQ5: Distributions of Alive Mutations
-As shown in Table 4, the mutation scores of rule-based approaches (i.e., PIT and Major) are lower than LLMs, indicating their mutations are harder to detect.
-Therefore, we intend to analyze the reasons for survival mutations.
-To address this RQ, we adopt a setting similar to the analysis of equivalent mutations.
-Specifically, for each approach, we randomly sample its alive mutations and manually classify them.
-We employ the categories of alive mutation causes defined in the study by Just et al.
+3.5. RQ5: Distributions of Surviving Mutants
+As shown in Table 4, the mutation scores of rule-based approaches (i.e., PIT and Major) are lower than those of LLMs, indicating their mutants are harder to detect.
+Therefore, we intend to analyze the reasons for surviving mutants.
+To address this RQ, we adopt a setting similar to the analysis of equivalent mutants.
+Specifically, for each approach, we randomly sample its surviving mutants and manually classify them.
+We employ the categories of surviving mutant causes defined in the study by Just et al.
 (Just et al., 2014a) as follows.
-(1) Not covered, i.e., the mutation is located in code never executed by the test suite. We refer to the mutation set of this category as .
-(2) Covered but no state infection, i.e., the mutation is executed but does not propagate any state change. We refer to the mutation set of this category as .
-(3) State infection without detection, i.e., the mutation alters the program state but any test oracle does not observe the deviation. We refer to the mutation set of this category as .
+(1) Not covered, i.e., the mutant is located in code never executed by the test suite. We refer to the mutant set of this category as .
+(2) Covered but no state infection, i.e., the mutant is executed but does not propagate any state change. We refer to the mutant set of this category as .
+(3) State infection without detection, i.e., the mutant alters the program state but any test oracle does not observe the deviation. We refer to the mutant set of this category as .
 These categories are mutually exclusive, with the following relationships: , , and .
-This classification enables us to analyze the underlying reasons for alive mutations and further evaluate the extent to which different approaches contribute to strengthening the test suite.
-Given an alive mutation, we first determine whether it is equivalent; if so, it is discarded.
-For non-equivalent mutations, we then check whether they are uncovered.
-If a mutation is covered, we further examine whether it alters the program state, i.e., whether any memory values are modified.
-Finally, the remaining alive mutations that are covered but do not lead to detectable differences are classified as not detected.
-This stepwise process ensures that every alive mutation is uniquely assigned to one category (equivalent, uncovered, not injecting state, or not detected).
+This classification enables us to analyze the underlying reasons for surviving mutants and further evaluate the extent to which different approaches contribute to strengthening the test suite.
+Given a surviving mutant, we first determine whether it is equivalent; if so, it is discarded.
+For non-equivalent mutants, we then check whether they are uncovered.
+If a mutant is covered, we further examine whether it alters the program state, i.e., whether any memory values are modified.
+Finally, the remaining surviving mutants that are covered but do not lead to detectable differences are classified as not detected.
+This stepwise process ensures that every surviving mutant is uniquely assigned to one category (equivalent, uncovered, not injecting state, or not detected).
 | Approach | #Sampled | #Eq. Mut. | NC | NI | ND | NC% | NI% | ND% |
@@
 | BugFarm (GPT-4o) | 94 | 9 | 62 | 11 | 12 | 72.94% | 12.94% | 14.12% |
+| BugFarm (DS-671b) | 94 | 11 | 52 | 15 | 16 | 62.65% | 18.07% | 19.28% |
 | LLMopheus (GPT-4o) | 95 | 11 | 26 | 44 | 14 | 30.95% | 52.38% | 16.67% |
+| LLMopheus (DS-671b) | 95 | 13 | 20 | 47 | 15 | 24.39% | 57.32% | 18.29% |
 | LLMut (GPT-3.5) | 95 | 12 | 66 | 12 | 5 | 79.52% | 14.46% | 6.02% |
@@
 | LLMut (GPT-4o-M) | 95 | 12 | 51 | 15 | 17 | 61.45% | 18.07% | 20.48% |
-| LLMut (DS-761b) | 95 | 12 | 45 | 16 | 22 | 54.22% | 19.28% | 26.51% |
+| LLMut (DS-671b) | 95 | 12 | 45 | 16 | 22 | 54.22% | 19.28% | 26.51% |
 | LLMut (DS-236b) | 95 | 13 | 54 | 15 | 13 | 65.85% | 18.29% | 15.85% |
@@
 | LLMut (CL-13b) | 96 | 11 | 68 | 12 | 5 | 80.00% | 14.12% | 5.88% |
-As shown in Table 13, the distributions of alive mutations differ significantly across approaches, where the columns NC%, NI%, and ND% represent the proportions of , , and within all non-equivalent mutations, respectively.
-First, LLM-based methods exhibit a substantially higher proportion of (i.e., not covered) mutations, meaning that more of their alive mutations reside in previously uncovered code lines.
-For instance, among the alive mutations generated by GPT-3.5 and BugFarm, 79.52% and 72.94% fall into the category, respectively, whereas PIT and Major exhibit much lower proportions of 46.24% and 45.65%.
+As shown in Table 13, the distributions of surviving mutants differ significantly across approaches, where the columns NC%, NI%, and ND% represent the proportions of , , and within all non-equivalent mutants, respectively.
+From the table, we can make the following observations.
+First, BugFarm and LLMut exhibit a substantially higher proportion of (i.e., not covered) mutants, while LLMopheus shows a higher proportion of (i.e., not injected) mutants.
+For instance, among the surviving mutants generated by BugFarm (GPT-4o) and BugFarm (DS-671b), 79.52% and 72.94% fall into the category, respectively.
+In contrast, the traditional approaches PIT, Major, and Bert exhibit much lower proportions of 46.24%, 45.65%, and 46.25%, respectively.
+This suggests that LLM-based approaches either tend to generate mutants located in previously untested code regions or, in the case of LLMorpheus, often produce mutations that do not alter the program state.
 This characteristic highlights the potential of LLMs in guiding test suite augmentation by directing attention to untested program areas.
-Second, traditional approaches, especially rule-based methods such as PIT and Major, produce a larger fraction of (i.e., not injected) and (i.e., not detected) mutations compared to LLM-based methods.
-For instance, PIT and Major exhibit proportions of 29.03% and 32.61%, respectively, whereas all LLMs based on our prompt remain below 20%.
-These types of alive mutations require not only generating inputs that execute the mutated code but also propagating and detecting subtle semantic differences, which is inherently more challenging than increasing coverage.
-Therefore, the overall implication is that LLM-generated alive mutations are more directly beneficial for strengthening unit tests, as they emphasize extending coverage, while rule-based approaches tend to stress the harder task of state propagation and detection.
+Notably, a recent study guides LLMs to generate unit test cases to kill surviving mutants (Harman et al., 2025; Wang et al., 2025b), indicating the usefulness of these mutants for other applications.
+Second, traditional approaches, especially rule-based methods such as PIT and Major, produce a larger fraction of (i.e., not injected) and (i.e., not detected) mutants compared to LLM-based methods.
+For instance, PIT and Major exhibit proportions of 29.03% and 32.61%, respectively, whereas both LLMut and BugFarm across different models remain below 20%.
+These types of surviving mutants require not only generating inputs that execute the mutated code but also propagating and detecting subtle semantic differences, which is inherently more challenging than increasing coverage.
+Therefore, the overall implication is that LLM-generated surviving mutants are more directly beneficial for strengthening unit tests, as they emphasize extending coverage, while rule-based approaches tend to stress the harder task of state propagation and detection.
 4. Discussion
-In this section, we explore the impact of different experiment setups for LLMs, then we discuss the implications.
+In this section, we examine the impact of various experiment setups for LLMs and discuss their implications.
 4.1. Sensitivity to Experimental Settings
@@
 Prompts play a critical role in the performance of LLMs.
-Therefore, we aim to investigate how different prompts influence mutation generation effectiveness.
+Therefore, we aim to investigate how different prompts influence mutant generation effectiveness.
 Due to budget constraints, we selected a subset of 306 bugs from our dataset, comprising 10 randomly sampled bugs from each Defects4J project and all bugs from the ConDefects dataset.
-For the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the recent open-source model DeepSeek-V3-761b, then switch different prompt templates.
-More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get five types of prompts, shown as follows:
-Prompt 1 (P1): The Default Prompt. P1 is the default prompt, shown as Figure 1.
-Prompt 2 (P2): P1 without Few-Shot Examples. P2 is based on P1 by removing all the few-shot examples.
-Prompt 3 (P3): P2 without Whole Java Method. P3 is based on P2 by removing the surrounding Java method code, only keeping the target code element to be mutated.
-Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of the corresponding unit tests of the target method into Context.
-Prompt 5 (P5): P1 with Whole Java File. P5 is a variant of P1, which uses the code of the entire Java file as the code context, rather than the entire method.
-| Metric | LLMut (GPT-3.5) | LLMut (DS-761b) | ||||||||
+Based on the prompt design of LLMut, we investigate how different sources of contextual information affect mutant generation.
+For this experiment, we select two representative LLMs—the closed-source GPT-4o and the recent open-source DeepSeek-V3-671b—and vary the prompt template used for mutation generation.
+Note that based on the evaluation results, both models are leading LLMs in various metrics.
+Starting from the prompt that includes the richest contextual information, we progressively remove specific sources of context in the Context section.
+This process yields five distinct prompt variants, summarized as follows:
+Prompt 1 (P1): P1 is the richest prompt with the code of the whole Java source file as the Context of the prompt.
+Prompt 2 (P2): P2 is the prompt with the code of the focal method and its corresponding unit tests as the Context of the prompt.
+Prompt 3 (P3): P3 is the prompt with the code of the focal method as the Context of the prompt, which is the default prompt in the previous section, shown as Figure 1.
+Prompt 4 (P4): P4 is the prompt with the code of the target code element to be mutated as the Context of the prompt.
+Prompt 5 (P5): P5 is the prompt with the code of the focal method as the Context, but remove all few-shot examples.
+| Metric | LLMut (GPT-4o) | LLMut (DS-671b) | ||||||||
 | P1 | P2 | P3 | P4 | P5 | P1 | P2 | P3 | P4 | P5 | |
-| Mut. # | 3681 | 3688 | 3350 | 3210 | 3496 | 3851 | 3847 | 3834 | 3844 | 3878 |
-| Mut. Score | 0.773 | 0.785 | 0.842 | 0.801 | 0.694 | 0.765 | 0.750 | 0.760 | 0.753 | 0.776 |
-| Real Bug Detec. | 79.5% | 78.4% | 72.2% | 72.5% | 79.5% | 85.5% | 85.8% | 83.9% | 84.7% | 83.1% |
-| Avg. Ochiai | 63.1% | 61.8% | 60.8% | 62.1% | 62.3% | 67.5% | 65.6% | 66.4% | 65.8% | 66.4% |
-| Coupling Rate | 52.1% | 52.2% | 47.4% | 49.6% | 47.3% | 53.2% | 51.7% | 50.9% | 50.0% | 49.1% |
-| Compling Rate | 66.9% | 66.8% | 60.0% | 61.2% | 70.3% | 82.8% | 83.5% | 78.0% | 82.0% | 79.5% |
-| Dup. Mut. Rate | 11.4% | 13.3% | 14.9% | 12.7% | 16.9% | 7.3% | 6.4% | 8.5% | 7.0% | 8.2% |
-| Avg. Gen. Time | 1.789 | 1.782 | 1.09 | 1.884 | 1.907 | 2.573 | 2.571 | 2.488 | 2.601 | 2.596 |
-| Avg. # Token | 105 | 95 | 44 | 219 | 393 | 104 | 92 | 43 | 219 | 392 |
-Note: The prompt template that achieves the best performance is highlighted in grey background color.
-Table 14 presents a comparative evaluation of different prompts for GPT-3.5 and DeepSeek-V3-761b across multiple metrics.
-Across both models, P1 delivers the best overall performance on all effectiveness metrics (i.e., real bug detectability, average Ochiai, and coupling rate), demonstrating its strength in generating mutations that are semantically closer to real bugs.
-For the validity metrics, P1 ranks second in compilation rate on GPT-3.5 and first on DeepSeek-V3-761b, while achieving the lowest duplicate mutation rate on GPT-3.5 and third-lowest on DeepSeek-V3-761b.
-Notably, compared with P5, which provides the entire file as context, P1 relies only on the full method yet achieves higher semantic similarity to real bugs, with reduced generation time and token usage, while maintaining comparable compilation and duplicate mutation rates.
-Across P1 to P5, we observe a clear context–performance trade-off.
-Removing relevant context (P3) substantially degrades effectiveness, as seen in lower real bug detectability, average Ochiai, and coupling rates.
-Providing a focused, method-level context (P1 and P2) delivers the best overall balance of effectiveness and validity.
-However, adding excessive context beyond the focal method (P4 and P5) does not yield further improvements and often increases token usage and generation time, while in some cases even reducing key effectiveness metrics.
-This suggests that moderate, relevant context benefits LLM mutation generation, whereas both insufficient and excessive context hurt efficiency and quality.
+| Mut. # | 3696 | 3692 | 3781 | 3807 | 3788 | 3878 | 3844 | 3851 | 3834 | 3847 |
+| Mut. Score | 0.841 | 0.801 | 0.773 | 0.832 | 0.785 | 0.776 | 0.753 | 0.765 | 0.760 | 0.750 |
+| Real Bug Detec. | 83.1% | 83.9% | 85.0% | 83.1% | 84.7% | 83.1% | 84.7% | 85.5% | 83.9% | 85.8% |
+| Avg. Ochiai | 64.8% | 65.0% | 66.4% | 64.8% | 65.3% | 66.4% | 65.8% | 67.5% | 66.4% | 65.6% |
+| Coupling Rate | 47.9% | 49.7% | 52.2% | 48.4% | 52.1% | 49.1% | 50.0% | 53.2% | 50.9% | 51.7% |
+| Compilability Rate | 78.4% | 81.6% | 80.1% | 77.4% | 82.5% | 79.5% | 82.0% | 82.8% | 78.0% | 83.5% |
+| Dup. Mut. Rate | 7.7% | 7.9% | 7.9% | 8.7% | 8.1% | 8.2% | 7.0% | 7.3% | 8.5% | 6.4% |
+| Avg. Gen. Time | 1.607 | 1.584 | 1.389 | 1.017 | 1.282 | 2.596 | 2.601 | 2.573 | 2.488 | 2.571 |
+| Avg. # Token | 392 | 218 | 105 | 44 | 94 | 392 | 219 | 104 | 43 | 92 |
+Note: P3 is the prompt used in the previous section. The prompt template that achieves the best performance is highlighted in grey background color.
+Table 14 presents a comparative evaluation of different prompts on two models across the main effectiveness, validity, and efficiency metrics.
+For the effectiveness metrics (i.e., real bug detectability, average Ochiai, and coupling rate), on both models, P3, the prompt used in the previous section, delivers the best overall performance, demonstrating its strength in generating mutants that are semantically closer to real bugs.
+P5, which removes the few-shot examples, ranks as the second-best prompt overall, suggesting that few-shot examples help guide the model toward generating mutants closer to real bugs.
+Interestingly, P1, the longest prompt with the whole Java file, performs the worst in most cases, indicating that providing more information does not necessarily help LLMs generate more bug-representative mutants.
+P2, which adds unit tests to the context compared with P3, shows slightly reduced effectiveness—suggesting that including unit tests does not help the model generate better mutants.
+For the validity metrics, different prompts express a similar trend in both models.
+For the compilability rate, P5, the one without real-world bug few-shot examples, achieves the highest compilability rate on both models.
+This indicates that removing few-shot examples reduces the syntactic complexity of the generated mutations, making them easier for LLMs to produce without introducing compilation errors.
+For the duplicated mutant rate, P1 achieves the best performance on GPT-4o, while P5 performs best on DS-671b, suggesting that duplication is more influenced by model behavior than by the prompts.
+For the efficiency costs, the results show a clear trend: prompts with more contextual information consume more time and tokens.
+For example, P1 (whole file) incurs the highest cost, followed by P3 (whole method), while P5 (only the code snippet to be mutated) is the most efficient.
+Another interesting finding is that P1 (whole file) is significantly longer than P2 (unit tests).
 We further explore the performance similarity of different prompts.
@@
 We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the -values are all below 0.05.
-Although P3 achieves the most balanced overall performance, the differences among prompts are not statistically significant.
+The results suggest that on both GPT-4o and DS-671b, although P3 achieves the most balanced overall performance, the differences among prompts are not statistically significant.
 4.1.2. Experiment Setup of the Context Length
 | Metric | 1-Line Context | 2-Line Context | 3-Line Context |
-| Real Bug Dectc. | 85.0% | 83.3% | 83.3% |
-| Avg. Ochiai | 33.3% | 35.0% | 30.0% |
-| Coupling Rate | 36.6% | 35.2% | 37.1% |
-| Compilability Rate | 62.9% | 63.2% | 67.2% |
-| Dup. Mut. Rate | 10.8% | 12.5% | 11.9% |
+| Real Bug Dectc. | 93.3% | 91.7% | 93.3% |
+| Avg. Ochiai | 46.7% | 50.0% | 48.3% |
+| Coupling Rate | 37.3% | 36.8% | 37.7% |
+| Compilability Rate | 80.7% | 81.4% | 82.1% |
+| Dup. Mut. Rate | 4.9% | 6.6% | 5.3% |
 | Settings | Spea. Coef | -value | Pear. Coef | -value |
-| GPT-3.5 P1 v.s. GPT-3.5 P2 | 0.9762 | 3.31E-05 | 0.9999 | 1.35E-31 |
-| GPT-3.5 P1 v.s. GPT-3.5 P3 | 0.9040 | 2.00E-03 | 0.9999 | 8.66E-23 |
-| GPT-3.5 P1 v.s. GPT-3.5 P4 | 0.9520 | 2.60E-04 | 0.9999 | 1.94E-31 |
-| GPT-3.5 P1 v.s. GPT-3.5 P5 | 0.9762 | 3.31E-05 | 0.9999 | 1.57E-26 |
-| DS-671b P1 v.s. DS-671b P2 | 1.0000 | 0.00E+00 | 0.9999 | 1.08E-31 |
-| DS-671b P1 v.s. DS-671b P3 | 1.0000 | 0.00E+00 | 0.9999 | 2.93E-29 |
-| DS-671b P1 v.s. DS-671b P4 | 1.0000 | 0.00E+00 | 0.9999 | 4.95E-31 |
-| DS-671b P1 v.s. DS-671b P5 | 1.0000 | 0.00E+00 | 0.9999 | 1.51E-30 |
-| 1-Line vs. 2-Line | 0.9762 | 3.31E-05 | 0.9980 | 2.12E-08 |
-| 2-Line vs. 3-Line | 0.9762 | 3.31E-05 | 0.9997 | 1.04E-10 |
-| 1-Line vs. 3-Line | 0.9524 | 2.60E-04 | 0.9978 | 2.69E-08 |
-| QB-6a (Default) v.s. QB-6b | 0.9429 | 4.81E-03 | 0.9984 | 1.00E-08 |
-| QB-6a (Default) v.s. QB-3 | 0.9856 | 3.09E-05 | 0.9999 | 1.62E-13 |
-| QB-6a (Default) v.s. QB-4 | 0.9643 | 4.54E-04 | 0.9999 | 7.28E-23 |
-| QB-6a (Default) v.s. QB-5 | 0.9643 | 4.54E-04 | 0.9999 | 9.62E-24 |
-| QB-6a (Default) v.s. QB-7 | 1.0000 | 0.0E+00 | 0.9999 | 9.39E-16 |
-| QB-6a (Default) v.s. QB-8 | 1.0000 | 0.0E+00 | 0.9999 | 3.34E-15 |
-| QB-6a (Default) v.s. QB-9 | 1.0000 | 0.00E+00 | 0.9999 | 5.01E-14 |
-| QB-6a (Default) v.s. QB-10 | 0.9856 | 3.09E-04 | 0.9999 | 2.89E-13 |
-| QB-6a (Default) v.s D4J-6 | 0.9524 | 2.60E-04 | 0.9980 | 2.13E-08 |
+| GPT-4o P3 v.s. GPT-4o P1 | 0.9500 | 8.76E-05 | 0.9969 | 5.37E-09 |
+| GPT-4o P3 v.s. GPT-4o P2 | 1.0000 | 0.00E+00 | 0.9995 | 8.56E-12 |
+| GPT-4o P3 v.s. GPT-4o P4 | 0.9500 | 8.76E-05 | 0.9998 | 8.32E-14 |
+| GPT-4o P3 v.s. GPT-4o P5 | 1.0000 | 0.00E+00 | 0.9999 | 5.61E-19 |
+| DS-671b P3 v.s. DS-671b P1 | 1.0000 | 0.00E+00 | 0.9999 | 1.51E-30 |
+| DS-671b P3 v.s. DS-671b P2 | 1.0000 | 0.00E+00 | 0.9999 | 4.95E-31 |
+| DS-671b P3 v.s. DS-671b P4 | 1.0000 | 0.00E+00 | 0.9999 | 2.93E-29 |
+| DS-671b P3 v.s. DS-671b P5 | 1.0000 | 0.00E+00 | 0.9999 | 1.08E-31 |
+| 1-Line vs. 2-Line | 0.9999 | 1.40E-24 | 0.9989 | 4.54E-05 |
+| 2-Line vs. 3-Line | 0.9999 | 1.40E-24 | 0.9995 | 1.46E-05 |
+| 1-Line vs. 3-Line | 0.9999 | 1.40E-24 | 0.9998 | 2.99E-06 |
+| QB-6a (Default) v.s. QB-6b | 1.0000 | 0.00E+00 | 0.9999 | 2.99E-15 |
+| QB-6a (Default) v.s. QB-3 | 1.0000 | 0.00E+00 | 0.9999 | 1.82E-14 |
+| QB-6a (Default) v.s. QB-4 | 1.0000 | 0.00E+00 | 0.9999 | 3.97E-15 |
+| QB-6a (Default) v.s. QB-5 | 1.0000 | 0.00E+00 | 0.9999 | 7.62E-14 |
+| QB-6a (Default) v.s. QB-7 | 1.0000 | 0.0E+00 | 0.9999 | 1.12E-14 |
+| QB-6a (Default) v.s. QB-8 | 1.0000 | 0.0E+00 | 0.9999 | 1.28E-13 |
+| QB-6a (Default) v.s. QB-9 | 1.0000 | 0.00E+00 | 0.9999 | 7.03E-15 |
+| QB-6a (Default) v.s. QB-10 | 1.0000 | 0.00E+00 | 0.9999 | 7.03E-15 |
+| QB-6a (Default) v.s D4J-6 | 1.0000 | 0.00E+00 | 0.9999 | 9.81E-17 |
 QB-N: Randomly samples N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b randomly samples 6 other examples. D4J-6 randomly samples 6 examples from Defects4J.
-To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches.
-To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance.
+To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutants for all the approaches.
 In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 15.
@@
 | Metric | QB-6a | QB-3 | QB-4 | QB-5 | QB-6b | QB-7 | QB-8 | QB-9 | QB-10 | D4J-6 |
-| Real Bug Detec | 0.917 | 0.933 | 0.883 | 0.883 | 0.917 | 0.900 | 0.900 | 0.917 | 0.867 | 0.900 |
-| Avg. Ochiai | 0.467 | 0.450 | 0.400 | 0.433 | 0.433 | 0.467 | 0.433 | 0.433 | 0.400 | 0.367 |
-| Coupling Rate | 0.136 | 0.128 | 0.126 | 0.127 | 0.132 | 0.112 | 0.113 | 0.121 | 0.136 | 0.139 |
-| Comp. Rate | 67.5% | 62.7% | 65.6% | 66.8% | 63.9% | 65.1% | 66.0% | 61.8% | 64.9% | 62.1% |
-| Dup.Mut.Rate | 13.7% | 12.8% | 12.6% | 9.4% | 12.8% | 11.4% | 13.2% | 13.8% | 13.6% | 13.9% |
+| Real Bug Detec | 0.933 | 0.900 | 0.917 | 0.933 | 0.917 | 0.900 | 0.883 | 0.933 | 0.933 | 0.933 |
+| Avg. Ochiai | 0.483 | 0.483 | 0.450 | 0.417 | 0.450 | 0.450 | 0.433 | 0.467 | 0.483 | 0.483 |
+| Coupling Rate | 0.379 | 0.379 | 0.385 | 0.380 | 0.373 | 0.362 | 0.348 | 0.367 | 0.352 | 0.388 |
+| Comp. Rate | 82.1% | 82.4% | 80.8% | 81.7% | 81.6% | 81.4% | 81.8% | 82.2% | 81.4% | 83.1% |
+| Dup.Mut.Rate | 5.3% | 5.6% | 6.4% | 5.6% | 5.8% | 5.6% | 7.5% | 6.4% | 6.2% | 5.8% |
 | Avg. # Token | 106 | 98 | 102 | 103 | 106 | 108 | 109 | 110 | 113 | 107 |
 QB-N: Randomly samples N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b randomly samples 6 other examples. D4J-6 randomly samples 6 examples from Defects4J.
-To measure the impact of different few-shot examples on the results, we conduct comparative experiments with various examples.
+To evaluate how different few-shot examples influence the results, we conduct comparative experiments by varying the few-shot examples used in the default LLMut (GPT-4o) prompt.
 First, we randomly select 3, 4, 5, 6, 7, 8, 9, and 10 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples.
@@
 Therefore, despite variations in few-shot examples, there is no statistically significant evidence indicating superior performance among the prompts.
-4.1.4. Performance of Using the Same Number of Mutations
-In Section 3, we did not restrict the number of mutations generated by each approach.
-To study the performance of each approach under fixed quantity conditions, we followed the setting of the existing studies (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by BERT (Degiovanni and Papadakis, 2022), which is 11397.
-This ensures a fair comparison across approaches by controlling for differences in mutation counts, thereby eliminating the potential bias that methods generating more mutations might appear to perform better simply due to quantity rather than quality.
-For counts exceeding this number, we randomly selected the specified number of mutations for analysis.
+4.1.4. Performance of Using the Same Number of Mutants
+In Section 3, we did not restrict the number of mutants generated by each approach.
+To study the performance of each approach under fixed quantity conditions, we followed the setting of the existing studies (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutants generated by all methods to the minimum produced by BERT (Degiovanni and Papadakis, 2022), which is 11397.
+This ensures a fair comparison across approaches by controlling for differences in mutant counts, thereby eliminating the potential bias that methods generating more mutants might appear to perform better simply due to quantity rather than quality.
+For counts exceeding this number, we randomly selected the specified number of mutants for analysis.
 We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 18.
-Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
-| Name | R. B. Detec. | Ochiai Coeffi. | Coup. Rate | Comp. Rate | Dup. Rate |
+Under a fixed number of mutants, we observe that LLMut (DS-671b) continues to deliver the best performance in both real bug detection rate and average Ochiai coefficient, with LLMut (GPT-4o) following closely.
+LLMorpheus (DS-671b) and LLMorpheus (GPT-4o) achieve the highest coupling rates.
+In terms of validity metrics—compilability rate and duplicate mutant rate—BugFarm (DS-671b) performs the best among all LLM-based approaches.
+These findings indicate that, when controlling for the number of generated mutants, different LLM-based approaches (e.g., LLMut, LLMorpheus, and BugFarm) exhibit distinct strengths rather than a single universally superior approach.
+| Name | R. B. Detec. | Avg. Ochiai | Coup. Rate | Comp. Rate | Dup. Rate |
 | PIT | 31.6% | 12.6% | 17.4% | — | — |
@@
 | BugFarm (GPT-4o) | 58.6% | 17.7% | 47.9% | 74.2% | 7.5% |
+| BugFarm (DS-671b) | 58.9% | 17.9% | 48.0% | 77.4% | 4.6% |
 | LLMorpheus (GPT-4o) | 67.5% | 23.6% | 50.2% | 68.2% | 7.3% |
+| LLMorpheus (DS-671b) | 68.5% | 24.9% | 51.3% | 67.8% | 6.8% |
 | LLMut (GPT-3.5) | 82.2% | 26.5% | 45.6% | 61.2% | 9.9% |
@@
 | LLMut (GPT-4o-M) | 88.6% | 31.8% | 47.9% | 74.3% | 7.1% |
-| LLMut (DS-761b) | 89.2% | 39.4% | 49.4% | 76.3% | 7.7% |
+| LLMut (DS-671b) | 89.2% | 39.4% | 49.4% | 76.3% | 7.7% |
 | LLMut (DS-236b) | 87.0% | 34.1% | 47.9% | 74.6% | 7.6% |
@@
 | LLMut (CL-13b) | 68.5% | 16.9% | 43.6% | 69.5% | 38.3% |
-| Metric | LLMut (GPT-3.5) | LLMut (DS-761b) | ||||
-| Default | More Detailed | Default + 10 Err. | Default | More Detailed | Default + 10 Err. | |
-| Real Bug Detec | 0.917 | 0.900 | 0.900 | 0.950 | 0.950 | 0.933 |
-| Avg. Ochiai | 0.467 | 0.483 | 0.467 | 0.500 | 0.500 | 0.517 |
-| Coupling Rate | 0.136 | 0.129 | 0.117 | 16.9% | 17.8% | 16.0% |
-| Comp. Rate | 67.5% | 71.1% | 68.7% | 78.4% | 76.4% | 77.5% |
-| Dup.Mut.Rate | 13.7% | 14.1% | 10.4% | 5.1% | 4.1% | 6.2% |
-Note: The prompt achieving the best performance in terms of a metric is highlighted in grey background color.
-4.2. Impact of Prompt Turbulence on LLM-Based Mutation Generation
-In software engineering practice, even when using the same information sources, users may design prompts differently, for example, varying the level of detail in the requirements description or introducing minor language errors.
-To investigate the impact of such variations, we derived two prompt variants from the default prompt shown in Figure 1.
-More Detailed Output Specification: replacing the output specifications of the default prompt by explicitly requiring: Prohibit generating identical mutants, i.e., ensure ‘aftercode’ ‘precode’, and all mutants are unique.
-Write all output mutants in a JSON file.
-Default + 10 Errors: randomly injecting 10 English grammar mistakes into the default prompt, such as subject–verb agreement errors (e.g., it are) and noun number errors (e.g., two mutation).
-Following the same experimental setup as before, we compared these two variants using GPT-3.5 and DeepSeek-V3-761b, with the results summarized in Table 19.
-For GPT-3.5, the More Detailed Output Specification variant slightly improves the compilation rate (+3.6%) but also increases the duplicate mutation rate.
-In contrast, the Default + 10 Errors variant reduces real bug detection and coupling rate compared to the default prompt.
-Interestingly, the Default + 10 Errors variant, presumably the lowest in prompt quality, achieves the best performance on duplicate mutation rate.
-For DeepSeek-V3-761b, the differences across variants are minimal.
-For instance, the default and the More Detailed Output Specification prompts achieve identical real bug detection and average Ochiai, and the More Detailed Output Specification variant even shows a small increase in average Ochiai.
-Overall, these results suggest that prompt wording has a certain degree of randomness in its effect on mutation generation performance.
-To further verify this, we conducted similarity analysis across the three prompts for each metric and found no statistically significant superiority of any one prompt over the other two.
-| Metric | LLMut (GPT-3.5) | LLMut (GPT-4o) | LLMut (DS-671b) | |||
-| FOM | HOM | FOM | HOM | FOM | HOM | |
-| Real Bug Detec. | 0.917 | 0.900 | 0.933 | 0.933 | 0.950 | 0.933 |
-| Avg. Ochiai | 0.467 | 0.483 | 0.467 | 0.483 | 0.500 | 0.483 |
-| Coupling Rate | 0.136 | 0.249 | 0.151 | 0.307 | 0.169 | 0.306 |
-| Compilability Rate | 67.5% | 63.4% | 76.3% | 70.2% | 78.4% | 73.3% |
-| Dup. Mut. Rate | 13.7% | 5.2% | 5.3% | 3.5% | 5.1% | 3.1% |
-Note: The order achieving the best performance in terms of a metric is highlighted in grey background color.
-4.3. Effectiveness on High-Order Mutation Generation
-Previously, we only evaluated LLMs in generating first-order mutations (FOM).
-To explore the effectiveness of LLMs in generating high-order mutations (HOMs) (Jia and Harman, 2009; Wong et al., 2020), we conducted a preliminary experiment by evaluating the recent mainstream models, including GPT-3.5, GPT-4o, and DeepSeek-V3-761b.
-We modified the original prompt in Figure 1 to explicitly request HOM generation.
+| Metric | LLMut (GPT-4o) | LLMut (DS-671b) | ||
+| FOM | HOM | FOM | HOM | |
+| Real Bug Detec. | 0.933 | 0.933 | 0.950 | 0.933 |
+| Avg. Ochiai | 0.483 | 0.483 | 0.500 | 0.483 |
+| Coupling Rate | 0.379 | 0.507 | 0.377 | 0.506 |
+| Compilability Rate | 82.1% | 70.2% | 84.1% | 73.3% |
+| Dup. Mut. Rate | 5.3% | 3.5% | 5.1% | 3.1% |
+Note: The order achieving the best performance in terms of a metric is highlighted in a grey background color.
+4.2. Effectiveness on High-Order Mutant Generation
+Previously, we only evaluated LLMs in generating first-order mutants (FOMs).
+To explore the effectiveness of LLMs in generating high-order mutants (HOMs) (Jia and Harman, 2009; Wong et al., 2020), we conducted a preliminary experiment by evaluating LLMut on the recent mainstream models, including GPT-4o and DS-671b.
+We modified the original prompt to explicitly request HOM generation, shown as Figure 4. The prompts directly instruct LLMs to generate HOMs, and each HOM is composed of several first-order mutants.
 Table 19 shows the results.
@@
 However, FOMs achieve higher real bug detectability on all models.
-For the validity metrics, HOMs exhibit lower duplicate mutation rates, suggesting improved mutation diversity.
+For the validity metrics, HOMs exhibit lower duplicate mutant rates, suggesting improved mutant diversity.
 However, FOMs exhibit significantly higher compilability rates, which aligns with the expectation that larger code changes are more prone to introducing compilation errors.
@@
 4.3. Implications
-Based on our findings, we discuss the pros and cons of traditional and LLM-based mutation generation approaches and suggest potential improvements.
+Based on our findings, we discuss the pros and cons of traditional and LLM-based mutant generation approaches and suggest potential improvements.
 4.3.1. Advantages of LLM-based Approaches
-Our study demonstrates that LLMs can generate high-quality mutations for mutation testing at an acceptable cost.
-The key advantage of LLMs lies in their ability to produce mutations with stronger coupling relationships with real-world bugs, which is crucial for improving test effectiveness.
-Additionally, LLMs could generate syntactically diverse mutations compared with traditional approaches, which is potentially the reason for generating high-coupling mutations.
-For example, Listing 1 presents two compilable mutations generated by GPT-3.5 for Chart-1, which are theoretically beyond the capabilities of traditional approaches.
-In Mutation-1, the LLM infers a relationship between variables (rightBlock implies the existence of leftBlock) and modifies the variable accordingly.
-This mutation is not killed by the original bug-triggering test, highlighting its potential for guiding test enhancement.
-Without natural language understanding (e.g., recognizing the correspondence between “right” and “left”), traditional approaches, both rule-based approaches and the approaches based on small-scale models, struggle to generate such context-aware mutations.
-In Mutation-2, the LLM transforms an else branch into an else if condition by synthesizing a logical condition using the integer variable seriesCount from the context, demonstrating an ability to introduce structurally meaningful changes that traditional approaches cannot achieve.
-Additionally, our study explores how different prompt designs, prompt turbulence, model choices, and contextual information influence mutation generation performance.
+Our study demonstrates that LLMs can generate high-quality mutants for mutation testing at an acceptable cost.
+The key advantage of LLMs lies in their ability to produce mutants with stronger coupling relationships with real-world bugs, which is crucial for improving test effectiveness.
+Additionally, LLMs could generate syntactically diverse mutants compared with traditional approaches, which is potentially the reason for generating high-coupling mutants.
+For example, Listing 1 presents two compilable mutants generated by GPT for Chart-1, which are theoretically beyond the capabilities of traditional approaches.
+In Mutant-1, the LLM infers a relationship between variables (rightBlock implies the existence of leftBlock) and modifies the variable accordingly.
+This mutant is not killed by the original bug-triggering test, highlighting its potential for guiding test enhancement.
+Without natural language understanding (e.g., recognizing the correspondence between “right” and “left”), traditional approaches, both rule-based approaches and the approaches based on small-scale models, struggle to generate such context-aware mutants.
+In Mutant-2, the LLM transforms an else branch into an else if condition by synthesizing a logical condition using the integer variable seriesCount from the context, demonstrating an ability to introduce structurally meaningful changes that traditional approaches cannot achieve.
+Additionally, our study explores how different prompt designs, model choices, and contextual information influence mutant generation performance.
 We find that selecting models with stronger code generation capabilities (e.g., GPT-4o) and designing prompts carefully are critical to achieving optimal results.
 Moreover, our study demonstrates that LLMs have clear potential in generating HOMs.
-Compared with first-order mutations (FOMs), HOMs generally achieve higher coupling rates and average Ochiai scores, suggesting that they are semantically closer to real bugs, while suffering from lower compilability rates and reduced real bug detectability.
+Compared with first-order mutants (FOMs), HOMs generally achieve higher coupling rates and average Ochiai scores, suggesting that they are semantically closer to real bugs, while suffering from lower compilability rates and reduced real bug detectability.
 The results highlight that strengthening LLM-based HOM generation is therefore a promising research direction.
 4.3.2. Advantages of Traditional Approaches
-Despite the advantages of LLM-generated mutations, traditional approaches still exhibit strengths in reliability and efficiency.
-They produce fewer non-compilable and duplicate mutations, with PIT achieving a 100% compilability rate and a 0% duplicate mutation rate, while Major follows closely with a 97.6% compilability rate and 0% duplicate mutation rate.
-In contrast, LLM-based approaches exhibit significantly lower compilability and higher duplicate mutation rates.
-Additionally, overall, traditional rule-based approaches generate a slightly lower equivalent mutation rate, further avoiding redundant costs.
+Despite the advantages of LLM-generated mutants, traditional approaches still exhibit strengths in reliability and efficiency.
+They produce fewer non-compilable and duplicate mutants, with PIT achieving a 100% compilability rate and a 0% duplicate mutant rate, while Major follows closely with a 97.6% compilability rate and 0% duplicate mutant rate.
+In contrast, LLM-based approaches exhibit significantly lower compilability and higher duplicate mutant rates.
+Additionally, overall, traditional rule-based approaches generate a slightly lower equivalent mutant rate, further avoiding redundant costs.
 This suggests that traditional approaches remain highly practical when strict compilability and low redundancy are prioritized.
@@
 Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate that there is still substantial room for improvement.
-One major limitation is the high rate of non-compilable mutations, which limits their practical application in mutation testing.
+One major limitation is the high rate of non-compilable mutants, which limits their practical application in mutation testing.
 Our analysis reveals that common causes include syntactic errors in generated expressions, incorrect method invocations, and type mismatches.
-To address these challenges, future research should explore the integration of LLM-based approaches with program analysis and repair techniques to refine mutation generation.
-Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutations before execution can enhance their reliability.
-Moreover, exploring the applications of LLM-generated mutation in the downstream tasks, such as mutation-based test case prioritization (MBTCP) and mutation-based fault localization (MBFL).
-Our study has shown that LLM-generated mutations often have a higher similarity with real bugs (e.g., have larger Ochiai coefficients), highlighting the potential of adopting them for these tasks.
+To address these challenges, future research should explore the integration of LLM-based approaches with program analysis and repair techniques to refine mutant generation.
+Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutants before execution can enhance their reliability.
+Moreover, exploring the applications of LLM-generated mutants in the downstream tasks, such as mutation-based test case prioritization (MBTCP) and mutation-based fault localization (MBFL).
+Our study has shown that LLM-generated mutants often have a higher similarity with real bugs (e.g., have larger Ochiai coefficients), highlighting the potential of adopting them for these tasks.
 Another important direction is cross-language generalization.
@@
 dynamically typed).
-Advancing towards multilingual mutation generation would not only broaden the applicability of LLMs in software testing but also provide stronger evidence for their robustness in heterogeneous, real-world development environments.
+Advancing towards multilingual mutant generation would not only broaden the applicability of LLMs in software testing but also provide stronger evidence for their robustness in heterogeneous, real-world development environments.
 4.4. Limitations of Our Study
@@
 First, while we evaluated a comprehensive set of static and behavioral metrics, we did not investigate more advanced dynamic indicators.
-For instance, the phenomenon of subsuming mutations (Garg et al., 2022; Diniz et al., 2021), where one mutation is always killed by any test that kills another, has been studied in the mutation testing literature as a way to assess better redundancy and the true fault-detection power of test suites.
-Incorporating such dynamic analyses would provide deeper insights into the effectiveness of LLM-generated mutations beyond the metrics reported in this paper.
-Additionally, while in the context of test suite enhancement, two mutants with identical failing test sets may be considered redundant, in mutation-based fault localization (Wang et al., 2025), the number of mutants per statement directly influences suspiciousness scores, since these are typically computed as averages across all associated mutants.
+For instance, the phenomenon of subsuming mutants (Garg et al., 2022; Diniz et al., 2021), where one mutant is always killed by any test that kills another, has been studied in the mutant testing literature as a way to assess better redundancy and the true fault-detection power of test suites.
+Incorporating such dynamic analyses would provide deeper insights into the effectiveness of LLM-generated mutants beyond the metrics reported in this paper.
+Additionally, while in the context of test suite enhancement, two mutants with identical failing test sets may be considered redundant, in mutation-based fault localization (Wang et al., 2025a), the number of mutants per statement directly influences suspiciousness scores, since these are typically computed as averages across all associated mutants.
 Second, our experiments are limited to the Java programming language.
 Although Java is the most widely studied language in mutation testing, it does not cover the diversity of programming paradigms.
-Mutation behaviors may differ significantly in languages such as C/C++, Python, JavaScript, and Rust, due to differences in type systems, memory models, and runtime environments.
-Future research should examine the generalizability of LLM-based mutation generation across multiple languages.
-Notably, modern LLMs are inherently multilingual, trained on large-scale cross-language corpora, which makes cross-language mutation generation a promising direction for broadening the applicability of our findings.
-Third, our study relies on fixed few-shot examples, which restricts LLMs to producing context-specific mutations for the given code snippets.
+Mutant behaviors may differ significantly in languages such as C/C++, Python, JavaScript, and Rust, due to differences in type systems, memory models, and runtime environments.
+Future research should examine the generalizability of LLM-based mutant generation across multiple languages.
+Notably, modern LLMs are inherently multilingual, trained on large-scale cross-language corpora, which makes cross-language mutant generation a promising direction for broadening the applicability of our findings.
+Third, our study relies on a fixed few-shot examples, which restricts LLMs to producing context-specific mutants for the given code snippets.
+Moreover, constructing few-shot examples requires manual effort and may introduce unintended biases into the model’s behavior during inference.
 In addition, we impose no restrictions on the mutation operators applied by LLMs, whereas rule-based approaches rely on predefined and relatively simple operators.
@@
 In this section, we illustrate the threats to the validity of our study.
-The selected programming language, mutation generation approaches, datasets, and LLMs could be a threat to the validity of our results.
-To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek).
-Moreover, our study further adopts two state-of-the-art LLM-based mutation generation approaches (i.e., BugFarm and LLMorpheus) as baselines.
+The selected programming language, mutant generation approaches, datasets, and LLMs could be a threat to the validity of our results.
+To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT models, StarCoder, CodeLlama, and DeepSeek models).
+Moreover, our study further covers two state-of-the-art LLM-based mutant generation approaches, BugFarm and LLMorpheus, as baselines, both of which are equipped with the recent mainstream models GPT-4o and DS-671b.
 The metrics we used may threaten the conclusion of our study.
@@
 Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014b) may be covered in the training set of the studied LLMs.
-To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024), which includes programs and faults that were made after the release time of the most LLMs we use (expect DeepSeek-V3-761b only) and thus have limited data leakage risk.
+To mitigate this threat, we employed another dataset, ConDefects (Wu et al., 2024), which includes programs and faults that were made after the release time of most LLMs we use (except DeepSeek-V3-671b only) and thus have limited data leakage risk.
 Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults.
-We hypothesize that, in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate.
-On Defects4J, the results are GPT-3.5 (133), GPT-4o (63), GPT-4o-Mini (138), DeepSeek-V3-761b (72), DeepSeek-V2-236b (66), BugFarm (59), LLMorpheus (26), StarChat (13), CodeLlama (28), LEAM (287), Bert (12), and Major (22), respectively.
-On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek-V3-761b (35), DeepSeek-V2-236b (32), BugFarm (22), LLMorpheus (19), StarChat (35), CodeLlama (56), LEAM (56), Bert (5), and Major (92), respectively.
-The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches.
-Interestingly, Major produces a similar number of exact matches with Codellama.
-Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except for GPTs and LEAM) due to exact matches (in the case of Defects4J).
+We hypothesize that, in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutant that is an exact match with the faults we investigate.
+On Defects4J, the results are LLMut (GPT-3.5) (133), LLMut (GPT-4o) (63), LLMut (GPT-4o-M) (138), LLMut (DS-671b) (72), LLMut (DS-236b) (66), LLMut (SC-16b) (13), LLMut (CL-13b) (28), BugFarm (GPT-4o) (59), BugFarm (DS-671b) (66), LLMorpheus (GPT-4o) (26), LLMorpheus (DS-671b) (31), LEAM (287), Bert (12), and Major (22), respectively.
+On ConDefects, the results are LLMut (GPT-3.5) (22), LLMut (GPT-4o) (25), LLMut (GPT-4o-M) (31), LLMut (DS-671b) (35), LLMut (DS-236b) (32), LLMut (SC-16b) (35), LLMut (CL-13b) (56), BugFarm (GPT-4o) (22), BugFarm (DS-671b) (29), LLMorpheus (GPT-4o) (19), LLMorpheus (DS-671b) (28), LEAM (56), Bert (5), and Major (92), respectively.
+The comparison indicates that on Defects4J, LEAM tends to produce significantly more exact matches than the other approaches.
+Interestingly, Major produces a similar number of exact matches with LLMut (CL-13b).
+Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except for LLM-based approaches with the GPT models and LEAM) due to exact matches (in the case of Defects4J).
 Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of exact matches, indicating a minor influence of any data leakage on the reported results.
-Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with -value less than 0.05, indicating their performance is similar on the two datasets.
+Nevertheless, the LLMs we studied exhibit the same trend on both datasets, achieving a Spearman coefficient of 0.943 and a Pearson correlation of 0.944, both with a -value less than 0.05, indicating their performance is similar on both datasets.
 The different experimental settings may also threaten the validity of our results.
-To address the threat, we explore the impacts of prompt design, prompt turbulence, context length, few-shot examples, and mutation numbers on the performance of LLMs.
+To address the threat, we explore the impacts of prompt design, context length, few-shot examples, and mutant numbers on the performance of LLMs.
 The results show that different settings are highly similar.
@@
 However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs.
-The subjective nature of human decisions in labeling for equivalent mutations, non-compilation errors, and alive mutations is another potential threat.
-To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutation.
-The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two annotators.
+The subjective nature of human decisions in labeling for equivalent mutants, non-compilation errors, and surviving mutants is another potential threat.
+To mitigate this threat, we follow a rigorous annotation process where two co-authors independently annotated each mutant.
+The final Cohen’s Kappa coefficient indicates a relatively high level of agreement between the two authors.
 6. Related Work
 In this section, we introduce the related work of our study.
-6.1. Mutation Testing and Mutation Generation
-Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites.
-It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test.
+6.1. Mutation Testing and Mutant Generation
+Mutation testing is a systematic testing approach that aims to guide the generation of high-utility test suites.
+It works by introducing syntactic modifications (i.e., mutants) into the source code of the programs under test.
 By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978), one can identify weaknesses of their test suites, i.e., mutants that are not killed.
-Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017).
-Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014c; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
-All these tasks require high-utility mutations.
-This is because it is required to compile and execute every mutation with tests, which is computationally expensive.
-Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022).
-Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019).
-Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to be coupled with real faults/bugs to allow these techniques to be effective.
-Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to as rule-based.
+Besides testing, mutation analysis is a tool that can support multiple applications in software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017).
+Most commonly, mutants are employed as substitutes for real-world bugs (Just et al., 2014c; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025a), test prioritization (Shin et al., 2019) and program repair (Hanna et al., 2025; Xiao et al., 2023, 2024).
+All these tasks require high-utility mutants.
+This is because it is required to compile and execute every mutant with tests, which is computationally expensive.
+Therefore, generating redundant or non-compilable mutants, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2016, 2018; Wang et al., 2017, 2021; Garg et al., 2022).
+Additionally, when mutants are used as substitutes for real defects, the mutants must be as syntactically close to the real bugs as possible, making the mutants more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019).
+Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of mutants must be coupled with real faults/bugs to make these techniques effective.
+Traditional mutation testing approaches generate mutants by pre-defined program transformation rules (i.e., mutation operators), which are referred to as rule-based.
 For example, one widely used mutation operator is to change one arithmetic binary operator with another one (e.g., ).
 During the last two decades, the majority of existing mutation testing approaches have been rule-based (Jia and Harman, 2010; Amalfitano et al., 2022).
-Researchers primarily focus on approaches in Java (e.g., PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), IBIR (Khanfir et al., 2023), etc.) and C/C++ (e.g., Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), WinMut (Wang et al., 2021), etc.), leveraging the mature ecosystems and tooling chains of these languages for mutation testing.
-Recently, with the emergence of machine learning techniques, researchers proposed using deep learning models to generate mutations.
+Researchers primarily focus on approaches in Java (e.g., PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), IBIR (Khanfir et al., 2023), MIN (Zhang et al., 2025), etc.) and C/C++ (e.g., Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), WinMut (Wang et al., 2021), etc.), leveraging the mature ecosystems and tooling chains of these languages for mutation testing.
+Recently, with the emergence of machine learning techniques, researchers proposed using deep learning models to generate mutants.
 For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion.
@@
 d.]).
-However, these approaches suffer from the challenges in generating syntactically correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations.
+However, these approaches suffer from challenges in generating syntactically correct mutants and mutants with diverse forms, often failing to produce syntactically valid mutants.
 Perhaps the closest to our work is that of Degiovanni et al., who proposed BERT (Degiovanni and Papadakis, 2022).
-BERT generates mutation by masking the code elements to be mutated and then predicting the mutation via BERT (Devlin et al., 2018).
-Several recent studies have also explored the use of LLMs for mutation generation.
+BERT generates mutants by masking the code elements to be mutated and then predicting the mutants via BERT (Devlin et al., 2018).
+Several recent studies have also explored the use of LLMs for mutant generation.
 Tip et al.
-concurrently investigated the use of LLMs for mutation generation in JavaScript (Tip et al., 2025), primarily focusing on the feasibility of leveraging LLMs to produce mutations.
+concurrently investigated the use of LLMs for mutant generation in JavaScript (Tip et al., 2025), primarily focusing on the feasibility of leveraging LLMs to produce mutants.
 In contrast, our study conducts a broader comparison by evaluating LLMs against a diverse set of baselines, including both rule-based and learning-based approaches.
@@
 Similarly, Endres et al.
-employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al.
+employed large models for mutant generation to aid deduction (Endres et al., 2024), while Deb et al.
 used GPT as one of the baselines in their approach (DEB et al., 2024).
@@
 (Dakhel et al., 2024) and Ibrahimzada et al.
-(Ibrahimzada et al., 2023) also proposed generating artificial bugs using LLMs.
+(Ibrahimzada et al., 2025) also proposed generating artificial bugs using LLMs.
 These approaches mainly evaluate from a certain perspective, such as mutation score or program repair.
-BugFarm (Ibrahimzada et al., 2023) intends to generate artificial bugs that are hard to kill, by first choosing code locations based on a trained small model and then mutating the code via an LLM.
-Compared with these approaches, our study is more comprehensive and evaluates existing mutation generation approaches from many aspects.
-Specifically, we measure the behavioral similarity between real bugs and mutations using established metrics, evaluate the syntactic diversity of mutations, and analyze the non-compilable mutations produced by LLMs.
-Our findings illustrate both the strengths and limitations of LLMs for mutation generation, and highlight directions for future improvement.
+BugFarm (Ibrahimzada et al., 2025) intends to generate artificial bugs that are hard to kill, by first choosing code locations based on a trained small model and then mutating the code via an LLM.
+Compared with these approaches, our study is more comprehensive and evaluates existing mutant generation approaches from many aspects.
+Specifically, we measure the behavioral similarity between real bugs and mutants using established metrics, evaluate the syntactic diversity of mutants, and analyze the non-compilable mutants produced by LLMs.
+Our findings illustrate both the strengths and limitations of LLMs for mutant generation, and highlight directions for future improvement.
 6.2. LLMs for Code
-Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023b).
+Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023; Schäfer et al., 2023).
 Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes.
@@
 7. Conclusion
-In this paper, we systematically investigate the performance of existing mutation generation appraoches across multiple dimensions, including their effectiveness, validity, and efficiency.
-Our study conducts extensive and detailed comparative experiments, covering both widely studied rule-based mutation generation approaches and LLM-based methods.
-Particularly, we evaluate not existing LLM-based approaches but also our new prompt for mutation generation.
-Our evaluation results highlight that LLMs significantly outperform traditional approaches in generating diverse mutations that closely mimic real bugs.
-However, the results also expose key limitations of LLMs, particularly their tendency to generate a high proportion of non-compilable mutations.
-Therefore, we analyze the error types of the LLM-generated non-compilable mutations and identify that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
+In this paper, we systematically investigate the performance of existing mutant generation approaches across multiple dimensions, including their effectiveness, validity, and efficiency.
+Our study conducts extensive and detailed comparative experiments, covering both widely studied rule-based mutant generation approaches and LLM-based methods.
+Particularly, we evaluate not only existing LLM-based approaches but also our new prompt for mutant generation.
+Our evaluation results highlight that LLMs significantly outperform traditional approaches in generating diverse s that closely mimic real bugs.
+However, the results also expose key limitations of LLMs, particularly their tendency to generate a high proportion of non-compilable mutants.
+Therefore, we analyze the error types of the LLM-generated non-compilable mutants and identify that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutants.
 Based on our findings, we emphasize the need for further research to enhance LLM-based mutation testing, addressing its current shortcomings while leveraging its strengths.
+Our implementation and the experimental data are available at our artifact: https://github.com/da-123-snake/Mutantion-Generation-Study.
+Acknowledgements.
+We are grateful to the anonymous reviewers. This work was supported by the National Natural Science Foundation of China under Grant No. 62202040.References
+- (1)
+- Achiam et al.
+(2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al.
+2023.
+Gpt-4 technical report.
+arXiv preprint arXiv:2303.08774 (2023).
+- Amalfitano et al. (2022) Domenico Amalfitano, Ana CR Paiva, Alexis Inquel, Luís Pinto, Anna Rita Fasolino, and René Just. 2022. How do Java mutation tools differ? Commun. ACM 65, 12 (2022), 74–89.
+- Andrews et al.
+(2005) James H Andrews, Lionel C Briand, and Yvan Labiche.
+2005.
+Is mutation an appropriate tool for testing experiments?.
+In Proceedings of the 27th international conference on Software engineering.
+402–411.
+- Beller et al.
+(2021) Moritz Beller, Chu-Pan Wong, Johannes Bader, Andrew Scott, Mateusz Machalica, Satish Chandra, and Erik Meijer.
+2021.
+What it would take to use mutation testing in industry—a study at facebook.
+In 2021 IEEE/ACM 43rd International Conference on Software Engineering: Software Engineering in Practice (ICSE-SEIP).
+IEEE, 268–277.
+- Brown et al.
+(2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al.
+2020.
+Language models are few-shot learners.
+Advances in neural information processing systems 33 (2020), 1877–1901.
+- Budd and Angluin (1982) Timothy A Budd and Dana Angluin. 1982. Two notions of correctness and their relation to testing. Acta informatica 18, 1 (1982), 31–45.
+- Chekam et al.
+(2020) Thierry Titcheu Chekam, Mike Papadakis, Tegawendé F.
+Bissyandé, Yves Le Traon, and Koushik Sen.
+2020.
+Selecting fault revealing mutants.
+Empir.
+Softw.
+Eng.
+25, 1 (2020), 434–487.
+- Chen et al.
+(2025) Xiangping Chen, Xing Hu, Yuan Huang, He Jiang, Weixing Ji, Yanjie Jiang, Yanyan Jiang, Bo Liu, Hui Liu, Xiaochen Li, Xiaoli Lian, Guozhu Meng, Xin Peng, Hailong Sun, Lin Shi, Bo Wang, Chong Wang, Jiayi Wang, Tiantian Wang, Jifeng Xuan, Xin Xia, Yibiao Yang, Yixin Yang, Li Zhang, Yuming Zhou, and Lu Zhang.
+2025.
+Deep Learning-based Software Engineering: Progress, Challenges, and Opportunities.
+Science China Information Sciences 68, 1 (2025), 1–88.
+- Coles et al.
+(2016) Henry Coles, Thomas Laurent, Christopher Henard, Mike Papadakis, and Anthony Ventresque.
+2016.
+Pit: a practical mutation testing tool for java.
+In Proceedings of the 25th international symposium on software testing and analysis.
+449–452.
+- Dakhel et al.
+(2024) Arghavan Moradi Dakhel, Amin Nikanjam, Vahid Majdinasab, Foutse Khomh, and Michel C Desmarais.
+2024.
+Effective test generation using pre-trained large language models and mutation testing.
+Information and Software Technology 171 (2024), 107468.
+- Daran and Thévenod-Fosse (1996) Murial Daran and Pascale Thévenod-Fosse.
+1996.
+Software error analysis: A real case study involving real faults and mutations.
+ACM SIGSOFT Software Engineering Notes 21, 3 (1996), 158–171.
+- de Mello et al.
+(2015) Rafael Maiani de Mello, Pedro Corrêa Da Silva, and Guilherme Horta Travassos.
+2015.
+Investigating probabilistic sampling approaches for large-scale surveys in software engineering.
+Journal of Software Engineering Research and Development 3 (2015), 1–26.
+- DEB et al. (2024) SOURAV DEB, KUSH JAIN, RIJNARD VAN TONDER, CLAIRE LE GOUES, and ALEX GROCE. 2024. Syntax Is All You Need: A Universal-Language Approach to Mutant Generation. (2024).
+- Degiovanni and Papadakis (2022) Renzo Degiovanni and Mike Papadakis.
+2022.
+bert: Mutation testing using pre-trained language models.
+In 2022 IEEE International Conference on Software Testing, Verification and Validation Workshops (ICSTW).
+IEEE, 160–169.
+- Delamaro et al. (1996) Márcio Eduardo Delamaro, José Carlos Maldonado, and A Mathur. 1996. Proteum-a tool for the assessment of test adequacy for c programs user’s guide. In PCS, Vol. 96. 79–95.
+- DeMillo et al. (1978) Richard A DeMillo, Richard J Lipton, and Frederick G Sayward. 1978. Hints on test data selection: Help for the practicing programmer. Computer 11, 4 (1978), 34–41.
+- DeMillo et al.
+(1979) Richard A DeMillo, Richard J Lipton, and Frederick G Sayward.
+1979.
+Program mutation: A new approach to program testing.
+Infotech State of the Art Report, Software Testing 2, 1979 (1979), 107–126.
+- Deng et al.
+(2024) Yinlin Deng, Chunqiu Steven Xia, Chenyuan Yang, Shizhuo Dylan Zhang, Shujing Yang, and Lingming Zhang.
+2024.
+Large language models are edge-case generators: Crafting unusual programs for fuzzing deep learning libraries.
+In Proceedings of the 46th IEEE/ACM International Conference on Software Engineering.
+1–13.
+- Devlin et al.
+(2018) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova.
+2018.
+Bert: Pre-training of deep bidirectional transformers for language understanding.
+arXiv preprint arXiv:1810.04805 (2018).
+- Diniz et al.
+(2021) João P Diniz, Chu-Pan Wong, Christian Kästner, and Eduardo Figueiredo.
+2021.
+Dissecting strongly subsuming second-order mutants.
+In 2021 14th IEEE Conference on Software Testing, Verification and Validation (ICST).
+IEEE, 171–181.
+- Endres et al.
+(2024) Madeline Endres, Sarah Fakhoury, Saikat Chakraborty, and Shuvendu K Lahiri.
+2024.
+Can Large Language Models Transform Natural Language Intent into Formal Method Postconditions?
+Proceedings of the ACM on Software Engineering 1, FSE (2024), 1889–1912.
+- Falleri et al.
+(2014) Jean-Rémy Falleri, Floréal Morandat, Xavier Blanc, Matias Martinez, and Martin Monperrus.
+2014.
+Fine-grained and accurate source code differencing.
+In Proceedings of the 29th ACM/IEEE international conference on Automated software engineering.
+313–324.
+- Fan et al.
+(2023) Angela Fan, Beliz Gokkaya, Mark Harman, Mitya Lyubarskiy, Shubho Sengupta, Shin Yoo, and Jie M.
+Zhang.
+2023.
+Large Language Models for Software Engineering: Survey and Open Problems.
+In 2023 IEEE/ACM International Conference on Software Engineering: Future of Software Engineering (ICSE-FoSE).
+IEEE, Los Alamitos, CA, USA, 31–53.
+- Garg et al.
+(2022) Aayush Garg, Milos Ojdanic, Renzo Degiovanni, Thierry Titcheu Chekam, Mike Papadakis, and Yves Le Traon.
+2022.
+Cerebro: Static subsuming mutant selection.
+IEEE Transactions on Software Engineering 49, 1 (2022), 24–43.
+- Gay and Salahirad (2023) Gregory Gay and Alireza Salahirad.
+2023.
+How Closely are Common Mutation Operators Coupled to Real Faults?.
+In 2023 IEEE Conference on Software Testing, Verification and Validation (ICST).
+IEEE, 129–140.
+- Guo et al.
+(2024) Qi Guo, Junming Cao, Xiaofei Xie, Shangqing Liu, Xiaohong Li, Bihuan Chen, and Xin Peng.
+2024.
+Exploring the potential of chatgpt in automated code refinement: An empirical study.
+In Proceedings of the 46th IEEE/ACM International Conference on Software Engineering.
+1–13.
+- Hamlet (1977) Richard G. Hamlet. 1977. Testing programs with the aid of a compiler. IEEE transactions on software engineering 4 (1977), 279–290.
+- Hanna et al.
+(2025) Carol Hanna, Aymeric Blot, and Justyna Petke.
+2025.
+Reinforcement learning for mutation operator selection in automated program repair.
+Automated Software Engineering 32, 2 (2025), 1–33.
+- Hariri et al.
+(2019) Farah Hariri, August Shi, Vimuth Fernando, Suleman Mahmood, and Darko Marinov.
+2019.
+Comparing mutation testing at the levels of source code and compiler intermediate representation.
+In 2019 12th IEEE conference on software testing, validation and verification (ICST).
+IEEE, 114–124.
+- Harman et al.
+(2025) Mark Harman, Jillian Ritchey, Inna Harper, Shubho Sengupta, Ke Mao, Abhishek Gulati, Christopher Foster, and Hervé Robert.
+2025.
+Mutation-Guided LLM-based Test Generation at Meta.
+In Proceedings of the 33rd ACM International Conference on the Foundations of Software Engineering.
+180–191.
+- Hindle et al. (2016) Abram Hindle, Earl T Barr, Mark Gabel, Zhendong Su, and Premkumar Devanbu. 2016. On the naturalness of software. Commun. ACM 59, 5 (2016), 122–131.
+- Huang et al.
+(2023) Lei Huang, Weijiang Yu, Weitao Ma, Weihong Zhong, Zhangyin Feng, Haotian Wang, Qianglong Chen, Weihua Peng, Xiaocheng Feng, Bing Qin, et al.
+2023.
+A survey on hallucination in large language models: Principles, taxonomy, challenges, and open questions.
+arXiv preprint arXiv:2311.05232 (2023).
+- Ibrahimzada et al.
+(2025) Ali Reza Ibrahimzada, Yang Chen, Ryan Rong, and Reyhaneh Jabbarvand.
+2025.
+Challenging Bug Prediction and Repair Models with Synthetic Bugs.
+In 2025 IEEE International Conference on Source Code Analysis & Manipulation (SCAM).
+IEEE, 133–144.
+- Jia and Harman (2008) Yue Jia and Mark Harman.
+2008.
+MILU: A customizable, runtime-optimized higher order mutation testing tool for the full C language.
+In Testing: Academic & Industrial Conference-Practice and Research Techniques (taic part 2008).
+IEEE, 94–98.
+- Jia and Harman (2009) Yue Jia and Mark Harman. 2009. Higher order mutation testing. Information and Software Technology 51, 10 (2009), 1379–1393.
+- Jia and Harman (2010) Yue Jia and Mark Harman. 2010. An analysis and survey of the development of mutation testing. IEEE transactions on software engineering 37, 5 (2010), 649–678.
+- Jiang et al.
+(2023) Nan Jiang, Thibaud Lutellier, Yiling Lou, Lin Tan, Dan Goldwasser, and Xiangyu Zhang.
+2023.
+Knod: Domain knowledge distilled tree decoder for automated program repair.
+In 2023 IEEE/ACM 45th International Conference on Software Engineering (ICSE).
+IEEE, 1251–1263.
+- Jimenez et al.
+(2018) Matthieu Jimenez, Thiery Titcheu Checkam, Maxime Cordy, Mike Papadakis, Marinos Kintis, Yves Le Traon, and Mark Harman.
+2018.
+Are mutants really natural?
+a study on how” naturalness” helps mutant selection.
+In Proceedings of the 12th ACM/IEEE International Symposium on Empirical Software Engineering and Measurement.
+1–10.
+- Just et al.
+(2014a) René Just, Michael D Ernst, and Gordon Fraser.
+2014a.
+Efficient mutation analysis by propagating and partitioning infected execution states.
+In Proceedings of the 2014 international symposium on software testing and analysis.
+315–326.
+- Just et al.
+(2014b) René Just, Darioush Jalali, and Michael D Ernst.
+2014b.
+Defects4J: A database of existing faults to enable controlled testing studies for Java programs.
+In Proceedings of the 2014 International Symposium on Software Testing and Analysis.
+437–440.
+- Just et al.
+(2014c) René Just, Darioush Jalali, Laura Inozemtseva, Michael D Ernst, Reid Holmes, and Gordon Fraser.
+2014c.
+Are mutants a valid substitute for real faults in software testing?.
+In Proceedings of the 22nd ACM SIGSOFT International Symposium on Foundations of Software Engineering.
+654–665.
+- Just et al. (2011) Rene Just, Franz Schweiggert, and Gregory M Kapfhammer. 2011. MAJOR: An efficient and extensible tool for mutation analysis in a Java compiler. In ASE. 612–615.
+- Kaufman et al.
+(2022) Samuel J Kaufman, Ryan Featherman, Justin Alvin, Bob Kurtz, Paul Ammann, and René Just.
+2022.
+Prioritizing mutants to guide mutation testing.
+In Proceedings of the 44th International Conference on Software Engineering.
+1743–1754.
+- Khanfir et al.
+(2023) Ahmed Khanfir, Anil Koyuncu, Mike Papadakis, Maxime Cordy, Tegawendé F Bissyandé, Jacques Klein, and Yves Le Traon.
+2023.
+IBiR: Bug-report-driven fault injection.
+ACM Transactions on Software Engineering and Methodology 32, 2 (2023), 1–31.
+- Kim et al.
+(2022) Jinhan Kim, Juyoung Jeon, Shin Hong, and Shin Yoo.
+2022.
+Predictive mutation analysis via the natural language channel in source code.
+ACM Transactions on Software Engineering and Methodology (TOSEM) 31, 4 (2022), 1–27.
+- Kintis et al.
+(2017) Marinos Kintis, Mike Papadakis, Yue Jia, Nicos Malevris, Yves Le Traon, and Mark Harman.
+2017.
+Detecting trivial mutant equivalences via compiler optimisations.
+IEEE Transactions on Software Engineering 44, 4 (2017), 308–333.
+- Kushigian et al.
+(2024) Benjamin Kushigian, Samuel J Kaufman, Ryan Featherman, Hannah Potter, Ardi Madadi, and René Just.
+2024.
+Equivalent Mutants in the Wild: Identifying and Efficiently Suppressing Equivalent Mutants for Java Programs.
+In Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis.
+654–665.
+- Le Goues et al.
+(2011) Claire Le Goues, ThanhVu Nguyen, Stephanie Forrest, and Westley Weimer.
+2011.
+Genprog: A generic method for automatic software repair.
+Ieee transactions on software engineering 38, 1 (2011), 54–72.
+- Li et al.
+(2023a) Raymond Li, Loubna Ben Allal, Yangtian Zi, Niklas Muennighoff, Denis Kocetkov, Chenghao Mou, Marc Marone, Christopher Akiki, Jia Li, Jenny Chim, et al.
+2023a.
+Starcoder: may the source be with you!
+arXiv preprint arXiv:2305.06161 (2023).
+- Li et al.
+(2023b) Tsz-On Li, Wenxi Zong, Yibo Wang, Haoye Tian, Ying Wang, Shing-Chi Cheung, and Jeff Kramer.
+2023b.
+Nuances are the key: Unlocking chatgpt to find failure-inducing tests with differential prompting.
+In 2023 38th IEEE/ACM International Conference on Automated Software Engineering (ASE).
+IEEE, 14–26.
+- Lin et al.
+(2017) Derrick Lin, James Koppel, Angela Chen, and Armando Solar-Lezama.
+2017.
+QuixBugs: A multi-lingual program repair benchmark set based on the Quixey Challenge.
+In Proceedings Companion of the 2017 ACM SIGPLAN international conference on systems, programming, languages, and applications: software for humanity.
+55–56.
+- Linares-Vásquez et al.
+(2017) Mario Linares-Vásquez, Gabriele Bavota, Michele Tufano, Kevin Moran, Massimiliano Di Penta, Christopher Vendome, Carlos Bernal-Cárdenas, and Denys Poshyvanyk.
+2017.
+Enabling mutation testing for android apps.
+In Proceedings of the 2017 11th Joint Meeting on Foundations of Software Engineering.
+233–244.
+- Liu et al.
+(2024) Aixin Liu, Bei Feng, Bing Xue, Bingxuan Wang, Bochao Wu, Chengda Lu, Chenggang Zhao, Chengqi Deng, Chenyu Zhang, Chong Ruan, et al.
+2024.
+Deepseek-v3 technical report.
+arXiv preprint arXiv:2412.19437 (2024).
+- Ma et al. (2005) Yu-Seung Ma, Jeff Offutt, and Yong Rae Kwon. 2005. MuJava: an automated class mutation system. Software Testing, Verification and Reliability 15, 2 (2005), 97–133.
+- Ma et al.
+(2024) Zeyang Ma, An Ran Chen, Dong Jae Kim, Tse-Hsun Chen, and Shaowei Wang.
+2024.
+LLMParser: An Exploratory Study on Using Large Language Models for Log Parsing.
+In 2024 IEEE/ACM 46th International Conference on Software Engineering (ICSE).
+IEEE Computer Society, 883–883.
+- Macklon et al.
+(2023) Finlay Macklon, Markos Viggiato, Natalia Romanova, Chris Buzon, Dale Paas, and Cor-Paul Bezemer.
+2023.
+A Taxonomy of Testable HTML5 Canvas Issues.
+IEEE Transactions on Software Engineering 49, 6 (2023), 3647–3659.
+- Moon et al.
+(2014) Seokhyeon Moon, Yunho Kim, Moonzoo Kim, and Shin Yoo.
+2014.
+Ask the mutants: Mutating faulty programs for fault localization.
+In 2014 IEEE Seventh International Conference on Software Testing, Verification and Validation.
+IEEE, 153–162.
+- Motwani and Brun (2023) Manish Motwani and Yuriy Brun.
+2023.
+Better automatic program repair by using bug reports and tests together.
+In 2023 IEEE/ACM 45th International Conference on Software Engineering (ICSE).
+IEEE, 1225–1237.
+- Namin and Kakarla (2011) Akbar Siami Namin and Sahitya Kakarla.
+2011.
+The use of mutation in testing experiments and its sensitivity to external threats.
+In Proceedings of the 2011 International Symposium on Software Testing and Analysis.
+342–352.
+- Offutt and Untch (2001) A Jefferson Offutt and Roland H Untch. 2001. Mutation 2000: Uniting the orthogonal. Mutation testing for the new century (2001), 34–44.
+- Ojdanic et al.
+(2023) Milos Ojdanic, Aayush Garg, Ahmed Khanfir, Renzo Degiovanni, Mike Papadakis, and Yves Le Traon.
+2023.
+Syntactic vs.
+semantic similarity of artificial and real faults in mutation testing studies.
+IEEE Transactions on Software Engineering (2023).
+- Papadakis et al.
+(2015) Mike Papadakis, Yue Jia, Mark Harman, and Yves Le Traon.
+2015.
+Trivial compiler equivalence: A large scale empirical study of a simple, fast and effective equivalent mutant detection technique.
+In ICSE.
+936–946.
+- Papadakis et al.
+(2019) Mike Papadakis, Marinos Kintis, Jie Zhang, Yue Jia, Yves Le Traon, and Mark Harman.
+2019.
+Mutation testing advances: an analysis and survey.
+In Advances in Computers.
+Vol.
+112.
+Elsevier, 275–378.
+- Papadakis and Le Traon (2015) Mike Papadakis and Yves Le Traon. 2015. Metallaxis-FL: mutation-based fault localization. Software Testing, Verification and Reliability 25, 5-7 (2015), 605–628.
+- Patra and Pradel (2021) Jibesh Patra and Michael Pradel.
+2021.
+Semantic bug seeding: a learning-based approach for creating realistic bugs.
+In Proceedings of the 29th ACM Joint Meeting on European Software Engineering Conference and Symposium on the Foundations of Software Engineering.
+906–918.
+- Petrović and Ivanković (2018) Goran Petrović and Marko Ivanković.
+2018.
+State of mutation testing at google.
+In Proceedings of the 40th international conference on software engineering: Software engineering in practice.
+163–171.
+- Petrović et al.
+(2021a) Goran Petrović, Marko Ivanković, Gordon Fraser, and René Just.
+2021a.
+Does mutation testing improve testing practices?.
+In 2021 IEEE/ACM 43rd International Conference on Software Engineering (ICSE).
+- Petrović et al.
+(2021b) Goran Petrović, Marko Ivanković, Gordon Fraser, and René Just.
+2021b.
+Practical mutation testing at scale: A view from google.
+IEEE Transactions on Software Engineering 48, 10 (2021), 3900–3912.
+- Potter et al.
+(2025) Hannah Potter, Ana CR Paiva, Domenico Amalfitano, Anna Rita Fasolino, Porfirio Tramontana, and René Just.
+2025.
+Evaluating the Impact of Scaffolding and Visualizations for Mutation Testing Exercises in Software Engineering Education.
+In Proceedings of the 33rd ACM International Conference on the Foundations of Software Engineering.
+823–834.
+- Roziere et al.
+(2023) Baptiste Roziere, Jonas Gehring, Fabian Gloeckle, Sten Sootla, Itai Gat, Xiaoqing Ellen Tan, Yossi Adi, Jingyu Liu, Tal Remez, Jérémy Rapin, et al.
+2023.
+Code llama: Open foundation models for code.
+arXiv preprint arXiv:2308.12950 (2023).
+- Schäfer et al.
+(2023) Max Schäfer, Sarah Nadi, Aryaz Eghbali, and Frank Tip.
+2023.
+An empirical evaluation of using large language models for automated unit test generation.
+IEEE Transactions on Software Engineering (2023).
+- Schuler and Zeller (2009) David Schuler and Andreas Zeller. 2009. Javalanche: efficient mutation testing for Java. In ESEC/FSE. 297–298.
+- Shi et al.
+(2019) August Shi, Jonathan Bell, and Darko Marinov.
+2019.
+Mitigating the effects of flaky tests on mutation testing.
+In Proceedings of the 28th ACM SIGSOFT International Symposium on Software Testing and Analysis.
+112–122.
+- Shin et al.
+(2019) Donghwan Shin, Shin Yoo, Mike Papadakis, and Doo-Hwan Bae.
+2019.
+Empirical evaluation of mutation-based test case prioritization techniques.
+Software Testing, Verification and Reliability 29, 1-2 (2019), e1695.
+- Siami Namin et al.
+(2008) Akbar Siami Namin, James H Andrews, and Duncan J Murdoch.
+2008.
+Sufficient mutation operators for measuring test effectiveness.
+In Proceedings of the 30th international conference on Software engineering.
+351–360.
+- Tian et al.
+([n.
+d.]) Zhao Tian, Junjie Chen, Dong Wang, Qihao Zhu, Xingyu Fan, and Lingming Zhang.
+[n.
+d.].
+LEAM++: Learning for Selective Mutation Fault Construction.
+ACM Transactions on Software Engineering and Methodology ([n.
+d.]).
+- Tian et al.
+(2022) Zhao Tian, Junjie Chen, Qihao Zhu, Junjie Yang, and Lingming Zhang.
+2022.
+Learning to construct better mutation faults.
+In Proceedings of the 37th IEEE/ACM International Conference on Automated Software Engineering.
+1–13.
+- Tian et al.
+(2024) Zhao Tian, Honglin Shu, Dong Wang, Xuejie Cao, Yasutaka Kamei, and Junjie Chen.
+2024.
+Large Language Models for Equivalent Mutant Detection: How Far Are We?.
+In Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis.
+1733–1745.
+- Tip et al. (2025) Frank Tip, Jonathan Bell, and Max Schäfer. 2025. Llmorpheus: Mutation testing using large language models. IEEE Transactions on Software Engineering (2025).
+- Touvron et al.
+(2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al.
+2023.
+Llama: Open and efficient foundation language models.
+arXiv preprint arXiv:2302.13971 (2023).
+- Tufano et al.
+(2019) Michele Tufano, Cody Watson, Gabriele Bavota, Massimiliano Di Penta, Martin White, and Denys Poshyvanyk.
+2019.
+Learning how to mutate source code from bug-fixes.
+In 2019 IEEE International conference on software maintenance and evolution (ICSME).
+IEEE, 301–312.
+- Vaswani et al.
+(2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Łukasz Kaiser, and Illia Polosukhin.
+2017.
+Attention is all you need.
+Advances in neural information processing systems 30 (2017).
+- Viera et al. (2005) Anthony J Viera, Joanne M Garrett, et al. 2005. Understanding interobserver agreement: the kappa statistic. Fam med 37, 5 (2005), 360–363.
+- Wang et al.
+(2021) Bo Wang, Sirui Lu, Yingfei Xiong, and Feng Liu.
+2021.
+Faster mutation analysis with fewer processes and smaller overheads.
+In 2021 36th IEEE/ACM International Conference on Automated Software Engineering (ASE).
+IEEE, 381–393.
+- Wang et al.
+(2025a) Bo Wang, Jinkang Wei, Mingda Chen, Chong Chen, Youfang Lin, and Jie M Zhang.
+2025a.
+A Systematic Exploration of Mutation-Based Fault Localization Formulae.
+Software Testing, Verification and Reliability 35, 1 (2025), e1905.
+- Wang et al.
+(2017) Bo Wang, Yingfei Xiong, Yangqingwei Shi, Lu Zhang, and Dan Hao.
+2017.
+Faster mutation analysis via equivalence modulo states.
+In Proceedings of the 26th ACM SIGSOFT International Symposium on Software Testing and Analysis.
+295–306.
+- Wang et al. (2025b) Guancheng Wang, Qinghua Xu, Lionel C Briand, and Kui Liu. 2025b. On Mutation-Guided Unit Test Generation. CoRR abs/2506.02954 (2025).
+- Wang et al.
+(2024) Junjie Wang, Yuchao Huang, Chunyang Chen, Zhe Liu, Song Wang, and Qing Wang.
+2024.
+Software testing with large language models: Survey, landscape, and vision.
+IEEE Transactions on Software Engineering (2024).
+- Wen et al.
+(2019) Ming Wen, Yepang Liu, Rongxin Wu, Xuan Xie, Shing-Chi Cheung, and Zhendong Su.
+2019.
+Exposing library API misuses via mutation analysis.
+In 2019 IEEE/ACM 41st International Conference on Software Engineering (ICSE).
+IEEE, 866–877.
+- Wong et al.
+(2020) Chu-Pan Wong, Jens Meinicke, Leo Chen, João P Diniz, Christian Kästner, and Eduardo Figueiredo.
+2020.
+Efficiently finding higher-order mutants.
+In Proceedings of the 28th ACM Joint Meeting on European Software Engineering Conference and Symposium on the Foundations of Software Engineering.
+1165–1177.
+- Wu et al.
+(2024) Yonghao Wu, Zheng Li, Jie M Zhang, and Yong Liu.
+2024.
+ConDefects: A Complementary Dataset to Address the Data Leakage Concern for LLM-Based Fault Localization and Program Repair.
+In Companion Proceedings of the 32nd ACM International Conference on the Foundations of Software Engineering.
+642–646.
+- Xia and Zhang (2022) Chunqiu Steven Xia and Lingming Zhang.
+2022.
+Less training, more repairing please: revisiting automated program repair via zero-shot learning.
+In Proceedings of the 30th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering.
+959–971.
+- Xiao et al.
+(2023) Yuan-An Xiao, Chenyang Yang, Bo Wang, and Yingfei Xiong.
+2023.
+ExpressAPR: Efficient patch validation for Java automated program repair systems.
+In 2023 38th IEEE/ACM International Conference on Automated Software Engineering (ASE).
+IEEE, 2038–2041.
+- Xiao et al.
+(2024) Yuan-An Xiao, Chenyang Yang, Bo Wang, and Yingfei Xiong.
+2024.
+Accelerating patch validation for program repair with interception-based execution scheduling.
+IEEE Transactions on Software Engineering (2024).
+- Yuan et al.
+(2024) Zhiqiang Yuan, Mingwei Liu, Shiji Ding, Kaixin Wang, Yixuan Chen, Xin Peng, and Yiling Lou.
+2024.
+Evaluating and improving chatgpt for unit test generation.
+Proceedings of the ACM on Software Engineering 1, FSE (2024), 1703–1726.
+- Zhang et al.
+(2016) Jie Zhang, Ziyi Wang, Lingming Zhang, Dan Hao, Lei Zang, Shiyang Cheng, and Lu Zhang.
+2016.
+Predictive mutation testing.
+In Proceedings of the 25th international symposium on software testing and analysis.
+342–353.
+- Zhang et al. (2018) Jie Zhang, Lingming Zhang, Mark Harman, Dan Hao, Yue Jia, and Lu Zhang. 2018. Predictive Mutation Testing. IEEE Transactions on Software Engineering 45, 9 (2018), 898–918.
+- Zhang et al.
+(2013) Lingming Zhang, Darko Marinov, and Sarfraz Khurshid.
+2013.
+Faster mutation testing inspired by test prioritization and reduction.
+In Proceedings of the 2013 International Symposium on Software Testing and Analysis.
+235–245.
+- Zhang et al.
+(2025) Peng Zhang, Zeyu Lu, Yang Wang, Yibiao Yang, Yuming Zhou, and Mike Papadakis.
+2025.
+Enriching mutation testing with innovative method invocation mutation: Filling the crucial missing piece of the puzzle.
+IEEE Transactions on Software Engineering (2025).
+- Zhang et al.
+(2022) Peng Zhang, Yang Wang, Xutong Liu, Yanhui Li, Yibiao Yang, Ziyuan Wang, Xiaoyu Zhou, Lin Chen, and Yuming Zhou.
+2022.
+Mutant reduction evaluation: what is there and what is missing?
+ACM Transactions on Software Engineering and Methodology (TOSEM) 31, 4 (2022), 1–46.
+- Zhu et al.
+(2024) Qihao Zhu, Daya Guo, Zhihong Shao, Dejian Yang, Peiyi Wang, Runxin Xu, Y Wu, Yukun Li, Huazuo Gao, Shirong Ma, et al.
+2024.
+DeepSeek-Coder-V2: Breaking the Barrier of Closed-Source Models in Code Intelligence.
+arXiv preprint arXiv:2406.11931 (2024).
+- Zhu et al.
+(2023) Qihao Zhu, Zeyu Sun, Wenjie Zhang, Yingfei Xiong, and Lu Zhang.
+2023.
+Tare: Type-aware neural program repair.
+In 2023 IEEE/ACM 45th International Conference on Software Engineering (ICSE).
+IEEE, 1443–1455.
+- Zou et al.
+(2019) Daming Zou, Jingjing Liang, Yingfei Xiong, Michael D Ernst, and Lu Zhang.
+2019.
+An empirical study of fault localization families and their combinations.
+IEEE Transactions on Software Engineering 47, 2 (2019), 332–347.